A Complete Bibliography of Publications in International Studies in the Philosophy of Science

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

28 September 2023 Version 1.10

Title word cross-reference

$45.00 [Bus19]. _{2} [Nee02]. ΔEΔt ≥ h [Ant97]. E = mc^{2} [Flo05, Flo06, Kra06].

0 [Ano02a]. 0-521-65027-5 [Ano02a]. 0-521-65976-0 [Ano02a]. 0-7923-6305-5 [Rad02].

1 [Rad02]. 1-4020-0247-5 [Rad02]. 18th [Rue02]. 18th-century [Rue02]. 19 [Kra06]. 1925 [Bor13].

20 [Flo06]. 20th [Ara20].

4th [Uga15].

5 [Rad02].

6 [Bus19].

978 [Bus19]. 978-0-19-871853-6 [Duw16]. 978-0-19-921323-8 [Pla10]. 978-0-19-965498-7 [Bor13]. 978-0-262-03841-6 [Bus19]. 978-0-7456-5062-3 [Dud17]. 978-0-7456-5063-0 [Dud17]. 978-0-7456-5374-7 [Zha15]. 978-0-7456-5375-4 [Zha15]. 978-1-107-00397-2 [Pau12].

Abandon [Sha07]. Abduction [Paa06, Bla96, Nee99]. Absolute [Bel07]. Abstinence [Rup05]. Abstract [Kuo09, RA16, Ye10, Bro89]. Academia [Kre17]. Academic [Jac11, Kre17, Lan11, Rad02]. Accelerating [Kre17]. acceptance [Jon00]. accidental [Kos00]. accorded [Feh90]. Account [Asa13, Bae21, Bus19, Dim16, Gij09, Kow14, Max09, Web09, Dow92, Sed96, Sza03]. Accounts [BTB16, Gin12, Gur22]. Acknowledgement [Ano21, Ano22]. Action [Bul09, Par15, Alb93, Ber90, Feh88, Fuc93, Til99]. Activities [Mac04]. acts [Til99]. actual [Mil91a]. Addendum [Ano88a]. Adding [CT19, Til98]. Addition [Isl17]. Adjusting [Rec17]. advancement [Elg02]. Advances [Col20, Gao15]. Aesthetic [RB02, Rue02, Tod08, Eng01, Eng05]. aesthetics [Dav98, Mam87, McA02]. affordances [Har90]. African [OM16]. After [Bor12, Jae17, Lew16]. Against [Fuc93, Gin05, Kha05, Liu16, Lyo13, PZ20, Rei01, Row10, Sol14, Tym21, DD02, Feh98a, Pay95, Sua03]. Age [Abr16]. Agency [Buz14, Pop16, Til99]. Agenda [Sch11]. agent [Jac93]. Agnes [Fla03]. Agnostic [Psi00]. AIDS [Bab15]. Aim [Gao17]. Aim-oriented [Gao17]. aimless [D’A90]. Alan [Lon97]. alchemy [New01]. algebra [Lar94]. Algebraic [EV14]. Almagest} [Rec17]. also [Gir94]. Ampliative [Bla96]. Analogies [MR88, Har88]. Analogy [VG15, Feh88, Wyl88]. Analyses [Bae10]. Analysing [Bog04]. Analysis [Car10b, CMF16, Gon12, Rus10, Feh87, PZ14]. analysis-synthesis [Feh87]. Analytic [Bre15, Pla10, vdS14]. Animal [LS94]. Animals [McF92, McF91]. anomaly [Fri03, Lew03]. Answer [Ped18, How11]. Ante [SM94]. anthropological [Kut87]. Anthropomorphism [Buz14]. Anti [Wra08, Len05]. anti-Platonism [Len05]. Anti-realism [Wra08, Wra21]. antinomy [Sl91]. Antiquity [Zem13]. Antireductionism [Roh04]. any [Nee04]. anything [Mil91a]. Apology [Ano92a]. aporia [Sal93]. Applicability [Rus13]. Application [Sve14, Bra94]. applications [Liu98]. Applied [Boo06, Mor06, Yag17]. Applying [Bod06, Hei06, Mor06, BBB06]. appreciation [Rue02]. Approach [Bar20, BB21, Mal07, Wag13, ZG16, Bra94, FL99, Ler01, Sce98, dCF93]. Approaches [Gon12, Rus10]. Approximate [Hun11]. Approximation [CT19]. apriorism [Mis01]. Arber [Fla03]. archaeological [Wyl88]. Archaeologies [Bab15]. Argument [All17, And15, Bel07, Bus11, Ken11, Miz13, Reu13, Saa07, Sko10, WĆ06, Wra08, YKMM22, DD02, Goł99, Gre94, Mat98, Nad93, Sik92]. argumentation [Feh98b]. Arguments [Mat21, Feh98a, Hoe00]. Aristotelian [MR88]. Aristotle [Len01, Mam16, New01, Uga15, Ano02a]. Armstrong [Pag02]. Arrow [CGL05, Reu13]. Art [Ban12, Elg02, Law93]. Artefact [JB15]. Arthur [Zha15]. Article [Ci03, Lel88, Rad02]. Articulating [Rou11]. Articulation [Lyo13]. Artifacts [Fra09]. Artificial [Kee12, Jac93]. Artifictional [Tym21]. Ashgate [Kno11]. Aspects [Rea20, Cul95, Got00]. Assess [CL19]. Assessing [Kra21]. Assessment [CG12, Mäk13, Pay15]. assisted [Aga92]. assumption [Wil91]. assumptions [Wyl88]. Astronomer [De 15b]. Astrophysical [WĆ06]. Astrophysics [Suá13]. Asymmetric [CGL05]. Asymmetry [Pal16]. Atom [Bor13]. Atomic [Bor13]. atoms [Mac01, Nee04]. attack [Feh98a]. Attempt [PV20, Sch09, Mat87]. Auction [Bol12]. Author [Ano10, Ano11, Ano12b, Ano13b, Ano16]. Autonomous [Obe90]. Autonomy [Kuu12]. Axiomatic [And15].

B [Duc05]. Bachelard [Cut88]. Background [Kow17, Mit89]. Backtracking [Kut11]. BACON [Wat92].

Bad [Feh98a, Kha10, Sch22, Sha16, Nad93]. Bailer [Mac07]. Bailer-Jones [Mac07]. Bakar [Nie08]. Balancing [Til99]. Bananaworld [Duw16]. Bandyopadhyay [How16]. Barrier [Bre15]. barriers [Pea02]. Based [Bae10, Chi19, Kou15, McF92, The14]. Basics [Kee12]. Basis [Per14, Vah94]. Bayesian [BTB16, Bra17, How16, Réd92]. Bayesianism [BB10, DS12]. BC [Uga15]. Be [Bou19, Ci06, Reu13, Tod08, Bus03, Cha97, Hon91, Lel89, Mis96, Nii90, Réd92, Tuc90, Ueb01]. Bearers [DF94]. beautiful [Eng02]. Beauty [De 02, Kos02, Tod08]. Before [HS09]. begin [Nee04]. Beginning [Zin08]. Beginnings [Bab10]. Behaviour [VG15]. Behaviourism [Sle91]. Belief [Nor07, Oka00a, Psi00, Vah94]. Beliefs [Sch22, Lin93]. believe [Per13b, Vai98]. Believing [Alc10]. Bell [Fer11]. Beneath [Tod08]. Benefit [Dou00]. Bentley [Pau12, Mac01]. Berkeley [Pap88b]. Bernard [LS94]. best [Jae02, Mar89]. best-system [Jae02]. Between [Bre15, Del22, Kin17, Okr98, CW09, Dor07, Kut10, Mac92, Ses11, Yan08]. Beyond [Ban12, Del22, Sle09, Sve14, Whi15]. Big [Row08]. bin [Nie08]. Biodiversity [Ese09]. Biography [Bab15]. biological [Tho92, Tis89, dS89]. Biology [Ano02a, Ken11, Mar15, Yan08, Jan89]. Biomedical [Leo12]. Bird [Row10]. birth [Bol87, Feh88, Til87b]. bivalence [Mis90]. Black [Bok11]. blend [Lel89]. Bloch [Woo07]. block [Ear08]. Bloomsbury [vdS14]. Bloor [Sle91]. blue [Law93]. bluff [Sle91]. Board [Ano12a, Ano13a, Ano14b, Ano87a, Ano87b, Ano88b, Ano88c, Ano89a, Ano89b, Ano90a, Ano90b, Ano90c, Ano91a, Ano91b, Ano91c, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano94a, Ano94b, Ano94c, Ano95a, Ano95b, Ano95c, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano98c, Ano99a, Ano99b, Ano99c]. Bodies [Chi10a]. Body [Mor11]. Bohm [Acu16, Myr03]. Bohmian [Sko10]. Bohr [Bor13, How18, MA13, Rio17]. Book [Abe91, AKLV98, Ano00a, Ano00b, Ano00c, Ano01a, Ano01b, Ano02a, Ano02b, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano05a, Ano05b, Ano05c, Ano06a, Ano06b, BSU97, BBD09, Bor13, Bus19, CHLL07, CIL96, CBG06, Cla13, CJD^{+}08, DLVW08, Dud17, Duw16, HF92, HMO^{+}99, Ili91, JSJ^{+}94, JBM^{+}95, Kno11, KSA07, LB98, LPR99, LLOA93, Man90, MV90, MWRB98, OKHT96, Okr98, Pau12, PLO94, Pla10, PS92, STL^{+}96, Sia19, Str17, SM91, VHO^{+}07, WR93, WMB97, YTP08, Zha15, vdS14, Mam10]. Bose [Das12]. Boskovic [Cul95, Cul93, Mar90]. Boudon [Aga93, BR93, Fuc93, Kli93, Lin93, Mon93a]. Boundaries [Pao17, Lon97]. boundedness [Kuk93]. Boxes [Kas04]. Boyle [Mac92, Mac01]. Bradford [Tre21]. Bradley [Man91]. Braess [Irv93, Mar96]. Brain [Abr16, Swa11]. brand [Bro88]. Break [Hop13]. breakdown [Pea98]. breaking [Ear04]. Bridge [CE11]. Bridging [Per12]. Brittan [How16]. Broglie [Acu16]. Bruno [Dud17]. Bub [Duw16, Pea02]. Bugbear [Kow17]. Bugbears [Bab10]. Butterfield [Pau12]. butterflies [Til90b].

calculus [dS91]. called [Fri03]. Calling [Bab15]. Cambridge [Ano02a, Bus19, Duc05, Dud17, Pau12, Zha15]. Can [Bou19, Bul09, Cha97, Ci06, Hon91, Lel89, Reu13, Sha16, Tuc90, Ueb01, VG15, Web09, Bro89, Nor00, Nug91, Piz93, Réd92]. Canada [Irv97]. Cancer [Mal07, Mar05, Tha02]. Capacities [Hoh03]. Capital [Des18]. Care [And20a, BV14]. career [Tem96]. Carl [Ano98d]. Carlo [BN12]. Carnap [Ueb01]. Cartwright [Okr98, Spu01]. Case [Gon15, Leo12, Pie12, Ruz11, Feh98a, Mar89, Rue02, Tem96]. Cases [CACA19, Pay95, Sik92]. Cat [Okr98]. Categories [Hul14]. Category [JWDD12]. Catherine [Bai98]. Causal [Bau09, Ben20, BG12, CG12, CMF14, CMF16, Kuo09, Lan14, Par15, Pin10, Pop16, RA16, Rus10, Via99, Vic06, Way15, Wil19, Woo04, Dow92, Feh87, Mil91a, Mit89]. Causalism [dR06, Tuc90]. Causality [Bog04, But07, Buz14, EV14, RW07, Tei13]. Causation [Cam14, HK14, Mac04, Par16b, Per13a, Per14, Web09]. Causes [CE11, Kat13, Mel09, Pap90]. Cavendish [Par16a]. Century [Leo11, Uga15, Zem13, Car87, PZ14, Rio17, Rue02, Ara20]. Ceteris [Kow14, Kow17]. Chagas [Joy13]. Challenge [Kin11, Par15, San96]. Challenges [And20b, CACA19]. Chance [Bae13, Pal16, Woo07]. chances [Mel90]. Change [Dim16, McA08c, Roe17, She21, Ven21, Mar99]. changes [PZ14]. Changing [Kre17, PS01, Ste11]. Character [Ken22]. Characterise [Yag17]. Characteristic [GY11]. Characterization [Gon15, Pin06]. Charades [Hun11]. Chemical [Thy22]. Chemistry [Sei16, Sim16, Car87, Nee99, Nee00, Nee04, Sce98]. Chemists [Thy22]. Cherries [Miz15]. Cherry [Miz15]. Cherry-picking [Miz15]. China [Sl91]. Chinese [Qin16]. Chirality [Hoe00]. Christian [Phi14]. circa [De 02]. Circularity [Ada04]. Circulation [SS15]. Claim [Bul09]. Claims [Hop13, Wil19]. Clarke [Mac01]. Classical [All17, Bae10, Dar07, How18, Kaz14, Lar06, Pie12, Cha97, Hug99, Lar01, Tem91, dS91]. Classification [Brz16, Hul14, dP15]. Classifications [Tso07]. Classificatory [Leo12]. Claude [LS94]. clear [Law93]. Clifton [Bub03]. Climate [CL19, RS10]. Climategate [RS10]. Closer [Bru13, Miz17]. coexistence [Pea98]. Cognition [Käs15, RR95, LM95, McF91, RB02, Via99].

Cognitive [Kha05, Kou15, Mar87, PZ20, Roh04, The12, Was08, Mil91b]. Cohen [Duc05]. coherent [Lel89, Mat87]. Cold [De 92]. collapse [Fri03, Lew03]. Collection [Thy22]. College [Nar01]. collision [Lar01]. Color [Chi10b, Dan20]. Colour [JWDD12]. Combining [Bri10]. Comment [Hes92, Aga93, Kli93, Duc05]. commentary [Tra92]. commentators [JL94b]. Comments [Aga92, BR93, Gil92, Hun94, Lin93, Luc97, Mon93a, JLL92]. committal [Pin10]. Commodification [Jac11]. commonsense [Mis01]. Communities [Roe17]. commutative [Réd92]. Companion [Kno11, vdS14]. comparative [Pec02, Til98, Wil91]. Compatibility [Iva11]. Compatible [Rus10]. competitor [Lyo09]. Complementarity [Bok11, dR06]. Completeness [Vic06, Spu99]. Complexity [The14]. Complexity-based [The14]. Component [Moo12]. Componential [Kuo09]. composition [Spu01]. Computer [BK14, Dur17, Jeb17, Kro08, Par08, Aga92]. computer-assisted [Aga92]. computerized [Aga92]. Computers [JM16, Tym21]. Concept [Des18, Tho14, Ven21, Gon95, Tet89, Bus19]. Conception [And15, Liu15, Ci03]. Conceptions [PB14, PS01]. Concepts [Bae10, Kuo09, Wag09]. Conceptual [All13, Bab03, Bel07, Rou11, Ven21, ZG16, Hac95, PZ14]. Concerning [SS14]. Conclusive [Mac00]. Concrete [Leo11]. Condition [DS21]. Conditionals [Suá13]. Conditions [Bul09, Lyo09]. Confirmation [Bra17, Bri10, Ped18]. conflict [BR93]. Connection [Col22, Kut10, Mil91b]. connectionist [dS91]. Consciousness [Nes22]. consequence [Hug99]. Consequences [Ral07, Wol15]. Conservation [Des18]. conservatism [Res94]. considerations [Eng05, Ram93]. Consistency [Bok11]. constant [Cha97]. Constituting [Pea10]. Constitution [Har10]. Constraints [JWDD12, The14]. Constructing [Mar01]. Construction [Mar08, RS10, Lep88, Ram93]. constructive [BO91]. Constructivism [Rea20]. Contemporary [Kno11, WĆ06]. Content [NL16, Pri00]. Contents [Ano10, Ano11, Ano12b, Ano13b, Ano14a, Ano16]. Contested [Bus19]. Context [Mam16, Run11, RH05, O’H94]. Continent [Joy13]. Continental [Bab10, Nel20]. Contingency [Kid13]. Continuity [ZG16, Mis01]. Contours [McA08a]. Contradictions [Kaz14, Mam10]. Contributors [Ano91f, Ano95d, Ano87c, Ano87d, Ano88d, Ano88e, Ano89c, Ano89d, Ano90d, Ano90e, Ano90f, Ano91e, Ano91g, Ano92e, Ano92f, Ano93d, Ano93e, Ano94d, Ano94e]. Controlled [Kra21]. Controversial [Pan17]. Convention [Ban12]. Conventional [Sed96]. Conventionalism [Kle01, Pet93a]. Convergence [Bae10]. cooperation [BR93]. Copenhagen [MA13]. Copernican [And91]. Copernicus [De 15b]. Corpuscular [New01]. correction [Bla95]. Correspondence [Ano03d, Ano05d, Lod01]. Corroboration [Row08]. Cosmic [De 15b]. Cosmopolitanism [Rol15]. cost [McF92]. cost-based [McF92]. could [Bus03]. count [Har88]. Countered [BTB16]. Counterexample [BTB16]. Counterfactual [Bog04, Was08]. Counterfactuals [Ful22, Woo04]. counterparts [Bri99]. Counting [Fri17, Fri03, Lew03]. countries [Car98b]. Creating [Das12]. Creation [Elg02]. Creations [Fra09]. creativity [Par90]. Credence [P´ 18]. Crisis [And20b]. Criterion [BG12]. Critical [Alm07, Ano01c, Auf12, Bab10, Bai98, Des18, Eme17, Pay15, vdB14, Dav98]. Criticisms [The14]. critics [Bou93a, Sim92b]. Critique [Alb06, Lyo13, OM16, BO91, Mis01]. Cube [Bog20]. Cultural [Bre15]. Cure [And20a]. Curie [Ear04]. Curing [Tha02]. Curious [Ben22, Lan11]. Cutting [Han09].

Dadaism [Ken22]. Daniel [New01]. Daniela [Mac07]. Dartmouth [Nar01]. Darwinian [Oft10, SS15, She21]. Darwinian-Selectionist [She21]. Data [Leo12, Oka00b, Zib94]. Data-intensive [Leo12]. Davidsonian [DD02]. day [Mam16]. dealing [Lew03]. Debate [Lyo09, NL16, Rea20, Rup05, Wra21, Dav99, Lod01]. Debates [Den15, Ken11, Run11]. Debating [Hop13]. Decades [She07]. Decision [Ede10, Bra94]. Decisive [Sim19]. Deduction [PB14, Mos94]. deductive [New94]. Default [Kha10]. Defeat [How16]. Defence [Bae21, DS12, Ful22, Wra08, Dow92]. defense [Car98a, Mis01]. Defining [McF91]. Definition [Mäk13, Ray09]. Deleuze [Wil11, vdT14]. Deliver [Per13a]. Demarcation [BB21, Fas18, Han09, Let19, Yan08]. Demetra [Bro94]. Democracy [Ci11, Sl91]. democratic [Ci03]. Demon [All13]. Demonstrative [Mag08]. DENDRAL [Til90a]. Denialism [Bab15]. Denkstil [Bab03]. dependence [Ada04]. Depth [Gij09]. Descartes [Kva16]. description [DF94, McA97a]. Descriptions [Lau15]. Design [Ram93, WĆ06]. Designation [Ina14]. Despite [Roh04]. Determination [Car10a, SC13]. Determinism [Bel95, But88, Gom91]. Developing [Bar94]. Development [Käs15, Svo15, vdB14, Bol87]. Developmental [CW13]. Deviant [Han20]. diagonal [Sik92]. Diagrams [Car10b, Mey08]. Dialectical [Kaz14, Ken22]. Dialogue [BB21, The12]. Diatribe [Del22]. Dichotomy [She21]. Dictator [JB15]. Did [Zin08, Nee04, Nug96]. Difference [Mor06]. Difficulties [Mäk13]. difficulty [Les96]. Digital [Kro08].

Dilemma [FHS21]. dimension [Til98]. Disambiguating [Ill11]. Disbelief [Nor07]. Disciplines [Han20, Yan08]. Discorsi [Feh98b]. Discovering [SH92]. Discovery [Auf17, Bri10, Gil92, Mar08, Mar10a, Paa06, Roe17, Sha07, De 92, Hes92, Los92, Nee02, NS92, Pet92, Pet93a, Sim92a, Sim92b, Tra92]. Discovery/Justification [Auf17]. Discriminant [Fas18]. discussion [How11]. Disease [Joy13]. disparity [Eng01]. Dispositional [Tug13]. Dispositions [Kat13, Mel09]. disputes [Ueb01]. dissolve [Wil95]. Distance [Par15]. Distances [Car10a]. Distant [HS09]. Distinction [Auf17]. disunities [Dav96]. disunity [Hoh03, Len01]. Diversity [Ci11]. Do [Cas17, Bro89, Mil91a, Nee04, Piz93]. Doctrinal [Let22]. Doctrines [Han20]. Does [BV14, Jan89, Lyr08, Per13a, dS91]. dogma [Rei01]. Dogmas [BB10]. doing [Das12]. Dominance [JB15]. domination [Bak91]. Dordrecht [Rad02]. Double [Fri17]. Doubt [Dou00]. Doubts [Mar92]. Downward [Par16b]. dream [Til88]. dreams [Til90b]. Dretske [Pag02]. Droplets [VG15]. Dual [Nor07]. duality [Har97c]. Dubrovnik [Bro99a]. due [Zib94]. Duhem [Sha20, Wei95]. Dynamical [Dor07, Rea20]. Dynamical/Geometrical [Rea20]. Dynamics [Hei06, PZ14, Rue03, Tem91].

Early [Bai98, Mar16, Lar94, Myr03]. Earman [Hoe00]. Earth [De 15b, KBdR05, vdT14]. EBM [Tre21]. ecology [Gil91]. Econocracy [Bag17]. Economic [CACA19, CMF14, CMF16, BO91, Bol12]. Economics [Bag17, JB15, Rol15, Mon93b]. Ecuador [SS15]. edge [Pea98]. editor [Ano03d, Ano97d, Ano02c, Ano03f, Ano05e, Ano06c, De 15a, FD10, HS15, McA07, McA08b, McA09, McA10, McA11, McA12, McA13, McA14, McA15, McA16, McA17].

Editorial [Ano87a, Ano87b, Ano88b, Ano88c, Ano89a, Ano89b, Ano90a, Ano90b, Ano90c, Ano91a, Ano91b, Ano91c, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano94a, Ano94b, Ano94c, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano98c, Ano99a, Ano99b, Ano99c, Ano03e, Ano03f, Ano04c, Ano12a, Ano13a, Ano14b, Bro01, Kin18, Kin19, Kin20, NS95, Wil92, McA01]. Eds [Pla10]. Educators [Thy22]. Effectiveness [Fil16, Tre21]. Effects [Tso07]. efficacious [Til99]. Effluvia [Par15]. Ehrenhaft [Col22]. eighteenth [Car87]. eighteenth-century [Car87]. Einstein [Kra06, Bro88, Das12, DiS92, Eng02, Eng05, Flo99, Flo05, Gha17, MHB22]. Electrodynamics [Pie12]. Electron [Cla13]. Element [Thy22]. Elimination [Bla95, Bai89]. Elongations [Rec17]. Elusive [DS21]. emancipation [Bak91]. Embodiment [Pin11b]. Emergence [HT00, Pin11b, The14]. Emergent [Vic13]. Emerging [Sch11]. Emotions [TK10]. Empirical [Ale09, Miz20, Pao17, Sar00, DiS92, Vah94]. empiricality [Ada04]. Empiricism [Gao17, Uga15, Psi00, Rei01]. empiricist [BO91, Dow92]. Enantiomorphy [Rom05]. Ends [Let19]. Engineering [Sch11, Zhu17, Mar99]. English [Wol15]. Enough [CT19, Miz17]. Enquiry [God14]. Entanglement [Lam13]. Entities [Ye10, Bro89]. Entropy [Reu13]. Environmental [OM17, Vaz16]. Epigenesis [vdB14]. Epiphenomenalism [Per13a]. Epistemic [Bar20, Fol87, Fre06, Käs15, Kuk93, PV20, P´ 18, Rol15, Row10, Sve14]. Epistemological [Gon12, Rat05]. Epistemology [Ben17, Bre15, Gao15, Leo11, Mac04, Par08, Ren13, Sha20, TK10, Bla95, Mac92, Res94].

Epochal [Hop13]. Equation [Flo05, Sko10, Flo06, Kra06]. equilibrium [Bac92, BO91]. equivalence [Sar00]. Errata [Ano91d]. error [Bai89]. Errour} [Hon91]. Erwin [Mal12]. Essay [Chi10b, Mar92]. Essays [CL11, Pin11a, Pin11b, Pla10, Sei16, Thy22]. Essential [Hoc13]. essentialism [Fuc93]. Essentialist [Bus19]. Establishing [Wil19]. Ethics [OM17, Wal13, Zhu17]. Euler [PZ14]. EUR [Dud17, Rad02, Zha15]. Europe [Bjö12]. European [Gon15, McA08a, Rad15, SS15]. evaluation [Got00]. Event [Bri17]. events [Har91]. Everett [Les96]. Evidence [BTB16, CG12, Fri17, How16, Ill11, Tei13, Vaz16, Web09]. Evidential [Bog20]. Evolution [Boo16, Bou19]. Evolutionary [Gon12, Lel89, Ren13, Bla95, O’H94, Rat05]. Evolving [Yan08]. Exception [Koc09]. Exceptional [Nel20]. Exclusion [Bau09]. exile [Ueb01]. Exist [Lyr08]. Existence [Sch17, Ghi03]. Existentialism [The12]. existents [Gol93]. Experience [Dan20, Uga15, Vah94, Gil91, Got00]. experiences [Vai98]. Experiencing [Vit16]. Experiment [Jeb17, Par16b, Per21, And91, Lep88, Tis89]. Experimental [Ale09, Bae13, Boo11, Col20, Gao15, JB15, Rin12, Rou11, Goo90]. experimentation [Boo04, LS94]. Experiments [CACA19, Col22, Yan13, Art99, BN12, Bon97, Bro87, Gra18, Kuj98, Mis92, Rue02]. Experts [Bag17]. Explain [Kro08]. Explaining [Iva14, JK15, Alb93]. Explanation [Ber90, Bri10, Dan20, Dur17, For96, Gij09, Jan15, KBdR05, Oka00a, Par16b, Ral07, Rat89, She21, Tei13, Woo04, Bro89, Cor89, Cul95, Det89, Dow92, Feh87, Flo99, Hoh03, Hun01, Jon00, Mar89, Mit89, Qiu89, Sea87]. Explanationist [Saa07]. Explanations [Dor07, Har10, Pin10, RA16, Rös89]. explanative [McA89]. Explanatory [Bul09, CF06, Dav96, Dur17, Ful22, Kha10, Way15, Nee04]. Exploring [HS09]. Expose [Sha16]. Extended [Ci06, Lod01]. Extensionalism [Cha11b]. extent [Cra00]. External [Yan13]. Extra [Dan20]. Extra-Mathematical [Dan20]. extragalactic [Roc98]. Extrapolation [Bae21]. eye [Fla03, dS91].

F [Flo06]. F. [Kra06, Man91]. Facilitates [Fas18]. Fact [Bab15, Nor00]. Facticity [Ral07]. Factivist [Gur22]. Facto} [Gol20]. Facts [RS10]. Factual [Bog04]. fails [Mat98]. fall [Bak91]. Falls [Auf17]. Falsification [Tra20]. Far [VG15]. Faradayan [Dar07]. Fear [Per13a]. Festschrift [Har97a]. few [Cor92]. Feyerabend [Col22, Del22, Far00, Ken22, Kid16, KF22, SS22]. Feynman [Mey08]. Fiction [Per13b]. Fictionalism [Liu16, Pap88a]. Fictions [Pin11a, Suá13]. Field [Dar07, EV14, Lam13, Sim16, Ven21, Eng01, Goł99]. fifth [Mar89]. Fine [Fri17, WĆ06]. Fine-tuning [Fri17]. Finite [And15]. first [Rio17]. Five [Chi10a]. Fleck [Okr98, Bab03, Bab15]. flight [Bis03]. Flores [Kra06]. Fluid [Hei06, PZ14]. Focus [Ara20]. Force [Moo12, Wil11, Mar89, Nug96]. Forces [Moo12, Liu03]. Fore [Gin12]. Fore-structure [Gin12]. forerunners [Ler01]. Forking [Bic22]. Form [Kuo09, Cha04, Fla03]. Formal [Bac11, Bul09, Kaz14, Sza03, Ste10]. Formalisation [DS21]. Formation [JWDD12]. Forms [Sci16, Mos94, Til88]. forward [Far00]. Foucauldian [Mar16]. foundation [Jan87]. Foundations [All13, Del22, Fel14, Fer11, Sia19, DiS92]. Four [Lew03, Wra15]. Fraassen [Gro91, Kin11, Mar89]. Fractal [Tem96]. fragmented [Mis96]. framework [Hac95]. Franca} [Wol15]. Franklin [Par08].

Fraternity [Yan08]. Free [Sch12]. Frege [Ant12]. Fregean [Det89]. Freud [Gom91]. Friedman [Dim16, Iva11, Kin11]. frigorifick [Mac01]. frontier [Lin93]. Function [All17, Laz17, Mat21]. Functional [ES07, Ful22, Kha05, Mit89]. Fundamental [Kos00, Per14, Spu99, Nug91]. Fundamentality [Fil16]. Further [Rea20]. fusion [De 92].

G [Ano02a, Cla13]. Gadamerian [Alb06]. Galilean [All17, Kut10]. Galileo [Feh98b, Kva16, Mat87, Sea87]. Game [JB15]. Gap [Per12]. Garden [Bic22]. Gaston [Cut88]. Gauge [Fil16, Liu03]. GDR [Sch98]. Gene [Bae10]. Genealogy [Sim16]. General [Bod06, Kat13, Pri00, Bro97]. General-purpose [Pri00]. Generalisations [CMF14, CMF16]. genesis [Nug88]. Genetic [CF06]. Genetics [Pio14]. Geometric [Lam12]. Geometrical [Rea20]. Geometry [Gau05, Sia19, DiS92, Kle01, Tem96]. Gerard [Dud17]. German [Car98b]. German-speaking [Car98b]. Getting [Miz17, Pih96]. Global [Lar06, Ruh13, RS10]. Globalized [Wol15]. Glock [Pla10]. goals [LM95]. God [Mac01, Pau12]. goes [Loc97]. Gold [Wei16]. Gone [HS09]. Good [Sch22, Sha20, Bou93a, Cre93, Feh98a, LR93, Nad93]. Gordian [Han09, Let19]. Grammar [Cha11a]. Graph [Tug13]. Graph-theoretic [Tug13]. Graveyard [Miz16]. gravitation [Lar01]. Gravitational [Roc98]. gravity [Liu03, Nug88]. Greek [Zem13]. Grounds [Wra08]. Group [All17]. growing [Ear08]. Growth [Mam10]. Guidance [Sko10]. Guide [Cam14, Run11].

H [BM22, Man91, Nee02]. H. [NS92]. Haack [Vah94]. Hacker [Pla10]. Hacking [Tso07, Mar16, Roc98]. Handbook [Chi19, Ruz11]. hands [Hoe00]. Hans [Pla10, Kle01, Phi14]. Hans-Johan [Pla10]. Hansonian [Paa06]. Happen [Sch22]. hardback [Ano02a, Bor13, Bus19, Dud17, Duw16, Pla10, Rad02, Zha15]. Harmanian [Paa06]. Harmony [Saa07]. Harré [Pay95]. Healey [Rob92]. Health [CG12, RW07]. Heaviside [Bol87]. Heidegger [Sch17]. Heisenberg [Cam05]. Helge [Bor13]. Hell [Bus11]. Helmholtz [Hac95]. Hempel [Ano98d, FHS21]. Herbert [Gil92, Pau12, Hes92, Los92, Tra92]. Here [CE11]. Hermann [Gau05]. Hermeneutic [Gin12]. heroes [Til90b]. Heuristic [Aga92, JLL92, Mam87, Pet92]. Hidden [Pie12]. hierarchies [Det89]. Higgs [Lyr08]. Higher [HK14]. Higher-level [HK14]. Hill [Tre21]. Historian [Ara17]. Historians [Thy22]. Historical [Bre15, Kid13, Miz15, Rös89, Car87]. Historically [Lyo13, Wag13]. historicism [Rey99]. Histories [Leo11]. Historiography [NL16, Kut87]. History [Abr16, Ara17, Cla13, Joy13, Lar94, Lew16, Miz16, Par20, Pau12, Qin16, Sch09, Val12, Zem13, McA97a, Sch98]. Hole [Bok11, YKMM22]. holistic [Rob92]. Holmes [Par08]. Homeopathy [Bab15]. Homer [Til90a]. Horizon [Mar10a]. Howell [Mar01]. Howson [BTB16, Koc09, Ste11]. Human [Bus19, Hul14, Sci16, Alb93, Par90, Til90b]. Humanities [Bor12]. Humanity [Tym21]. Hume [How11, Sho18, Ste10]. Humean [Jae02]. Hunting [CE11]. Hybrid [Liu16]. Hyman [Pla10]. Hypotheses [dP15, Got00, How16]. Hypothesis [DS21, Miz20].

Ian [Mar16]. Ibn [Ray09]. Iconic [Amb14]. Ideal [Kra97]. Idealization [Pin11a]. Idealizations [JK15]. Ideas [PG06, SS15, Car87, De 92]. Identification [Mäk13]. Identities [Ina14]. identity [Che02, vBG88]. If [Ste10]. II [Liu98]. Image [Per12, Sim19]. Imagery [Mil91b]. images [Til87a]. Imagination [Per13b]. Immature [Hib16]. Impact [Kid13, Car87]. imperatives [Nii90]. Imperialism [CW09, CW13, Kid13, Mäk13, Rol15]. Implications [Tra20, Wyl88]. Importance [Fey22]. Imprints [Mar16]. Incident [Ben22]. Inclusive [Rad15]. Incognita [KBdR05]. incoherence [Kuk95]. Incommensurability [Alb06, Kin17, MS09, Pol17, Zhe88, Bab03, DD02, Hun01, Mil91a, San98, Mil91a]. Incongruent [Bri99]. Incorrigible [Let22]. Indeterminism [Lar01]. Index [Ano10, Ano11, Ano12b, Ano13b, Ano16]. Indexicality [Ruy21]. Indifferent [Tso07]. Indigenous [Wag09]. Indispensability [Bus11, Rus13]. Indistinguishable [Ben22]. Individuality [dS89]. individuation [Har91]. Induction [Gol20, Gre94, Mag08, Mos94, OM94, Sho18, Ste10, Wra21, Bar94, Bis03, Coh94, Gir94, Hun94, JL94a, JL94b, Mül15, Smi94]. Inductions [Miz15, Wra15, Ruh13]. Inductive [Hun11, New94, Gir94]. Industry [Gha12]. Ineffable [Zwi12]. Inertial [Acu16]. Inference [Bri10, Isl17, Kha10, Mag08, Mar89, Pag02, Réd92]. inferences [BN12]. Inferential [Bae21, Bic22, New12, New13]. Inferentialist [CMF16]. Inferring [Lan14]. inflating [Liu15]. Influence [Kut11]. Information [Fel14, Sar16, Søe19, Til90a, Tet89]. Informational [Ben20]. Informed [Lyo13, Wag13]. Inheritance [Bou19]. Instituting [Mar08]. Instrumentalism [PZ20, Sea90]. instruments [Boo04, Fje91, Kut87]. Int [Flo06, Kra06]. Integral [Rad02]. Integrate [Sch09]. Integrating [BR93]. Intellectual [Bro98, Mal12]. Intellectuals [PG06]. Intelligence [Kee12, Tym21, Jac93]. intelligibility [Luk93]. intensive [Leo12]. Intentional [Hoc13, RR95]. intentionality [Why90]. inter [Mar90]. inter-theory [Mar90]. Interaction [Kuo09, Lar06, Whi15, Mac92]. Interactions [Bok11]. Interactive [Bus19, Tso07]. Interdisciplinarity [Alv11, Pol17]. Interdisciplinary [CW13]. International [Ano15, Ano17]. Interphenomena [Bog20]. Interpretation [All17, Jan15, MA13, Per12, Thé15, Ber90, Flo06, Kra06, Pec02, Rob92, Vai98]. Interpretations [Bar20, Flo05, Kra06, Fri03]. Interpreting [RW07]. Interventionism [EB14, HK14, Reu13]. Interventionist [Bau09, Par16b, Per13a]. Interventionists [Reu13]. Interventions [Was08]. Intrinsic [BG12]. Introducing [Lin95]. Introduction [And20a, BBB06, Bro99b, Bry12, Ekw15, Eme17, Har97b, Kap10, Leg94, NSW87a, NSW87b, NSW88a, NSW88b, NSW89a, NSW89b, Pau16, Pet93b, Pio14, Rin12, TO22, Wal13, Zhu17]. Intuitions [Lor15]. Invariance [Rus14, Ear02]. invention [Bai98, Gra18]. Investigation [Auf12, Bab15, Hac95]. invisible [Bai98]. Iran [Nav07]. Iranian [PG06]. irreality [Roc98]. Irreversibility [CGL05]. ISBN [Ano02a, Bor13, Bus19, Dud17, Duw16, Pau12, Pla10, Rad02, Zha15]. Islamic [Pay15]. Islamisation [Pay15]. Isolation [GY11]. isomorphism [Sua03]. issue [Bro99b, Wil92]. Issues [And20b, Kou15, Liu98]. ivory [Mam90].

J [Cla13]. J. [Bra94, Har97a]. James [Ano02a]. Jeffrey [Duw16]. Jesus [SS15]. Johan [Pla10]. John [Pla10, Gat00]. Johnson [Coh94]. Jones [Mac07, Oka00a]. Jordi [Okr98]. Journey [Boo16, Roe17]. Judgements [Tod08, Log97]. Judgment [Lan11]. Jung [Kar15]. Jure} [Gol20]. Justice [Bjö12]. Justifiable [Sha07].

Justification [Auf17, Käs15, Nee13, San96]. Justificatory [Sch09].

Kant [Hoe00, Wat92, vdB14]. Kantian [Cam05, Fje91]. Kantianism [Dim16]. Karl [PG06]. Kathleen [Ano06d]. Keynes [Bra94]. Kind [Bir12, Nee12, Tso07]. Kindness [BJS17]. Kinds [Brz16, Ful22, God14, Hul14, dS89]. Kitcher [Psi97]. Kluwer [Rad02]. Knot [Han09, Let19]. know [dS91]. Knowability [CL11]. Knowledge [Der10, Gha12, Iva14, JM16, Koc08, Kow17, Lin22, Mam10, Mar08, O’H94, Pay15, P´ 18, Sch17, SS15, Was08, Ant97, Bol87, Kra92, Mac92, Oka00b, Pet93a, Rec17]. Kochan [How09]. Kragh [Bor13]. Krajewski [Big07, Flo06]. Kries [Hei01]. Kripke [How18]. Kronfelder [Bus19]. Kuhn [Alb06, Bab03, Ghi03, Kuu09, May22, Pol20, Ren13]. Kuhnian [Hun01, Kin17, Sim19]. Kuhnianism [Dim16].

laboratory [Mar01]. ladenness [Fje91]. Laird [Coh94]. Lakatos [Bro90]. Lakatosian [Kuu17]. Language [Gil91, Kva16]. Last [JB15]. Latour [Bab15, Dud17]. Laudan [Feh98a, Kno02, Pay95, San96]. Lawlike [BG12]. Lawmakers [Vet11]. Laws [Gol93, HS09, Hun90, Jae17, Kow14, Kow17, Lau15, Max09, McA97a, Mel90, Reu11, SM94, Vet11, Bro94, Car98a, Jae02, Nor00, Nug91, Pag02, Spu99, Spu01, Why90]. Layers [BK14]. leads [Ram93]. learn [Nor00]. Learning [Ste10]. least [Feh88]. Leaving [Bag17]. Lee [Tha02]. Left [Bus19]. Legacy [Buz10, LS94]. Leibniz [Fut02, Gom91, Lod01, Mor11, Zha15]. Lelas [Rad02, Ano03g]. Lennox [Ano02a]. lensing [Roc98]. Lesson [Ken22, Rös89]. Letter [Ano03d, De 15a, FD10, HS15]. level [HK14]. Levels [Par16b]. Lexicons [Pol20]. Life [Ano02a, Far19, Leo11, Mal12, Pau12, Sci16, Swa11, vdT14]. Lifeboat [OM16]. Light [PB14]. Like [Fas18]. Limit [All17]. Limits [CF06, Jac01, Luk93]. Lingua [Wol15]. Linguistic [Alb06, Bjö12, Pol20]. Linguistics [Gon12]. Living [Dav99, Saa07]. Local [Ruh13]. Locke [Nar01]. Logic [Kaz14, Sha07, Kuk95]. logical [Hei01, Rei01]. Lola [Okr98]. Lonely [HS09]. Long [Ken11]. Longish [Bur19]. Look [Bru13, Vic13]. Looping [Tso07]. Loops [CF06]. Loose [Let19]. Lots [Kha10]. Lucas [Har97a]. Lure [Bri17].

M [Bra94, Pla10]. MA [Bus19]. Machines [Ste06]. macrocosm [Gou87]. Macroscopic [Bur19, VG15]. Made [Bus11, Tod08]. Magnetic [Col22, Fey22]. Major [Tra20]. Make [Per13b]. Make-believe [Per13b]. Makes [Koc09]. making [Bra94, Nov10, Won97]. Mallards [God14]. Mandelstam [Pec02]. manifold [Hug99]. Manipulationism [Kow17]. Manipulative [Gel03]. Many [Baç10, Ken11, Les96, Vai98]. many-worlds [Les96, Vai98]. March [Gar94]. Maria [Bus19]. Marriage [Bus11]. Marxism [She07]. masculine [Til87b]. Massimi [Bic22]. masterpiece [Bro90]. Material [Bri10]. materialism [Jac93]. Mathematical [Ban09, Boo11, Bus11, Dan20, Duc05, Mar99, Pap88a, Van10, Wag09, Wag13, Det89, Sza03]. Mathematics [JM16, Ken13, Rus13, Tho92, Tro11]. Mathematisation [Kva16]. Mathesis [Til87b]. matter [Car87]. Matters [Yan08, Ear02, Psi00]. Maupertuis [Feh88]. Maximum [Rec17]. Maxwell [All13]. May [Tod08]. McAllister [Dav98]. Meaning [CMF14, Laz17, Swa11, BS95, Cor92, Mil91b, Mon93b, vBG88]. Means [ES07, SC13]. measured [Cha97]. Measurement [KF22]. measurements [Bok03]. Mechanical [Käs19, DF94]. Mechanically [Per12]. Mechanicism [MHB22]. Mechanics [All13, All17, Bar20, CGL05, Del22, Duw16, Fel14, Ken22, Kut10, Lar06, Laz17, MA13, Pan17, Sko10, Uga15, Cha97, For00, Fri03, Hod97, Hug99, Lew03, Luc95, Nor01, Pec02, Rob92, Rös89, Tem91]. Mechanism [Kuo09, Lyr08]. Mechanisms [Mac04, Was08]. Mechanistic [Bae13, Har10, Ill11, Web09]. Medicine [And20a, Tei13, Wil19, Qiu89]. Meets [How18]. memories [Nor00]. Memory [And15]. men [Wil91]. Mental [Mis92, Kis99, New94]. Mentzou [Bro94]. Mereological [Pin06]. Meta [Mül15, Ruh13, Sho18]. Meta-Induction [Sho18, Mül15]. Meta-inductions [Ruh13]. Metacriterion [Fas18]. Metaphor [OM16, Feh88, vBG88]. Metaphors [Wag13]. Metaphysical [Ben17, Gra18, Mar05, Rup05, Bis03, Gro91]. Metaphysics [Bur19, Den15, Liv21, Mac04, Max09, Mei14, Vet11, Pap88b]. Meteorology [New01]. Method [Del22, Gha17, Ruz11, Thé15, Vis14, Feh87, For93]. Methodological [JB15, Mat21, Res94, Car87, San96]. Methodology [Bol12, Bac92, BO91, For93, Lar94, Ram93]. mice [Wil91]. Michael [Pau12]. Microbiology [Kai15]. microcosm [Gou87]. microscope [Bai98]. Middle [Bur19]. Middle-sized [Bur19]. Mill [Nar01]. Mimesis [Ban12]. mimetic [Rue02]. Mimicking [VG15]. Mind [Fra09, HS09, Phi14, Sle09, Fla03]. Mingled [Chi10a]. minimal [Ada04]. Minkowski [Kut10]. Minkowskian [Gau05]. Miracles [Hun11, Mat98]. Mirage [Ses11]. mixed [Pap90]. Mizrahi [CT19, Isl17]. Modal [Esf09, Rea20, Rue03, Bri99]. Model [Bod06, Bor13, Chi19, Kou15, Moo12, New12, New13, Par15, RH05, Ear08, Gir94, JL94a, JL94b, Lin95, Smi94, dCF93]. Model-based [Chi19, Kou15]. Model-Theoretic [RH05]. Modeling [Pin11a]. Modelling [Sve14]. Models [BK14, Bul09, CL19, Cas17, GY11, Hei06, Hoc13, JK15, Knu10, Liu97, Liu98, Liu16, Paa06, Per13b, Qiu89, Roe17, Ste06, Tug13, Was08, BJ03, Bon97, Duc05, Gar94, Har88, Kra97, Mis92, New94, Nug88, dS91]. Modern [Abr16, Jac11, Max09, Pea10, Sar16, Sch12, Sve14, Bai98, MR88, Nug88, Rad02, Rös89, Sl91]. Modernity [SS15, Rad02]. Modes [Ale09]. modest [Ci03]. Modus [How09, Lyo13]. Molecular [Bae10]. Molyneux [Par20]. Monad [Mor11]. Monopoles [Col22]. monster [Hon91]. Monte [BN12]. most [Eng02]. motion [Mat87]. motivation [McF91]. move [Gro91]. Movement [De 15b]. MR0599360 [Duc05]. MR2106565 [Mam10]. MR2201925 [Kra06]. MR2253243 [Flo06]. MR2680848 [How11]. MR2845407 [Ste11]. Multidisciplinarity [Alv11]. Multiple [Bru13, Ful22, Kis99, SC13, Str17]. Multisemiosis [MS09]. Multiverse [Fri17]. music [Gou87]. Myth [Miz16].

N [Gat00]. Nancy [Okr98]. Nath [Das12]. nationalism [Mis97]. Natural [Bai89, Bir12, BV14, Bou19, Bru13, Brz16, Des18, Far19, God14, Hul14, Nee12, Oft10, dP15, Bon97, Kut87, Liu03, McA97a, Pag02, Pih96, RR95, dCF93]. Naturalism [Ban09, Fre06, WĆ06, Ye10, Kno02, San96, Sed96, Sla00]. naturalist [Lel89]. Naturalistic [BB21, Bor12, Kow14]. Naturalized [Koc12, Sce98]. Nature [Bri17, Bus19, Esf09, Jae17, Jan15, Kva16, Ses11, Vet11, Whi15, Bak91, DF94, Gol93, McA97a, Rad02, SM94, Phi14]. Navier [PZ14]. Navigating [Bic22]. necessity [Hun01, Mac01]. need [Jan89]. Needham [Bir12, Sce99, Sce00]. Negative [BB21]. neglected [Wyl00]. Neo [Dim16, Reu13]. Neo-Kantianism [Dim16]. Neo-Russellians [Reu13]. Neoplatonism [Mar10b]. Nesić [Ben22]. net [Fre97a, Fre97b, Fre98a, Fre98b, Fre99]. Netherlands [McA97b]. network [Mis01]. Networks [Pol20]. Neurath [Okr98, Rei01, Ueb01]. Neurobiological [Har10, Yan13]. Neurophilosophy [Sin17]. Neuroscience [Abr16, Ven21]. Neutral [Nes22, Obe90]. Neutral-Structuralist [Nes22]. neutron [Vai98]. Newcomb [Irv93, Mar96]. news [Ano97e]. Newton [Bel07, BM22, Car87, DiS92, Duc05, Feh87, Gou87, Kva16]. Newton-Smith [BM22]. Newtonian [Duc05, Die01, Duc05]. nexus [Goo90]. Niels [How18, Rio17, Bor13]. Nineteenth [Zem13, PZ14]. nineteenth-century [PZ14]. No [CE11, HS09, How09, How11, Hun11, Kas04, Flo06, Kra06, Mat98, Mil91a]. no-miracles [Mat98]. Nominal [Ray09]. Nominalism [Saa07]. Nomological [All17]. Non [BTB16, Bau09, Gur22, How16, Lyo09, Moo12, Pin10, Reu11, SS15, Sve14, Ant97, Fut02, Lar01, Réd92]. Non-Bayesian [BTB16, How16]. non-collision [Lar01]. Non-committal [Pin10]. non-commutative [Réd92]. Non-competitor [Lyo09]. Non-epistemic [Sve14]. Non-European [SS15]. Non-Factivist [Gur22]. non-observational [Ant97]. Non-reductive [Bau09, Moo12]. non-tensed [Fut02]. Non-universal [Reu11]. Nonlocality [Woo07, Pea98]. Nonrelativistic [All17]. Normative [Fre06, San96, Sla00, Ste10, Jan87, Kno02, Til87a]. Normativity [BS22]. Northpoles [Fey22]. notebook [Mar01]. Notes [Ano87c, Ano87d, Ano88d, Ano88e, Ano89c, Ano89d, Ano90d, Ano90e, Ano90f, Ano91e, Ano91f, Ano91g, Ano92e, Ano92f, Ano93d, Ano93e, Ano94d, Ano94e, Ano95d, Ano97e]. Nothing [Bac11]. Notice [Ano01c, Bai98, Ano98d, Dav98]. notices [Ano97f]. Notion [Mam16, Per13a]. Novelty [Jeb17]. Nozick [Wyl00]. Number [Sin17]. numbers [Šik96]. Nurture [Ses11, Whi15].

O [Nee02]. obeyed [Nii90]. Obituary [Ano97f, Ano98d, Ano03g, Big07, Bub03, Gat00]. objections [Myr03]. Objective [BB10, DS12]. Objectivity [Amb10, Fil16, Kuu12, Pea10, Pop16, Tul16, Wyl00]. Objects [Bur19, Haw10, Kra92, Kra97, Sza03]. Observable [She21]. Observable-Unobservable [She21]. Observation [Per21, Goo90]. observational [Ant97]. observations [Fje91]. Obsolete [Mül15]. Oct [Bus19]. Oil [VG15]. Okasha [Jon00]. Old [Fri17, Miz15, Nug96]. Oliver [Bol87]. Omniscienter [Kos02]. Once [Wil95]. One [Baç10, Ken11]. Only [FHS21, Nor00]. Ontic [Esf09]. Ontological [CACA19, Har97c]. ontologically [Pih96]. Ontologies [Leo12, Rob95]. Ontology [Ben17, Chi10b, Laz17, Rus13, Slo05, Bel95, Har90, Har95, Pea02]. Open [Leo12, Reu13]. Opposite [Bog04]. Optics [Zem13]. Optimality [Sho18]. option [Wyl00]. Order [De 15b, Jae17, Pin11b]. Organicism [Mal07, vdB14, Mar05]. oriented [CMF14, CMF16, Gao17]. Origin [Alc10, Rad02]. Origins [Ano02a, Hei01]. Ørsted [Phi14]. Orthodoxy [dR16]. Osman [Nie08]. other [BS95, Mos94, Smi94]. Otto [Rei01, Okr98]. Outcome [BB21]. outsider [Das12]. Overall [Kra21]. Overlooked [And20b]. Overpopulation [OM16]. Overview [All10]. Oxford [Bor13, Duw16, Pla10].

P [Cla13, Pla10]. page [Ano97d]. Paleontology [Bry12]. paper [Aga93, BR93, Kli93, Lin93, Mar01, Mon93a]. paperback [Ano02a, Dud17, Rad02, Zha15]. Paraconsistenciti [SS14]. paradigm [Bab03, Pea98]. Paradigms [Asa13, Hun01].

Paradox [CL11, Ped18, Irv93, Mar96, Woo07]. Paradoxes [Kas04, Mas09]. Paribus} [Kow14, Kow17]. Parochialism [Wol15]. particulars [Har95]. Past [Nee13, Nor00]. path [Tha02]. Paths [Bic22]. Patrick [Tha02]. Patterns [Ben17, Feh98b]. Paul [Far00, Ken22, Sce00]. peaceful [Pea98]. Peirce [PB14]. People [Sch22]. perceive [Mac00]. perception [Hac95]. perceptualist [Sed96]. Perfect [Bou19, Kli93]. Perils [Bag17]. peripheral [Das12]. Persistence [Kat13]. Perspectival [Bic22]. Perspective [Mas09, BO91, Pec02]. Perspectives [Pea10, Rio17, Tul16, Hun01]. Perspectivist [Ruy21]. Pertinence [Mal12]. Pessimistic [Mül15, Ruh13, Wra15, Bis03]. Peter [SS14]. Phenomena [Ste06]. Phenomenological [Bar20]. Phenomenology [The12]. Philos [Flo06, Kra06]. Philosopher [All10, De 15b, Miz16, PG06]. Philosophers [Pay95, Thy22, Feh90]. Philosophical [Alc10, Auf12, Bry12, Cas17, Cha11a, Ken22, Pau16, Pin11a, Sim16]. Philosophically [Wag13]. Philosophy [Alm07, And20a, Ano02a, Ano15, Ano17, Ara17, Ara20, Bab10, Bol12, Bre15, Cam05, Cha11b, Chi10a, Col20, Der10, Ekw15, Far19, Gin05, Gon15, Hol09, Irv97, Kai15, Kap10, Käs19, Ken11, Kno11, Knu10, Koc12, Lew16, Mar15, McA97b, McA08a, Nav07, Nel20, Okr98, Par20, Pio14, Pla10, Pol20, Rad15, Rin12, Rio17, Sch11, Sch09, Sei16, Sha16, SS22, Sle09, Tei13, The12, Val12, Van10, Vaz16, Wil11, Wol15, Yan08, vdB14, Bai98, Car98b, Cut88, Far00, Kle01, Man91, Mis01, Nar01, Nie08, PS01, Sch98, vdS14]. Physical [Thé15, De 02, Gol93, Har90, Hun01, Nor00, Sza03, Why90]. Physicalism [Bau09, Pao17, Pin06, Jac93]. Physicalism* [FHS21]. physicalist [Sza03]. Physics [Gra18, Mam16, Pao17, Pea10, Per14, Rio17, Sar16, SS22, Vic06, Way15, Die01, Eng02, HT00, Jan87, MR88, Nug91, Nug96, Rös89, Spu99, Kno11]. picking [Miz15]. pictorial [Sed96]. Picture [Ler01]. pions [Hoe00]. Place [Das17, Mam16]. Planck [Cha97]. Planetary [Car10a, Mar92]. Planets [God14]. Platonic [Bro89]. Platonism [Bro94, Len05]. Platonist [De 15b]. please [Far00]. Pluralism [Cla97]. Pluralist [Bus19]. Pluralistic [Far00]. Poincaré [dP15]. Policy [CMF14, CMF16, Vaz16, McA01]. Policy-oriented [CMF14, CMF16]. Political [Bak91]. Politics [Gin05, Okr98]. Polity [Dud17, Zha15]. Pollock [Jac93]. pomegranites [Mar92]. Ponens} [How09, Lyo13]. Popper [Auf12, Bla95, Buz10, Del22, PV20, Par90, PG06, Sce98]. Popperian [Far00, Sla00]. Populations [Oft10]. Possibility [Lam12, Hoh03]. Possible [Bri17]. Post [Bus19, Pol20, Sle09]. Post-Essentialist [Bus19]. Post-‘Linguistic [Pol20]. Post-Reductionist [Sle09]. Postcolonial [SS15]. Posteriori [JM16]. Postmodern [Gin05]. Postmodernism [Bor12]. Postmodernist [Kid16]. Power [Bul09]. powerful [Har95]. pp [Ano02a, Bor13, Bus19, Dud17, Duw16, Pau12, Pla10, Rad02, Zha15]. Practical [Ruz11]. Practice [Brz16, Cha11a, Gol20, Lew16, Mar10a, Miz20, OM17, Per21, Sch09, Van10, Wyl88]. Practices [Amb14, Boo11, Lin22, Mar05]. Pragmatic [Ale09, BB21, vBG88]. Pragmatic-Naturalistic [BB21]. Pragmatism [Alm07, MA13, Mis90]. Praise [Far19]. Precautionary [Vaz16]. predicates [Lin95]. Prediction [Lip90, Gon95]. Preferential [RH05]. pregnancy [Won97]. prejudice [Lip90]. Premises [Tra20]. preparation [Rue03]. Present [Mam16, Cra00]. Present-day [Mam16]. Press [Ano02a, Bor13, Bus19, Duc05, Dud17, Duw16, Pau12, Pla10, Zha15]. presuppositions [Mar05]. Prevail [Rup05]. Primacy [How18]. Primate [Alc10]. Primates [Duw16]. Primer [Cas17]. Principia [Duc05]. Principle [Dar07, PV20, Ste10, Vaz16, BS95, Feh88, Rei01, Dou00, Ear04]. Principles [Max09, San96]. priori [Art99, Fje91, Iva11, McA08c]. Privilege [Kha10]. Probabilistic [Web09, Got00]. probabilities [Pap90]. Probability [Lan14, Hei01, Log97, OM94]. probably [Hug99]. Problem [BB21, BS22, EB14, Fas18, Fer11, Gil92, Sho18, Ste10, Ste11, Acu16, Hes92, Irv93, Jac01, Mar96, NS92, Nor01, Pag02, Rat05, Rue03, Sim92a, Sim92b, Tra92, Vah94, Vor89]. Problems [Tra92, Hun01]. Processes [Bur19, Kro08]. Produced [Per12]. Producing [Pay15]. Production [SS15]. Professors [Lan11]. Program [Dan20]. Programme [Koc08, Feh98a, Oka00b, Sle91, Tem91]. Programmes [Pay15, For93]. Progress [CT19, CW13, Gao17, Isl17, Mar10a, Miz17, Mül15, Row10]. Project [Aga92]. Projectability [Hun90]. Projectivism [Pop16]. Prometheanism [Bak91]. Proof [Bro90]. Proofs [Car10b, For93]. Proper [CW09]. Properties [Lee14, Vic13, DF94, Kis99, Mel90]. property [Fri03]. Propositions [Sal93]. Prospects [Luc95, Ear08]. prothrombin [Mar01]. Provisional [EB14]. Pseudo [How16]. Pseudo-hypotheses [How16]. Pseudoscience [Bab15, BB21, Let22]. Pseudosciences [Fas18]. Psychiatric [Tso07]. Psychic [Gom91]. psycho [Why90]. psycho-physical [Why90]. psychological [Car98a]. Psychology [Käs15, Val12, Rat05, Sea90, Wil91]. Psychopaths [BJS17]. Ptolemaic [Car10a]. Ptolemy [Rec17]. Public [Ci06, Feh90]. Publishers [Rad02]. Pure [Yag17]. purpose [Pri00]. Pursuit [Sha20]. Push [VG15]. Putnam [Nor01]. Puzzle [Ban09, P´ 18]. Puzzles [Kas04].

Quanta [Bac11]. Quantities [For00].

Quantum

[All10, All17, Bar20, Bog20, Bok03, Bor13, CGL05, Del22, Duw16, EV14, Eng01, Fel14, Kas04, Ken22, KF22, Kut10, Lam13, Laz17, MA13, Pan17, Pea98, VG15, Acu16, DF94, For96, For00, Fri03, Hod97, Lew03, Luc95, Nor01, Pea02, Pec02, Rob92, Rös89, Vai98]. Question [Kar15, Par20]. Questioning [Nee13]. Questionnaires [CACA19]. Questions [Kra92, SH92]. Quine [Cha11b, Wei95].

R [Har97a]. Racist [Mnc17]. Radical [Rob95, She21]. Ramsey [Cru04]. Random [Tem91]. Randomised [Kra21]. rapid [Tem96]. Rational [Kuu17, P´ 18, Til99, Fuc93]. Rationalism [Auf12]. Rationality [Aga93, Ede10, LR93, PV20, Van93, Wor90, Wra21, Bou93b, Bro99b, Cre93, D’A90, Dav99, Fol87, Kli93, Lin93, Log97, Mar87, Mon93b, Sal93, Zhe88]. Ravens [Ped18, Wil95]. Raymond [Aga93, BR93, Fuc93, Kli93]. Re [Liu15]. Re-inflating [Liu15]. Reactivity [JB15]. Reading [Phi14]. Real [Chi10b, Kro08]. Real-World [Kro08]. Realisability [Ful22]. Realisation [Bru13]. Realism [Asa12, Asa13, Ben17, Ben20, Bog20, Bus11, Eme17, Esf09, Har95, Hod97, Hun11, Iva11, Koc08, Lyo09, Lyo13, Mat21, McA08c, Miz13, Pan17, Roh04, Saa07, Slo05, Cha04, Ci03, Cru04, Kle01, Kra92, Lel89, Luc95, McA89, Nor01, Psi00, Tuc90, Via99, Wra08, Wra21]. Realism/Anti [Wra21]. Realism/Anti-realism [Wra21]. realist [Fri03]. realities [Til90b]. Reality [Luc97, Roc10, Sar16, Tro11, Mam90, Sea99]. Realization [Str17, Kis99]. really [Har88, Kno02]. Reason [Gha17, Roc10, Mis96]. reasonableness [Van93]. Reasoning [Boo11, Bri10, Kou15, PB14, Sci16, Gir94, New94, dCF93]. Reasons [Bie16, Alb93, Bou93a, Cre93, Dav99, LR93, Mac00, Nad93, Piz93]. Reassessing [Ear08]. Reassessment [KF22]. rebus [SM94]. Receptive [Ven21]. Reconciled [dR06, Tuc90]. Reconciliation [Gur22, PV20]. reconfiguration [Elg02]. Reconsideration [Kin17]. reconsidered [Wei16]. Reconsidering [But07, Col22]. Reconstructing [Auf17]. Reconstruction [Bor12, Kuu17, Mat87]. recourse [McA89]. Red [Chi10b]. Reduction [ES07, KBdR05, Nee99, Nee00, RH05, Sle09, Kis99, Ram93, Sce98, Sla00]. Reductionism [CF06, Kha05, Mal07, Mar05, Rup05]. Reductionist [Sle09]. reductive [Bau09, Moo12]. Reference [Asa12, Bis03, Cru04, Mil91a, New01, Nor01, Psi97]. Referential [Bae10]. Referring [Bir12]. Refining [New13]. Reflections [Alc10, Har91, Ara17]. Reflective [Bac92]. Reforms [And20b]. refutations [For93]. Regularity [Bae13]. Regulation [BS22]. Rehabilitating [Bel99]. Reichenbach [Auf17, Bog20, Gon95, Kle01]. Reinflating [FL99]. Rejoinder [CM95]. Relating [Smi94]. Relation [Kin17, Pan17]. Relationalism [Rea20, Bel99]. relationism [Bri99, Die01]. Relations [CG12, CW09, Way15, Mar90]. Relative [Miz13]. Relativised [Iva11, McA08c]. Relativism [Zib94, San96, Tuc90, Fre06, Sci16]. Relativistic [But07, EV14, Die01, Jan89]. Relativity [Dor07, Fer11, Gha17, Kat13, BS95, Bro97, Moz00]. relevance [Wyl88]. Reliabilism [Koc08]. Reliability [CL19, Sol14, Yan13, Ada04]. Religion [Alc10, Run11]. religious [Mac92]. Remarkable [Boo16]. remarks [Alb93, Car87]. reovirus [Tha02]. Replicability [And20b]. Reply [Ben22, Bir12, Bou93a, CT19, How09, How16, Isl17, JL94b, Koc09, Oka00a, Sim92b, Ste11, Wra21, Bro94, Flo06, Fuc93, Kra06]. Report [Ano05e, Ano06c, Kin18, Kin20, McA07, McA08b, McA09, McA10, McA11, McA12, McA13, McA14, McA15, McA16, McA17, Ano02c, Ano03f]. Represent [Mey08, BJ03]. Representation [Ban12, Fra09, Gre94, Ken13, Liu15, Mas09, Mei14, Per12, Per13b, Too12, YKMM22, Cor92, LM95, Sed96, Sua03, dS91]. representational [Hun01]. Representations [Amb14]. Representative [Amb14]. Reproducibility [Bae13]. Rereading [Kuu09]. res} [SM94]. Research [CACA19, CMF14, CMF16, Gin12, Jac11, Mal07, Ruz11, For93, Mar05, Smi94, Tem91]. Resilience [Tho14]. Resiliency [Log97]. Resisting [Mis97]. resolved [Ueb01]. Response [All15, Sce99, Flo06, Jon00, Nee99, Nee00, Sce00]. Responses [Kin11]. Restoration [Lyo13]. Resultant [Moo12]. Results [Sol14]. Rethinking [Buz10, Jae17, Rad10, Wyl00]. Retroduction [PB14]. reversal [Ear02]. Review [Ano02a, Bor13, Bus19, Ci03, Cla13, Col20, Dud17, Duw16, Kno11, Lel88, Okr98, Pau12, Pla10, Rad02, Zha15, vdS14, Abe91, Ili91, Mam10, Man90]. Reviewers [Ano21, Ano22]. Reviewing [RH05]. Reviews [AKLV98, Ano00a, Ano00b, Ano00c, Ano01a, Ano01b, Ano02b, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano05b, Ano05c, Ano06a, Ano06b, BBD09, CHLL07, CBG06, CJD^{+}08, DLVW08, JSJ^{+}94, KSA07, SM91, VHO^{+}07, YTP08, Ano05a, BSU97, CIL96, HF92, HMO^{+}99, JBM^{+}95, LB98, LPR99, LLOA93, MV90, MWRB98, OKHT96, PLO94, PS92, STL^{+}96, WR93, WMB97]. revision [Jan89]. Revisited [Alb06, YKMM22, Dav99, Liu97, Wei95]. Revisiting [Kva16, Pan17]. Revolution [De 15b, Far19, NL16, And91, Bro87, Lin95, Duc05]. Revolutionary [PG06]. Revolutions [Kin11]. Richard [Zha15]. Riemannian [Gau05]. Rightful [Das17]. Rigid [Ina14]. rise [Bak91, Mis97]. Rises [Auf17]. Risky [DS21]. Road [All13]. Rob [Bub03]. Robert [Mac92]. robots [McF92]. Robustness [Sol14, Log97]. Role [JK15, dP15, Bro97, Das12, Feh88, Feh90, Fje91, Kuj98, Lep88, Mam87, Til87a, Wyl88]. Roles [Ci11]. Room [Gur22]. Rule [Isl17, Koc09]. Russellians [Reu13]. Russo [Ill11, Tre21].

S [Pla10]. Salmon [dR06]. Same [Miz15]. Samir [Jon00]. Satyendra [Das12]. Saul [How18]. Scerri [Nee99, Nee00]. schemes [Bab03]. schizophrenic [Vai98]. Schrödinger [Mal12]. Sci [Flo06, Kra06]. Science [Ale09, Alm07, Ano15, Ara17, Ara20, Bab10, Bab15, Ban12, Bod06, Bol12, Boo06, Bre15, CACA19, Cas17, Chi19, Ci11, Col20, Das17, Ekw15, Fey22, Gin05, Gon15, Han20, Hib16, Hol09, Hop13, Iva14, Jac11, JLL92, Kap10, Kar15, Kha05, KBdR05, Knu10, Koc12, Kou15, Kuu12, Leo12, Let22, Lew16, Liv21, Mar08, Mar12, Max09, McA08a, Mis01, Miz16, Mnc17, Mor06, Nav07, Nee13, NL16, Okr98, Pap88b, Pau12, Pay15, Pin11a, Qin16, Rad10, Rad15, Run11, Rus13, RS10, Sch09, Sch12, Sch17, Sha16, She07, Sim19, Sle09, Sve14, The12, Tod08, Tul16, Vaz16, Vet11, Wil11, Wol15, Yag17, dP15, Art99, Bac92, CM90, Car98b, Ci03, Cut88, D’A90, Das12, Dav96, Dav98, De 02, Elg02, Feh90, Irv97, Jac01, Kra97, Law93]. science [Len01, Lin93, Mam87, Man91, McA97b, McA02, Mil91a, Mil91b, Nar01, PS01, Rad02, Sl91, Til87a, BBB06, Ano02a, Rad02, Ano17, Rad02]. Science/Knowledge [Pay15]. Sciences [Alv11, CG12, CW09, Gon15, Hul14, RW07, vdT14, Bon97, Kut87, Tho92]. Scientific [Alc10, And15, And20b, Asa13, Ber96, BB21, Boo11, Bri10, Brz16, Bus11, Cha11a, Cha11b, Ci06, CW09, Cru04, De 92, De 15b, Dim16, Dur17, Eme17, Gao17, Gij09, Gin12, God14, Gur22, Hoc13, Isl17, Iva14, Jan15, Ken13, Kid13, Kin11, Knu10, Koc08, Kut87, Lin22, Liu15, Liu16, Lyo09, Lyo13, Mäk13, Mar08, Mar10a, Mas09, Miz17, Miz20, Mül15, New12, New13, NL16, Oka00a, Pan17, Per13b, Per21, PB14, Roe17, Row10, RS10, Sch17, Sea87, Sim16, Sim92a, Sim92b, Sol14, Sua03, Til88, Too12, Vis14, BJ03, Boo04, Bro87, Cha04, Fje91, Fol87, For93, Hun01, Kos02, Ler01, Lin95, Los92, Mac92, Mar01, Mar05, Mar87, Pet92, Pet93a, Psi00, Ram93, SH92, Zhe88, Gil92, Hes92, NS92, Tra92]. Scientism [dR16]. Scourge [Joy13]. Seagulls [Fas18]. Search [Iva11, Laz17, Sha07, JLL92]. Second [Pin11b, Sce00, Nee00]. Second-Order [Pin11b]. Selecting [Dar07]. Selection [BV14, Bou19, Bru13, Oft10, Bai89, RR95]. Selectionist [Gon12, She21]. Selective [Ped18]. Selfhood [Nes22]. Selves [Ben22]. Semantic [Isl17, FL99, Ler01, Liu97]. semantics [Rue03]. Sennert [New01]. Sense [Bou19, Sha20]. Senses [Chi10a, Kut87]. sentences [Cru04]. sets [Mac00]. Sfendoni [Bro94]. Sfendoni-Mentzou [Bro94]. Shaped [PG06]. Shared [Roe17]. Shech [Wra21]. shift [Pea98]. Should [Mis96, Nii90, Rup05, Vai98]. Shutters [Kas04]. Similarity [Too12, Sua03]. Simon [Aga92, Gil92, Hes92, JLL92, Los92, Mar92, NS92, Tra92]. Simple [Wyl88]. Simulation [Jeb17, Par08]. Simulations [BK14, Dur17, Kro08, BN12]. Simultaneity [Buz14]. since [Bro87]. Single [Fey22]. singularities [Lar01]. sized [Bur19]. Skeleton [Mag08]. sky [Law93]. slain [Hon91]. Smith [BM22]. so-called [Fri03]. Sober [Bog20]. Social [Bul09, CACA19, Ci11, Gon15, Hol09, Hul14, Lew16, Lin22, Ren13, Sch17, Sha20, Tis89, Gil91, Lin93, Res94]. Society [SS15]. sociobiology [dS90]. Sociology [Koc08, Sch17, Oka00b, Pet93a, Wor90, dS90]. Sokal [Lon97]. Solipsism [Bog20]. Solution [EB14, Fas18]. Solved [Sho18]. solves [Irv93, Mar96]. solving [Jac01, Sim92a, Sim92b, Gil92, Hes92, NS92, Tra92]. Some [Alb93, Car87, Cul95, Har91, JLL92, Let19, Myr03]. Sorites [How09]. Source [JM16]. sources [Har90]. Southpoles [Fey22]. Soviet [Bak91]. Space [Bel07, Die01, Gau05, Kut10, Ses11, Wei13, Dav99, Har97c, Hoe00, Til88]. Spaces [SS15, ZG16, Hun01]. Spaceships [Fer11]. Spacetime [Slo05, YKMM22, DiS92]. Span [Dur17]. spatial [PZ14]. speaking [Car98b]. Special [Fer11, Bro97, Bro99b, Moz00, New01, Wil92]. Specialisation [Pol17]. Specialised [Roe17]. speciation [Vor89]. Species [Wal13, Vor89, dS89]. Sphere [Ci06]. spontaneous [Ear04, Lew03]. Springer [Chi19]. Srdan [Rad02, Ano03g]. Stance [Ruy21, RR95]. State [PG06, For96, Rue03]. statement [McA01]. Statements [BG12]. Statistical [All13, Rat89, Réd92]. statistics [Das12]. Status [Hun11]. Stellar [Suá13]. step [Far00]. Stewart [Nar01]. Stimulating [The12]. Stokes [PZ14]. Stone [Abr16]. Story [Boo16]. straighten [Ant97]. Strategies [Way15, Lew03, Sol14]. strategy [Gre94]. String [Vis14]. Strong [BB10, Koc08, Feh98a, Jac93, Oka00b, Sle91]. Structural [Ben17, Ben20, Dor07, Esf09, Iva11, McA08c, Slo05]. Structuralism [Cha04]. Structuralist [Nes22]. Structure [Bor13, Haw10, Kre17, Lam13, Lan14, Mei14, ZG16, Car87, Feh87, Fri03, Gin12, Nug88]. Structures [Esf09, Tug13, Bus03]. Struggle [Whi15]. Stud [Flo06, Kra06]. Studies [Ano02a, Ano15, Ano17, Bol12, Koc12, Lew16, She07, Tul16]. Study [Miz20, Pie12, Pol20, Ruz11, Sik92]. Studying [Sch09]. Stuff [Wei13]. Style [Duc05]. Styles [Boo11, Sci16]. Sub [ES07].

Sub-types [ES07]. subjunctives [Rat89]. Substance [Mor11, Lod01]. substantival [Hoe00]. Substantivalism [But88, Goł99, Hug99]. Success [Iva14, Gel03]. supertasks [Bok03]. Supervenience [EB14, Jae02]. Support [Sol14]. Supposed [Tod08]. Surgery [Abr16]. Surrender [Tym21]. Survey [Alm07]. Sweep [She07]. Symbolic [Til88]. symbols [Tra92]. symmetric [Wei16]. Symmetries [Fil16, Ruy21, BS95, Kos00]. Symmetry [Kas04, Sko10, Che02, Ear04, Wor90]. Symposium [BBB06]. Synthesis [Mam10, Feh87]. synthetic [Bou93b, Cre93, Sal93]. System [Car10a, Kuo09, Rus10, Jae02]. Systemic [Mal07]. Systems [Lam13, Pin11b, Reu13, Rou11, Mar92, Sza03].

T [Zha15]. Tainted [Sha16]. Taper

[How16]. Taxonomic [San98]. Taxonomies [Pol20]. Taxonomy [Tra20]. Taymiyya [Ray09]. team [McA01]. Technological [Boo04, Til90b, Nii90]. Technologies [Nel20]. Technology [Bol12, Boo06, Kou15, Nel20, Obe90, Qin16, Tul16, Til90a]. teenage [Won97]. Teleology [CW13, Tet89]. tell [Nug91]. Temporal [Pal16]. tense [Moz00]. tensed [Fut02]. Terms [Max09, Cru04, Dou00]. Terra [KBdR05]. Test [Row08, Rus14]. Testing [DS21, Miz20]. Their [CL19, Fey22, Fra09, dP15, Car87, Ueb01]. Them [CE11]. themes [Far00]. Theologians [Kap10]. Theological [Alc10]. Theorem [Ant12]. Theoretic [RH05, Tug13, dCF93]. Theoretical [Dou00, Ina14, Kou15, Cru04, Flo99]. Theories [And15, Dar07, Fra09, MHB22, Mar90, Max09, Miz16, The14, Eng01, Eng02, Eng05, Flo99, Fuc93, Ler01, Lew03, Liu97, Liu98, Rat89, Til88]. Theory [All10, Bol12, Buz14, Cha11b, Dor07, EV14, Ede10, ES07, Fel14, Gha12, Goo90, How16, Hun11, Kas04, KF22, Leo12, MHB22, McA08c, Nes22, OM17, Pin11b, Rad02, Reu11, Sha20, She21, Ste10, Svo15, Thé15, Vis14, ZG16, Acu16, Ada04, Bon97, Bou93b, BO91, Coh94, Cre93, Fje91, Flo99, Fut02, Gir94, Hac95, Hei01, Hun94, JLL92, JL94a, JL94b, Jon00, Lep88, Les96, Mar90, Mat87, Mil91a, Myr03, Nug88, Oka00b, Pag02, Sal93, Smi94, Tis89, Vai98, Zib94, Rec17]. theory-acceptance [Jon00]. theory-dependence [Ada04]. theory-ladenness [Fje91]. There [BG12, Bou19, Bre15, CE11, Gur22, Kuk95, Mam90]. Thesis [Ill11, Kva16, Tre21, dR06, Wei95, Wor90]. Thingamajigs [Bir12, Nee12]. Think [Lan11]. Thinking [Jan15]. Third [Par15, Sia19, Rei01, Wyl00]. Thomas [Okr98, Ghi03]. Thomson [Cla13]. Thought [Bro87, Far19, Pau12, Art99, Bon97, Kuj98, Mis92, Mos94]. Threat [How09]. Three [Asa13]. threshold [dS91]. Time [Ant97, CGL05, Den15, Gau05, Kas04, Kre17, Moz00, Pau16, Rom05, Sim19, Vit16, Wei13, Zin08, Die01, Ear02, Fut02, Har97c, Loc97, Luc97, Til87b, Wei16]. Time-Asymmetric [CGL05]. time-symmetric [Wei16]. times [Kut10]. tithing [Bro98]. token [Jac93]. Tooley [Pag02]. Tools [Bac11, Hoc13, Liv21]. tower [And91, Mam90]. Tracks [Har90]. trading [Das12]. tradition [New01]. Traditions [Boo11]. Traits [CACA19]. trajectories [Acu16]. Transcendental [Pea10]. Transdisciplinarity [Alv11]. Transformation [Cam05, Sim19]. Transformed [Hop13]. transgressing [Lon97]. Translation [Kin17]. Travelling [May22]. Treatment [SS14, Tre21, Tha02]. trial [Bai89, Sea87]. Trials [Kra21]. Truth [Asa12, Baç10, BV14, CT19, Der10, Gur22, Hun11, Käs15, Miz17, Nov10, Roc10, Tod08, Bro90, Cor89, Nor01, Nug91, Pay95, Psi00, Sza03]. Truth-making [Nov10]. Truthmaking [Asa13]. Truths [CT19]. Tuning [WĆ06, Fri17]. Turn [Alb06, Lew16, Pol20]. Twentieth [Leo11]. Twentieth-Century [Leo11]. Twenty [Rio17]. Twenty-first-century [Rio17]. Two [BB10, Boo11, Gin12, Kuo09]. twofold [Mon93b]. Tychomancy [Lan14]. Tying [Let19]. types [ES07, Flo99, Kra92]. Tyranny [Mar12].

Uebel [Okr98]. unconstrained [Cla97].

Underconsideration [Kha10, Miz13, Wra08]. Underdetermination [Jon00, Pie12, Oka00b, Sar00, Zib94]. Understand [Whi15]. Understanding [Cat11, Cor89, Gao17, Gur22, Iva14, Jan15, New12, New13, Rou11, Kos02, Pen97]. unexpected [Acu16]. Unification [Ben17, CW13, Liu03]. Unificationism [dR06]. unified [Bon97]. uniformity [Hun90]. Unifying [Tho14]. unity [CM90, Dav96, Len01]. Univ [Duc05]. universal [Reu11]. universe [Gra18, Wei16, Ear08]. University [Ano02a, Bor13, Duw16, Jac11, Pau12, Pla10]. Unmasking [Tod08]. Unmet [And20b]. Unobservable [She21]. unobservables [Cul93]. unpredictability [Rei01]. unreal [Gel03]. Updated [Kuu17]. US$129 [Rad02]. US$22.95 [Ano02a]. US$24.95 [Dud17, Zha15]. US$42 [Rad02]. US$44.95 [Duw16]. US$62.99 [Bor13]. US$64.95 [Ano02a]. US$69.95 [Dud17, Zha15]. Use [Rec17, Web09]. User [Cam14]. Uses [Gur22]. Using [CE11, SC13].

V [Ano06d]. Validity

[Ale09, CACA19, Kra21, Yan13]. Value [Ara17, Ber96]. Values [Ci06, Rad10, Sve14, Mar87]. Variables [Wag09, Har97c]. variation [Kuj98]. Varieties [Wra15]. Varying [Dur17].

Venus [Rec17]. Verge [Das17]. Verification [Tra20]. verificationism [Bro88]. Versus [Whi15, Psi00, RA16, San96]. vi [Bor13]. Vickers [SS14]. View [Isl17, Liu16, Row10, Duc05, Liu97]. Viewpoints [Tre21]. Views [PG06, Sla00]. viii [Dud17]. vindicated [Wat92]. Virtual [Bac11, Wil11]. Virtue [Sha20]. Visual [Ven21]. Vocabulary [Roe17]. void [Mac01]. Volder [Lod01]. Volume [Ano10, Ano11, Ano12b, Ano13b, Ano14a, Ano16, Qin16]. Vries [Dud17]. vs [Bai89].

W [Gat00, Zha15]. W. [Flo06, Oka00a]. Waismann [Hei01]. War [vdT14, Wil92]. Warming [RS10]. Was [Kid16]. water [Nee02]. Waterbears [Bab10]. Watkins [Gat00]. Wave [All17, Laz17, Mat21]. Wave-Function [Mat21]. Wesley [dR06]. Western [Irv97]. Weyl [Gau05]. Where [Har88, HS09, Vic13]. Whigs [Bab10]. Wigner [Ban09]. Wilkes [Ano06d]. Will [Far00, Sch12]. William [BM22]. Williamson [Ill11, Tre21]. Wilson [Bai98]. Within [Zhu17]. Without [Bog20, Bou19]. Wittgenstein [Hei01, Pla10]. Władysław [Big07]. words [Cor92]. work [McA02, Nee04]. workable [Kli93]. Works [Iva14, Mar16]. World [Bjö12, Kro08, Lan11, Mei14, Rou11, Bai98, Ghi03, McA97a, Par90]. Worlds [Der10, HS09, May22, Les96, Vai98]. Worldview [OM16]. Worrall [San96]. worry [Ada04, Len05]. writings [Ray09]. wrong [Kno02].

xiii [Rad02]. xiv [Duw16]. xv [Pau12]. xvi [Zha15]. xxii [Pla10]. xxiii [Ano02a].

yes [Mil91a]. York [Duw16].

zone [Das12]. zoo [Bis03].

References

Abela:1991:BR

[Abe91] Paul Abela. Book review. International Studies in the Philosophy of Science, 5(3):281–283, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                              //www.tandfonline.com//doi/abs/
                                              10.1080/02698599108573400.

Abraham:2016:HBS

[Abr16] Tara H. Abraham. A history of the brain: From stone age surgery to modern neuroscience. International Studies in the Philosophy of Science, 30(1): 91–92, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                              URL http://www.tandfonline.com/
                                              doi/full/10.1080/02698595.2016.
                                              1240477.

Acuna:2016:ITB

[Acu16] Pablo Acuña. Inertial trajectories in de Broglie–Bohm quantum theory: an unexpected problem. International Studies in the Philosophy of Science, 30(3):201–230, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                              //www.tandfonline.com/doi/full/
                                              10.1080/02698595.2017.1316107.

Adam:2004:WWA

[Ada04] Matthias Adam. Why worry about theory-dependence? Circularity, minimal empiricality and reliability. International Studies in the Philosophy of Science, 18(2–3):117–132, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                              //www.tandfonline.com/doi/full/
                                              10.1080/0269859042000296486.

Agassi:1992:HCA

[Aga92] Joseph Agassi. Heuristic computerassisted, not computerized: Comments on Simon’s Project. International Studies in the Philosophy of Science, 6(1):15–18, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573404.

Agassi:1993:RCR

[Aga93] Joseph Agassi. Rationality: a comment on Raymond Boudon’s paper. International Studies in the Philosophy of Science, 7(1):21–23, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573440. See

[Bou93b] and reply [Bou93a].

Andreou:1998:BR

[AKLV98] Chrisoula Andreou, Andris Krumins, Brendan Larvor, and Andre Vellino. Book reviews. International Studies in the Philosophy of Science, 12(3):281–292, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573602.

Albert:1993:SRR

[Alb93] Hans Albert. Some remarks on reasons in explaining human action. International Studies in the Philosophy of Science, 7(1):25–27, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573441.

Albedah:2006:GCK

[Alb06] Amani Albedah. A Gadamerian critique of Kuhn’s linguistic turn: Incommensurability revisited. International Studies in the Philosophy of Science, 20(3):323–345, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600961000.

Alcorta:2010:BPS

[Alc10] Candace S. Alcorta. The believing primate: Scientific, philosophical, and theological reflections on the origin of religion. International Studies in the Philosophy of Science, 24(2):233–236, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2010.
                                        484562.

Alegre:2009:EVP

[Ale09] Maria Caamaño Alegre. Experimental validity and pragmatic modes in empirical science. International Studies in the Philosophy of Science, 23(1):19–45, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590902843377.

Allori:2010:QTP

[All10] Valia Allori. Quantum theory: a philosopher’s overview. International Studies in the Philosophy of Science, 24(3):330–333, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2010.522416.

Allori:2013:RMD

[All13] Valia Allori. The road to Maxwell’s Demon: Conceptual foundations of statistical mechanics. International Studies in the Philosophy of Science, 27(4):453–456, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.868179.

Allori:2015:R

[All15] Valia Allori. Response. International Studies in the Philosophy of Science, 29(1):94–98, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1079077.

Allori:2017:NAN

[All17] Valia Allori. A new argument for the nomological interpretation of the wave function: The Galilean group and the classical limit of nonrelativistic quantum mechanics. International Studies in the Philosophy of Science, 31(2):177–188, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1424759.

Almeder:2007:PPS

[Alm07] Robert Almeder. Pragmatism and philosophy of science: a critical survey. International Studies in the Philosophy of Science, 21(2):171–195, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701498100.

Alvargonzalez:2011:MIT

[Alv11] David Alvargonzález. Multidisciplinarity, interdisciplinarity, transdisciplinarity, and the sciences. International Studies in the Philosophy of Science, 25(4):387–403, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.623366.

Ambrosio:2010:O

[Amb10] Chiara Ambrosio. Objectivity. International Studies in the Philosophy of Science, 24(1):125–128, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903467184.

Ambrosio:2014:IRR

[Amb14] Chiara Ambrosio. Iconic representations and representative practices. International Studies in the Philosophy of Science, 28(3):255–275, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.959831.

Andersson:1991:TEC

[And91] Dr Gunnar Andersson. The tower experiment and the Copernican revolution. International Studies in the Philosophy of Science, 5(2):143–152, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573386.

Andreas:2015:FMA

[And15] Holger Andreas. A finite memory argument for an axiomatic conception of scientific theories. International Studies in the Philosophy of Science, 29(2):113–127, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1119410.

Andreoletti:2020:CCI

[And20a] Mattia Andreoletti. Care & cure. an introduction to philosophy of medicine. International Studies in the Philosophy of Science, 33(1):59–62, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1796038.

Andreoletti:2020:RCS

[And20b] Mattia Andreoletti. Replicability crisis and scientific reforms: Overlooked issues and unmet challenges. International Studies in the Philosophy of Science, 33(3):135–151, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.1943292.

Anonymous:1986:EBa

[Ano87a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 1(1):??, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573277.

Anonymous:1986:EBb

[Ano87b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 1(2):??, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573287.

Anonymous:1986:NCa

[Ano87c] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 1(1):142, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598608573286.

Anonymous:1986:NCb

[Ano87d] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 1(2):262–263, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573296.

Anonymous:1987:A

[Ano88a] Anonymous. Addendum. International Studies in the Philosophy of Science, 2(2):2, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573308.

Anonymous:1987:EBa

[Ano88b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 2(1):??, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573297.

Anonymous:1987:EBb

[Ano88c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 2(2):??, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573307.

Anonymous:1987:NCa

[Ano88d] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 2(1):115, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573306.

Anonymous:1987:NCb

[Ano88e] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 2(2):243–244, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573318.

Anonymous:1988:EBa

[Ano89a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 3(1):??, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573319.

Anonymous:1988:EBb

[Ano89b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 3(2):??, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573329.

Anonymous:1988:NCa

[Ano89c] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 3(1):117–118, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573328.

Anonymous:1988:NCb

[Ano89d] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 3(2):248–249, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598908573340.

Anonymous:1990:EBa

[Ano90a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 4(1):??, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573341.

Anonymous:1990:EBb

[Ano90b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 4(2):??, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573351.

Anonymous:1990:EBc

[Ano90c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 4(3):??, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573362.

Anonymous:1990:NCa

[Ano90d] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 4(1):113–114, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573350.

Anonymous:1990:NCb

[Ano90e] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 4(2):226, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573361.

Anonymous:1990:NCc

[Ano90f] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 4(3):325, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573372.

Anonymous:1991:EBa

[Ano91a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 5(1):??, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573373.

Anonymous:1991:EBb

[Ano91b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 5(2):??, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573382.

Anonymous:1991:EBc

[Ano91c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 5(3):??, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573392.

Anonymous:1991:E

[Ano91d] Anonymous. Errata. International Studies in the Philosophy of Science, 5(2):192, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573391.

Anonymous:1991:NCa

[Ano91e] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 5(1):93, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573381.

Anonymous:1991:NCb

[Ano91f] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 5(2):191, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573390.

Anonymous:1991:NCc

[Ano91g] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 5(3):285, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573401.

Anonymous:1992:A

[Ano92a] Anonymous. Apology. International Studies in the Philosophy of Science, 6(3):173–174, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573427.

Anonymous:1992:EBa

[Ano92b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 6(1):??, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573402.

Anonymous:1992:EBb

[Ano92c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 6(2):??, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573418.

Anonymous:1992:EBc

[Ano92d] Anonymous. Editorial board. International Studies in the Philosophy of Science, 6(3):??, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573425.

Anonymous:1992:NCa

[Ano92e] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 6(2):163, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573424.

Anonymous:1992:NCb

[Ano92f] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 6(3):255, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573436.

Anonymous:1993:EBa

[Ano93a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 7(1):??, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573437.

Anonymous:1993:EBb

[Ano93b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 7(2):??, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573456.

Anonymous:1993:EBc

[Ano93c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 7(3):??, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573465.

Anonymous:1993:NCa

[Ano93d] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 7(2):208, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573464.

Anonymous:1993:NCb

[Ano93e] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 7(3):273, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573471.

Anonymous:1994:EBa

[Ano94a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 8(1):??, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573472.

Anonymous:1994:EBb

[Ano94b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 8(2):??, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573486.

Anonymous:1994:EBc

[Ano94c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 8(3):??, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573493.

Anonymous:1994:NCa

[Ano94d] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 8(2):169, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573492.

Anonymous:1994:NCb

[Ano94e] Anonymous. Notes on contributors. International Studies in the Philosophy of Science, 8(3):277, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573501.

Anonymous:1995:EBa

[Ano95a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 9(1):??, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573502.

Anonymous:1995:EBb

[Ano95b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 9(2):??, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573510.

Anonymous:1995:EBc

[Ano95c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 9(3):??, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573516.

Anonymous:1995:NC

[Ano95d] Anonymous. Notes on contributors.

International Studies in the Philosophy of Science, 9(3):198, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573518.

Anonymous:1996:E

[Ano96a] Anonymous. Editorial. International Studies in the Philosophy of Science, 10(2):101, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573532.

Anonymous:1996:EBa

[Ano96b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 10(1):??, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573525.

Anonymous:1996:EBb

[Ano96c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 10(2):??, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573531.

Anonymous:1996:EBc

[Ano96d] Anonymous. Editorial board. International Studies in the Philosophy of Science, 10(3):??, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573538.

Anonymous:1997:EBa

[Ano97a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 11(1):??, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573545.

Anonymous:1997:EBb

[Ano97b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 11(2):??, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573555.

Anonymous:1997:EBc

[Ano97c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 11(3):??, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573566.

Anonymous:1997:EP

[Ano97d] Anonymous. Editor’s page. International Studies in the Philosophy of Science, 11(2):117, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573556.

Anonymous:1997:NN

[Ano97e] Anonymous. Notes and news. International Studies in the Philosophy of Science, 11(1):109, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573554.

Anonymous:1997:ON

[Ano97f] Anonymous. Obituary notices. International Studies in the Philosophy of Science, 11(3):315, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573575.

Anonymous:1998:EBa

[Ano98a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 12(1):??, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573576.

Anonymous:1998:EBb

[Ano98b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 12(2):??, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573585.

Anonymous:1998:EBc

[Ano98c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 12(3):??, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573595.

Anonymous:1998:ONC

[Ano98d] Anonymous. Obituary notice: [Carl Hempel (1905–1997)]. International Studies in the Philosophy of Science, 12(2):205, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573594.

Anonymous:1999:EBa

[Ano99a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 13(1):??, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573603.

Anonymous:1999:EBb

[Ano99b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 13(2):??, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573610.

Anonymous:1999:EBc

[Ano99c] Anonymous. Editorial board. International Studies in the Philosophy of Science, 13(3):??, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573620.

Anonymous:2000:BRa

[Ano00a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 14(1):83–99, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900111927.

Anonymous:2000:BRb

[Ano00b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 14(2):199–210, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590050045764.

Anonymous:2000:BRc

[Ano00c] Anonymous. Book reviews. International Studies in the Philosophy of Science, 14(3):327–336, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437845.

Anonymous:2001:BRa

[Ano01a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 15(1):99–120, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590020029332.

Anonymous:2001:BRb

[Ano01b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 15(3):311–318, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120080235.

Anonymous:2001:CN

[Ano01c] Anonymous. Critical notice. International Studies in the Philosophy of Science, 15(1):93–98, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590020029323.

Anonymous:2002:BRAa

[Ano02a] Anonymous. Book review: Aristotle’s Philosophy of Biology: Studies in the Origins of Life Science, James G. Lennox, Cambridge, Cambridge University Press, 2001. xxiii + 321 pp, £40.00, US$64.95 (hardback), ISBN 0-521-65027-5; £14.95, US$22.95 (paperback), ISBN 0-521-65976-0. International Studies in the Philosophy of Science, 16(3):291–295, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859022000013364.

Anonymous:2002:BRa

[Ano02b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 16(2):187–198, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145115.

Anonymous:2002:ER

[Ano02c] Anonymous. Editor’s report, 2001. International Studies in the Philosophy of Science, 16(2):109–110, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145052.

Anonymous:2003:BRa

[Ano03a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 17(1):95–106, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590305231.

Anonymous:2003:BRb

[Ano03b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 17(2):195–203, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859031000160658.

Anonymous:2003:BRc

[Ano03c] Anonymous. Book reviews. International Studies in the Philosophy of Science, 17(3):313–319, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859032000181583.

Anonymous:2003:CLE

[Ano03d] Anonymous. Correspondence: Letter to the editor. International Studies in the Philosophy of Science, 17(2):191–194, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859031000160649.

Anonymous:2003:E

[Ano03e] Anonymous. Editorial. International Studies in the Philosophy of Science, 17(1):3–5, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590305234.

Anonymous:2003:EER

[Ano03f] Anonymous. Editorial: Editor’s report, 2002. International Studies in the Philosophy of Science, 17(2):115–116, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859031000160577.

Anonymous:2003:OSL

[Ano03g] Anonymous. Obituary: Srdan Lelas (1939–2003). International Studies in the Philosophy of Science, 17(3):309–311, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859032000169497.

Anonymous:2004:BRa

[Ano04a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 18(1):95–108, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590412331289288.

Anonymous:2004:BRb

[Ano04b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 18(2–3):231–239, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000296549.

Anonymous:2004:E

[Ano04c] Anonymous. Editorial. International Studies in the Philosophy of Science, 18(2–3):115–116, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000296477.

Anonymous:2005:BRa

[Ano05a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 19(1):85–100, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500051183.

Anonymous:2005:BRb

[Ano05b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 19(2):213–222, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500249555.

Anonymous:2005:BRc

[Ano05c] Anonymous. Book reviews. International Studies in the Philosophy of Science, 19(3):335–344, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500462505.

Anonymous:2005:C

[Ano05d] Anonymous. Correspondence. International Studies in the Philosophy of Science, 19(2):209–212, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500249548.

Anonymous:2005:ER

[Ano05e] Anonymous. Editor’s report, 2004. International Studies in the Philosophy of Science, 19(2):101–103, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500249381.

Anonymous:2006:BRa

[Ano06a] Anonymous. Book reviews. International Studies in the Philosophy of Science, 20(1):113–124, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600641107.

Anonymous:2006:BRb

[Ano06b] Anonymous. Book reviews. International Studies in the Philosophy of Science, 20(2):233–242, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600814514.

Anonymous:2006:ER

[Ano06c] Anonymous. Editor’s report, 2005. International Studies in the Philosophy of Science, 20(2):125–127, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600814258.

Anonymous:2006:KVW

[Ano06d] Anonymous. Kathleen V. Wilkes, 1946–2003. International Studies in the Philosophy of Science, 20(1):109–111, 2006. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600641081.

Anonymous:2010:CAI

[Ano10] Anonymous. Contents and author index, volume 24, 2010. International Studies in the Philosophy of Science, 24(4):453–458, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.543361.

Anonymous:2011:VCA

[Ano11] Anonymous. Volume contents & author index. International Studies in the Philosophy of Science, 25(4):427– 431, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
624715.

Anonymous:2012:EB

[Ano12a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 26(4):??, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.757085.

Anonymous:2012:VCA

[Ano12b] Anonymous. Volume contents and author index. International Studies in the Philosophy of Science, 26(4):471– 475, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
749616.

Anonymous:2013:EB

[Ano13a] Anonymous. Editorial board. International Studies in the Philosophy of Science, 27(4):??, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.868628.

Anonymous:2013:VCA

[Ano13b] Anonymous. Volume contents and author index, 2013. International Studies in the Philosophy of Science, 27(4):463–467, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.868187.

Anonymous:2014:CV

[Ano14a] Anonymous. Contents of volume 28, 2014. International Studies in the Philosophy of Science, 28(4):449–453, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2014.
                                        979666.

Anonymous:2014:EB

[Ano14b] Anonymous. Editorial board. International Studies in the Philosophy of Science, 28(4):??, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.1007619.

Anonymous:2015:ISP

[Ano15] Anonymous. International studies in the philosophy of science. International Studies in the Philosophy of Science, 29(4):448–452, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1195154.

Anonymous:2016:VCA

[Ano16] Anonymous. Volume contents and author index, 2016. International Studies in the Philosophy of Science, 30(4):419–423, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1335936.

Anonymous:2017:ISP

[Ano17] Anonymous. International Studies in the Philosophy of Science. International Studies in the Philosophy of Science, 31(4):439–443, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565202.

Anonymous:2021:RA

[Ano21] Anonymous. Reviewers acknowledgement. International Studies in the Philosophy of Science, 34(4):??, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2123491.

Anonymous:2022:RA

[Ano22] Anonymous. Reviewers acknowledgement. International Studies in the Philosophy of Science, 35(3–4):??, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2235854.

Antonopoulos:1997:TNO

[Ant97] Constantin Antonopoulos. Time as non-observational knowledge: How to straighten out ΔEΔt ≥ h. International Studies in the Philosophy of Science, 11(2):165–183, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573561.

Antonelli:2012:FT

[Ant12] G. Aldo Antonelli. Frege’s Theorem. International Studies in the Philosophy of Science, 26(2):219–222, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.703481.

Arabatzis:2017:WIH

[Ara17] Theodore Arabatzis. What’s in it for the historian of science? Reflections on the value of philosophy of science for history of science. International Studies in the Philosophy of Science, 31(1): 69–82, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2017.
                                        1370924.

Arabatzis:2020:CPS

[Ara20] Theodore Arabatzis. 20th Century philosophy of science in focus. International Studies in the Philosophy of Science, 33(1):53–57, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2020.1784585.

Arthur:1999:TEP

[Art99] Richard Arthur. On thought experiments as a priori science. International Studies in the Philosophy of Science, 13(3):215–229, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573622.

Asay:2012:TRR

[Asa12] Jamin Asay. Truth, reference and realism. International Studies in the Philosophy of Science, 26(3):345–348, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
731735.

Asay:2013:TPS

[Asa13] Jamin Asay. Three paradigms of scientific realism: a truthmaking account. International Studies in the Philosophy of Science, 27(1):1–21, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.783971.

Aufrecht:2012:PCR

[Auf12] Monica Aufrecht. Popper’s critical rationalism: a philosophical investigation. International Studies in the Philosophy of Science, 26(2):223–225, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
703482.

Aufrecht:2017:RFR

[Auf17] Monica Aufrecht. Reichenbach Falls — and rises? Reconstructing the discovery/justification distinction. International Studies in the Philosophy of Science, 31(2):151–176, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1424760.

Babich:2003:FDK

[Bab03] Babette E. Babich. From Fleck’s Denkstil to Kuhn’s paradigm: Conceptual schemes and incommensurability. International Studies in the Philosophy of Science, 17(1):75–92, 2003. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698590305236.

Babich:2010:TCP

[Bab10] Babette Babich. Towards a critical philosophy of science: Continental beginnings and bugbears, whigs, and waterbears. International Studies in the Philosophy of Science, 24(4):343–391, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
543349.

Babich:2015:CSP

[Bab15] Babette Babich. Calling science pseudoscience: Fleck’s archaeologies of fact and Latour’s ‘biography of an investigation’ in AIDS denialism and homeopathy. International Studies in the Philosophy of Science, 29(1):1–39, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1071550.

Baccarini:1992:REM

[Bac92] Elvio Baccarini. Reflective equilibrium and methodology of science. International Studies in the Philosophy of Science, 6(3):175–180, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573428.

Bac:2010:TOM

[Baç10] Murat Baç. Truth as one and many.

International Studies in the Philosophy of Science, 24(1):122–125, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903467176.

BacelarValente:2011:VQN

[Bac11] Mario Bacelar Valente. Are virtual Quanta nothing but formal tools? International Studies in the Philosophy of Science, 25(1):39–53, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.552417.

Baetu:2010:RCG

[Bae10] Tudor M. Baetu. The referential convergence of gene concepts based on classical and molecular analyses. International Studies in the Philosophy of Science, 24(4):411–427, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.543351.

Baetu:2013:CER

[Bae13] Tudor M. Baetu. Chance, experimental reproducibility, and mechanistic regularity. International Studies in the Philosophy of Science, 27(3):253– 271, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2013.
                                        825492.

Baetu:2021:DIA

[Bae21] Tudor M. Baetu. In defence of an inferential account of extrapolation. International Studies in the Philosophy of Science, 34(2):81–100, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2021.2015678.

Bagatur:2017:EPL

[Bag17] Sine Bagatur. The econocracy: The perils of leaving economics to the experts. International Studies in the Philosophy of Science, 31(3):324–327, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2018.
                                        1463690.

Baigrie:1988:NSV

[Bai89] Brian S. Baigrie. Natural selection vs trial and error elimination. International Studies in the Philosophy of Science, 3(2):157–172, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598908573333.

Baigrie:1998:CNC

[Bai98] Brian S. Baigrie. Critical notice: Catherine Wilson’s The invisible world: Early modern philosophy and the invention of the microscope. International Studies in the Philosophy of Science, 12(2):165–174, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573590.

Bakhurst:1991:PED

[Bak91] David Bakhurst. Political emancipation and the domination of nature: The rise and fall of Soviet Prometheanism. International Studies in the Philosophy of Science, 5(3):215–226, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573395.

Bangu:2009:WPM

[Ban09] Sorin Bangu. Wigner’s puzzle for mathematical naturalism. International Studies in the Philosophy of Science, 23(3):245–263, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903196015.

Bantinaki:2012:BMC

[Ban12] Katerina Bantinaki. Beyond mimesis and convention: Representation in art and science. International Studies in the Philosophy of Science, 26(1):114– 118, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/

doi/full/10.1080/02698595.2012.
653120.

Bara:1994:DI

[Bar94] Bruno G. Bara. Developing induction. International Studies in the Philosophy of Science, 8(1):31–34, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573475.

Barzegar:2020:PAE

[Bar20] Ali Barzegar. A phenomenological approach to epistemic interpretations of quantum mechanics. International Studies in the Philosophy of Science, 33(3):175–187, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.1950483.

Baumgartner:2009:ICE

[Bau09] Michael Baumgartner. Interventionist causal exclusion and nonreductive physicalism. International Studies in the Philosophy of Science, 23(2):161–178, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903006909.

Bandyopadhyay:2010:TDS

[BB10] Prasanta S. Bandyopadhyay and Gordon Brittan, Jr. Two dogmas of strong objective Bayesianism. International Studies in the Philosophy of Science, 24(1):45–65, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903467119.

Blancke:2021:PNO

[BB21] Stefaan Blancke and Maarten Boudry.

Pseudoscience as a negative outcome of scientific dialogue: a pragmaticnaturalistic approach to the demarcation problem. International Studies in the Philosophy of Science, 34(3):183– 198, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2057777.

Bod:2006:ISA

[BBB06] Rens Bod, Mieke Boon, and Marcel Boumans. Introduction to the symposium ‘Applying Science’. International Studies in the Philosophy of Science, 20(1):1–3, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600640943.

Bergstrom:2009:BR

[BBD09] Lars Bergström, Lisa Bortolotti, and Willem B. Drees. Book reviews. International Studies in the Philosophy of Science, 23(1):109–118, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590902843419.

Belot:1995:DO

[Bel95] Gordon Belot. Determinism and ontology. International Studies in the Philosophy of Science, 9(1):85–101, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698599508573508.

Belot:1999:RR

[Bel99] Gordon Belot. Rehabilitating relationalism. International Studies in the Philosophy of Science, 13(1):35–52, 1999. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573606.

Belkind:2007:NCA

[Bel07] Ori Belkind. Newton’s conceptual argument for absolute space. International Studies in the Philosophy of Science, 21(3):271–293, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701589551.

Beni:2017:SRM

[Ben17] Majid Davoody Beni. Structural realism, metaphysical unification, and the ontology and epistemology of patterns. International Studies in the Philosophy of Science, 31(3):285–300, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2018.1463691.

Beni:2020:CIS

[Ben20] Majid D. Beni. Causal informational structural realism. International Studies in the Philosophy of Science, 33(2):117–134, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2020.1848346.

Beni:2022:CII

[Ben22] Majid D. Beni. The curious incident of indistinguishable selves: a reply to Nesić. International Studies in the Philosophy of Science, 35(3–4):261–268, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2153215.

Bergstrom:1990:EIA

[Ber90] Lars Bergström. Explanation and interpretation of action. International Studies in the Philosophy of Science, 4(1):3–15, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573342.

Bergstrom:1996:SV

[Ber96] Lars Bergström. Scientific value. International Studies in the Philosophy of Science, 10(3):189–202, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573539.

Blondeau:2012:TIC

[BG12] Julien Blondeau and Michel Ghins. Is there an intrinsic criterion for causal lawlike statements? International Studies in the Philosophy of Science, 26(4):381–401, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.748496.

Bica:2022:NMP

[Bic22] Daian Bica. Navigating Massimi’s perspectival garden with inferential forking paths. International Studies in the Philosophy of Science, 35(3–4):291– 303, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2023.
                                        2191168.

Bielik:2016:RW

[Bie16] Lukáš Bielik. Reasons why. International Studies in the Philosophy of Science, 30(4):407–409, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1331974.

Bigaj:2007:OWK

[Big07] Tomasz Bigaj. Obituary: Władysław Krajewski, 1919–2006. International Studies in the Philosophy of Science, 21(1):91–93, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701305818.

Bird:2012:RNK

[Bir12] Alexander Bird. Referring to natural kind thingamajigs, and what they are: a reply to Needham. International Studies in the Philosophy of Science, 26(1):103–109, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.653111. See

[Nee12].

Bishop:2003:PIF

[Bis03] Michael A. Bishop. The pessimistic induction, the flight to reference and the metaphysical zoo. International Studies in the Philosophy of Science, 17(2):161–178, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859031000160621.

Bailer-Jones:2003:WSM

[BJ03] Daniela M. Bailer-Jones. When scientific models represent. International Studies in the Philosophy of Science, 17(1):59–74, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590305238.

Bjorkman:2012:LJE

[Bjö12] Beyza Björkman. Linguistic justice for Europe and for the world. International Studies in the Philosophy of Science, 26(3):354–357, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.731738.

Brzovic:2017:KP

[BJS17] Zdenka Brzović, Marko Jurjako, and Predrag Sustar. The kindness of psychopaths. International Studies in the Philosophy of Science, 31(2):189–211, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2018.
1424761.

Boyer-Kassem:2014:LMC

[BK14] Thomas Boyer-Kassem. Layers of models in computer simulations. International Studies in the Philosophy of Science, 28(4):417–436, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.979670.

Blachowicz:1995:ECP

[Bla95] James Blachowicz. Elimination, correction and Popper’s evolutionary epistemology. International Studies in the Philosophy of Science, 9(1):5–17, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573503.

Blachowicz:1996:AA

[Bla96] James Blachowicz. Ampliative abduction. International Studies in the Philosophy of Science, 10(2):141–157, 1996. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573535.

Brown:2022:WHN

[BM22] James Robert Brown and Cheryl Misak. William H. Newton-Smith (1943–2023). International Studies in the Philosophy of Science, 35(2):205– 208, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2023.
                                        2208416.

Beisbart:2012:WMC

[BN12] Claus Beisbart and John D. Norton. Why Monte Carlo simulations are inferences and not experiments. International Studies in the Philosophy of Science, 26(4):403–422, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.748497.

Boylan:1991:CET

[BO91] Thomas A. Boylan and Pascal F.

O’Gorman. The critique of equilibrium theory in economic methodology: a constructive empiricist perspective. International Studies in the Philosophy of Science, 5(2):131–142, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573385.

Bod:2006:TGM

[Bod06] Rens Bod. Towards a general model of applying science. International Studies in the Philosophy of Science, 20(1):5–25, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600640950.

Bogen:2004:ACO

[Bog04] Jim Bogen. Analysing causality: The opposite of counterfactual is factual. International Studies in the Philosophy of Science, 18(1):3–26, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590412331289233.

Boge:2020:RIR

[Bog20] Florian J. Boge. Realism without interphenomena: Reichenbach’s cube, Sober’s evidential realism, and quantum solipsism. International Studies in the Philosophy of Science, 33(4):231– 246, 2020. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2021.
1964335.

Bokulich:2003:QMS

[Bok03] Alisa Bokulich. Quantum measurements and supertasks. International Studies in the Philosophy of Science, 17(2):127–136, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859031000160595.

Bokulich:2011:ICB

[Bok11] Peter Bokulich. Interactions and the consistency of black hole complementarity. International Studies in the Philosophy of Science, 25(4):371–386, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2011.
                                        623365.

Bolotovsky:1986:BDN

[Bol87] Boris M. Bolotovsky. The birth and development of new knowledge: Oliver Heaviside. International Studies in the Philosophy of Science, 1 (2):143–175, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573289.

Boldyrev:2012:PSS

[Bol12] Ivan A. Boldyrev. Philosophy of science or science and technology studies? Economic methodology and auction theory. International Studies in the Philosophy of Science, 26(3):289– 307, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2012.
                                        731732.

Boniolo:1997:UTM

[Bon97] Giovanni Boniolo. On a unified theory of models and thought experiments in natural sciences. International Studies in the Philosophy of Science, 11(2):121–142, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573558.

Boon:2004:TIS

[Boo04] Mieke Boon. Technological instruments in scientific experimentation. International Studies in the Philosophy of Science, 18(2–3):221–230, 2004. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000296530.

Boon:2006:HSA

[Boo06] Mieke Boon. How science is applied in technology. International Studies in the Philosophy of Science, 20(1):27–47, 2006. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590600640992.

Boon:2011:TSR

[Boo11] Mieke Boon. Two styles of reasoning in scientific practices: Experimental and mathematical traditions. International Studies in the Philosophy of Science, 25(3):255–278, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.605248.

Boon:2016:RJS

[Boo16] Eva Boon. A remarkable journey: The story of evolution. International Studies in the Philosophy of Science, 30(1):88–90, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1240475.

Borutti:2012:APN

[Bor12] Silvana Borutti. After postmodernism: a naturalistic reconstruction of the humanities. International Studies in the Philosophy of Science, 26(4):467–470, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.748503.

Borrelli:2013:BRN

[Bor13] Arianna Borrelli. Book review: Niels Bohr and the Quantum Atom: The Bohr Model of Atomic Structure 1913–1925 [Helge Kragh, Oxford, Oxford University Press, 2012. vi + 410 pp. ISBN 978-0-19-965498-7, £35.50, US$62.99 (hardback)]. International Studies in the Philosophy of Science, 27(2):222–224, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.813250.

Boudon:1993:MGR

[Bou93a] Raymond Boudon. More on “good reasons”: Reply to critics. International Studies in the Philosophy of Science, 7(1):87–102, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573455. See

[Bou93b, Aga93, BR93, Fuc93, Kli93, Lin93, Mon93a].

Boudon:1993:TST

[Bou93b] Raymond Boudon. Toward a synthetic theory of rationality. International Studies in the Philosophy of Science, 7(1):5–19, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573439.

Bourrat:2019:WSC

[Bou19] Pierrick Bourrat. In what sense can there be evolution by natural selection without perfect inheritance? International Studies in the Philosophy of Science, 32(1):13–31, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1623485.

Ben-Rafael:1993:ICC

[BR93] Eliezer Ben-Rafael. Integrating cooperation and conflict: Comments on Raymond Boudon’s paper. International Studies in the Philosophy of Science, 7(1):29–31, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573442. See

[Bou93b] and reply [Bou93a].

Brady:1994:AJM

[Bra94] Michael E. Brady. On the application of J. M. Keynes’s approach to decision making. International Studies in the Philosophy of Science, 8(2):99–112, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573487.

Bradley:2017:WBC

[Bra17] Darren Bradley. What is Bayesian confirmation for? International Studies in the Philosophy of Science, 31(3):229–241, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2018.1463692.

Brenner:2015:TCB

[Bre15] Anastasios Brenner. Is there a cultural barrier between historical epistemology and analytic philosophy of science? International Studies in the Philosophy of Science, 29(2):201–214, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1119421.

Brighouse:1999:ICM

[Bri99] Carolyn Brighouse. Incongruent counterparts and modal relationism. International Studies in the Philosophy of Science, 13(1):53–68, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573607.

Brigandt:2010:SRM

[Bri10] Ingo Brigandt. Scientific reasoning is material inference: Combining confirmation, discovery, and explanation.

International Studies in the Philosophy of Science, 24(1):31–43, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903467101.

Brioschi:2017:NEL

[Bri17] Maria Regina Brioschi. Nature as event: The lure of the possible. International Studies in the Philosophy of Science, 31(4):427–429, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565208.

Brown:1986:TES

[Bro87] James Robert Brown. Thought experiments since the scientific revolution. International Studies in the Philosophy of Science, 1(1):1–15, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573279.

Brown:1987:EBV

[Bro88] James Robert Brown. Einstein’s brand of verificationism. International Studies in the Philosophy of Science, 2(1):33–54, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573301.

Brown:1988:PEW

[Bro89] James Robert Brown. Platonic explanation: Or, what abstract entities can do for you. International Studies in the Philosophy of Science, 3(1):51–67, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573324.

Brown:1990:PTL

[Bro90] James Robert Brown. Proof and truth in Lakatos’s masterpiece. International Studies in the Philosophy of Science, 4(2):117–130, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573352.

Brown:1994:PLR

[Bro94] James Robert Brown. Platonism and laws: a reply to Demetra Sfendoni-Mentzou. International Studies in the Philosophy of Science, 8(3):243–246, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573498. See

[SM94].

Brown:1997:RSR

[Bro97] Harvey R. Brown. On the role of special relativity in general relativity. International Studies in the Philosophy of Science, 11(1):67–81, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573551. Festschrift

for J. R. Lucas.

Brown:1998:IT

[Bro98] James Robert Brown. Intellectual tithing. International Studies in the Philosophy of Science, 12(1):5–6, 1998. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573577.

Brown:1999:D

[Bro99a] James Robert Brown. “Dubrovnik”.

International Studies in the Philosophy of Science, 13(2):101, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573611.

Brown:1999:ISI

[Bro99b] James Robert Brown. Introduction to the special issue on rationality. International Studies in the Philosophy of Science, 13(3):213, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573621.

Brown:2001:E

[Bro01] James Robert Brown. Editorial. International Studies in the Philosophy of Science, 15(2):125–126, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120058988.

Brunnander:2013:NSM

[Bru13] Björn Brunnander. Natural selection and multiple realisation: a closer look. International Studies in the Philosophy of Science, 27(1):73–83, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.783972.

Brysse:2012:PPI

[Bry12] Keynyn Brysse. Paleontology: a philosophical introduction. International Studies in the Philosophy of Science, 26(1):111–114, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.653112.

Brzovi:2016:NKC

[Brz16] Zdenka Brzovi. Natural kinds and classification in scientific practice. International Studies in the Philosophy of Science, 30(4):412–415, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331975.

Brown:1995:MRP

[BS95] Harvey R. Brown and Roland Sypel. On the meaning of the relativity principle and other symmetries. International Studies in the Philosophy of Science, 9(3):235–253, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573522.

Bolton:2022:RNP

[BS22] Derek Bolton and Predrag Sustar. Regulation and the normativity problem. International Studies in the Philosophy of Science, 35(2):135–151, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2149050.

Belot:1997:BR

[BSU97] Gordon Belot, Spas Spassov, and Alasdair Urquhart. Book reviews. International Studies in the Philosophy of Science, 11(3):305–313, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573574.

Bandyopadhyay:2016:NBA

[BTB16] Prasanta S. Bandyopadhyay, Mark L.

Taper, and Gordon Brittan, Jr. Non-Bayesian accounts of evidence: Howson’s counterexample countered. International Studies in the Philosophy of Science, 30(3):291–298, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2017.

1316109. See reply [How16].

Bub:2003:ORC

[Bub03] Jeffrey Bub. Obituary: Rob Clifton (1964–2002). International Studies in the Philosophy of Science, 17(1):93– 94, 2003. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698590305237.

Bulle:2009:UWC

[Bul09] Nathalie Bulle. Under what conditions can formal models of social action claim explanatory power? International Studies in the Philosophy of Science, 23(1):47–64, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590902843385.

Bursten:2019:MMM

[Bur19] Julia R. S. Bursten. Macroscopic metaphysics: Middle-sized objects and longish processes. International Studies in the Philosophy of Science, 32(1):

63–64, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.
1615659.

Busch:2003:WSC

[Bus03] Jacob Busch. What structures could not be. International Studies in the Philosophy of Science, 17(3):211–225, 2003. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859032000169433.

Busch:2011:SRI

[Bus11] Jacob Busch. Scientific realism and the indispensability argument for mathematical realism: a marriage made in Hell. International Studies in the Philosophy of Science, 25(4):307– 325, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
623361.

Buskell:2019:BRW

[Bus19] Andrew Buskell. Book review: What’s Left of Human Nature? A Post-Essentialist, Pluralist, and Interactive Account of a Contested Concept, by Maria Kronfelder, Cambridge, MA, The MIT Press, Oct. 2018, 336 pp, $45.00, £38.00 (hardback), ISBN 978-0-262-03841-6. International Studies in the Philosophy of Science, 32(2):137– 140, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.
1704616.

Butterfield:1987:SD

[But88] Jeremy Butterfield. Substantivalism and determinism. International Studies in the Philosophy of Science, 2(1):10–32, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573300.

Butterfield:2007:RRC

[But07] Jeremy Butterfield. Reconsidering relativistic causality. International Studies in the Philosophy of Science, 21(3):295–328, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701589585.

Buzzoni:2010:RPH

[Buz10] Marco Buzzoni. Rethinking Popper and his legacy. International Studies in the Philosophy of Science, 24(3):309– 321, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
522413.

Buzzoni:2014:ATC

[Buz14] Marco Buzzoni. The agency theory of causality, anthropomorphism, and simultaneity. International Studies in the Philosophy of Science, 28(4):375– 395, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
979668.

Boudry:2014:NSD

[BV14] Maarten Boudry and Michael Vlerick. Natural selection does care about truth. International Studies in the Philosophy of Science, 28(1):65–77, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915651.

Caamano-Alegre:2019:OTV

[CACA19] María Caamaño-Alegre and José Caamaño-Alegre. From ontological traits to validity challenges in social science: The cases of economic experiments and research questionnaires. International Studies in the Philosophy of Science, 32(2):101–127, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1682773.

Camilleri:2005:HTK

[Cam05] Kristian Camilleri. Heisenberg and the transformation of Kantian philosophy. International Studies in the Philosophy of Science, 19(3):271–287, 2005. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590500462273.

Campaner:2014:CUG

[Cam14] Raffaella Campaner. Causation: a user’s guide. International Studies in the Philosophy of Science, 28(2):221– 223, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
932532.

Carrier:1986:NIS

[Car87] Martin Carrier. Newton’s ideas on the structure of matter and their impact on eighteenth-century chemistry: Some historical and methodological remarks. International Studies in the Philosophy of Science, 1(1):85–105, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573283.

Carrier:1998:DPL

[Car98a] Martin Carrier. In defense of psychological laws. International Studies in the Philosophy of Science, 12(3):217–232, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573597.

Carrier:1998:PSG

[Car98b] Martin Carrier. The philosophy of science in German-speaking countries. International Studies in the Philosophy of Science, 12(1):45–86, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573582.

Carman:2010:DPD

[Car10a] Christián C. Carman. On the determination of planetary distances in the Ptolemaic system. International Studies in the Philosophy of Science, 24(3):257–265, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522408.

Carter:2010:DPA

[Car10b] Jessica Carter. Diagrams and proofs in analysis. International Studies in the Philosophy of Science, 24(1):1–14, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590903467085.

Cassini:2017:HDS

[Cas17] Alejandro Cassini. How to do science with models: a philosophical primer. International Studies in the Philosophy of Science, 31(1):105–107, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1370926.

Cat:2011:UU

[Cat11] Jordi Cat. On understanding understanding. International Studies in the Philosophy of Science, 25(4):405–411, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2011.
                                        623368.

Cat:2006:BR

[CBG06] Jordi Cat, Jonathan Bain, and John Gascoigne. Book reviews. International Studies in the Philosophy of Science, 20(3):347–357, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600961018.

Cartwright:2011:HCU

[CE11] Nancy Cartwright and Sophia Efstathiou. Hunting causes and using them: Is there no bridge from here to there? International Studies in the Philosophy of Science, 25(3):223– 241, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
605245.

Carrier:2006:ELL

[CF06] Martin Carrier and Patrick Finzer.

Explanatory loops and the limits of genetic reductionism. International Studies in the Philosophy of Science, 20(3):267–283, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600960937.

Campaner:2012:EAC

[CG12] Raffaella Campaner and Maria Carla Galavotti. Evidence and the assessment of causal relations in the health sciences. International Studies in the Philosophy of Science, 26(1):27–45, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
653113.

Castagnino:2005:TAI

[CGL05] Mario Castagnino, Manuel Gadella, and Olimpia Lombardi. Time’s arrow and irreversibility in time-asymmetric quantum mechanics. International Studies in the Philosophy of Science, 19(3):223–243, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500460806.

Chang:1997:CPC

[Cha97] Hasok Chang. Can Planck’s constant be measured with classical mechanics? International Studies in the Philosophy of Science, 11(3):223–243, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573568.

Chakravartty:2004:SFS

[Cha04] Anjan Chakravartty. Structuralism as a form of scientific realism. International Studies in the Philosophy of Science, 18(2–3):151–171, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/0269859042000296503.

Chang:2011:PGS

[Cha11a] Hasok Chang. The philosophical grammar of scientific practice. International Studies in the Philosophy of Science, 25(3):205–221, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.605244.

Chatti:2011:EST

[Cha11b] Saloua Chatti. Extensionalism and scientific theory in Quine’s philosophy. International Studies in the Philosophy of Science, 25(1):1–21, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.552415.

Chester:2002:SI

[Che02] Marvin Chester. Is symmetry identity? International Studies in the Philosophy of Science, 16(2):111–124, 2002. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145061.

Chimisso:2010:FSP

[Chi10a] Cristina Chimisso. The five senses:

a philosophy of mingled bodies. International Studies in the Philosophy of Science, 24(2):226–228, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.484559.

Chirimuuta:2010:RRE

[Chi10b] Mazviita Chirimuuta. The red and the real: An essay on color ontology. International Studies in the Philosophy of Science, 24(3):339–342, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522419.

Chiffi:2019:SHM

[Chi19] Daniele Chiffi. Springer handbook of model-based science. International Studies in the Philosophy of Science, 32(1):65–67, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1615662.

Carter:2007:BR

[CHLL07] Jessica Carter, Jussi Haukioja, Mariska E. M. P. J. Leunissen, and Brendan Larvor. Book reviews. International Studies in the Philosophy of Science, 21(2):213–225, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701498175.

Ci:2003:RAM

[Ci03] Jiwei Ci. Review article: From modest realism to a democratic conception of science. International Studies in the Philosophy of Science, 17(3):301–307, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859032000169488.

Ci:2006:CSV

[Ci06] Jiwei Ci. Can scientific values be extended to the public sphere? International Studies in the Philosophy of Science, 20(2):219–231, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600814506.

Ci:2011:SSD

[Ci11] Jiwei Ci. Social science and the diversity of its roles for democracy. International Studies in the Philosophy of Science, 25(2):179–190, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.574860.

Caruana:1996:BR

[CIL96] Louis Caruana, Sam Ingus, and Brendan Larvor. Book reviews. International Studies in the Philosophy of Science, 10(2):173–179, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573537.

Code:2008:BR

[CJD^{+}08] Lorraine Code, Struan Jacobs, Deepanwita Dasgupta, Charles R. Twardy, and Rafaela Hillerbrand. Book reviews. International Studies in the Philosophy of Science, 22(1):97– 114, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698590802280969.

Costa-Leite:2011:NEK

[CL11] Alexandre Costa-Leite. New essays on the knowability paradox. International Studies in the Philosophy of Science, 25(2):194–196, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.574865.

Carrier:2019:CMH

[CL19] Martin Carrier and Johannes Lenhard.

Climate models: How to assess their reliability. International Studies in the Philosophy of Science, 32(2):81–100, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.
1644722.

Clarke:1997:PU

[Cla97] Steve Clarke. Pluralism unconstrained. International Studies in the Philosophy of Science, 11(2):143–146, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573559.

Clarke:2013:BRH

[Cla13] Imogen Clarke. Book review: A History of the Electron: J. J. and G. P. Thomson. International Studies in the Philosophy of Science, 27(1):104–107, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2013.
783973.

Carrier:1990:US

[CM90] Martin Carrier and Jürgen Mittelstrass. The unity of science. International Studies in the Philosophy of Science, 4(1):17–31, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573343.

Carrier:1995:R

[CM95] Martin Carrier and Jürgen Mittelstrass. Rejoinder. International Studies in the Philosophy of Science, 9(1):103–104, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573509.

Claveau:2014:MCG

[CMF14] François Claveau and Luis Mireles-Flores. On the meaning of causal generalisations in policy-oriented economic research. International Studies in the Philosophy of Science, 28(4):397–416, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.979669.

Claveau:2016:CGP

[CMF16] François Claveau and Luis Mireles-Flores. Causal generalisations in policy-oriented economic research: An inferentialist analysis. International Studies in the Philosophy of Science, 30(4):383–398, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331976.

Cohen:1994:JLT

[Coh94] L. Jonathan Cohen. Johnson– Laird’s theory of induction. International Studies in the Philosophy of Science, 8(1):35–36, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573476.

Colaco:2020:RAE

[Col20] David Colaço. Review of advances in experimental philosophy of science. International Studies in the Philosophy of Science, 33(1):57–59, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1795781.

Collodel:2022:EEM

[Col22] Matteo Collodel. Ehrenhaft’s experiments on magnetic monopoles: Reconsidering the Feyerabend–Ehrenhaft connection. International Studies in the Philosophy of Science, 35(1):69– 94, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2123099.

Corbi:1988:UTE

[Cor89] Josep Corbí. Understanding, truth, and explanation. International Studies in the Philosophy of Science, 3(1):19–34, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573322.

Cordeschi:1992:FWR

[Cor92] Roberto Cordeschi. A few words on representation and meaning. International Studies in the Philosophy of Science, 6(1):19–21, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573405.

Craig:2000:EP

[Cra00] William Craig. The extent of the present. International Studies in the Philosophy of Science, 14(2):165–185, 2000. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590050045746.

Crespi:1993:GRS

[Cre93] Franco Crespi. The good reasons of the synthetic theory of rationality. International Studies in the Philosophy of Science, 7(1):33–35, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573443.

Cruse:2004:SRR

[Cru04] Pierre Cruse. Scientific realism, Ramsey sentences and the reference of theoretical terms. International Studies in the Philosophy of Science, 18(2–3):133–149, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000296495.

Cevolani:2019:WAT

[CT19] Gustavo Cevolani and Luca Tambolo. Why adding truths is not enough: A reply to Mizrahi on progress as approximation to the truth. International Studies in the Philosophy of Science, 32(2):129–135, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.

1623502. See [Miz17].

Culjak:1993:BU

[Cul93] Zvonimir Culjak. Bosković’s unobservables. International Studies in the Philosophy of Science, 7(3):211–224, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698599308573466.

Culjak:1995:SAE

[Cul95] Zvonimir Culjak. Some aspects of explanation in boskovic. International Studies in the Philosophy of Science, 9(1):73–84, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573507.

Cutting:1987:GBP

[Cut88] Gary Cutting. Gaston Bachelard’s philosophy of science. International Studies in the Philosophy of Science, 2(1):55–71, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573302.

Clarke:2009:SIP

[CW09] Steve Clarke and Adrian Walsh. Scientific imperialism and the proper relations between the sciences. International Studies in the Philosophy of Science, 23(2):195–207, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903007170.

Clarke:2013:IPD

[CW13] Steve Clarke and Adrian Walsh. Imperialism, progress, developmental teleology, and interdisciplinary unification. International Studies in the Philosophy of Science, 27(3):341–351, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2013.
825493.

DAgostino:1990:ARS

[D’A90] Fred D’Agostino. The aimless rationality of science. International Studies in the Philosophy of Science, 4(1):33–50, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573344.

Danne:2020:EMP

[Dan20] Nicholas Danne. An extramathematical program explanation of color experience. International Studies in the Philosophy of Science, 33(3):153–173, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.1951637.

Darrigol:2007:FPS

[Dar07] Olivier Darrigol. A Faradayan principle for selecting classical field theories. International Studies in the Philosophy of Science, 21(1):35–55, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701305768.

Dasgupta:2012:CPT

[Das12] Deepanwita Dasgupta. Creating a peripheral trading zone: Satyendra Nath Bose and Bose–Einstein statistics, doing science in the role of an outsider. International Studies in the Philosophy of Science, 26(3):259–287, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.731731.

Dasgupta:2017:RPS

[Das17] Deepanwita Dasgupta. The rightful place of science: Science on the verge. International Studies in the Philosophy of Science, 31(1):108–110, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1370927.

Davies:1996:EDU

[Dav96] David Davies. Explanatory disunities and the unity of science. International Studies in the Philosophy of Science, 10(1):5–21, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573526.

Davies:1998:MAS

[Dav98] David Davies. McAllister’s aesthetics in science: a critical notice. International Studies in the Philosophy of Science, 12(1):25–32, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573580.

Davies:1999:LSR

[Dav99] David Davies. Living in the “space of reasons”: The “rationality debate” revisited. International Studies in the Philosophy of Science, 13(3):231–244, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573623.

daCosta:1993:MTA

[dCF93] Newton C. A. da Costa and Steven French. A model theoretic approach to ‘natural’ reasoning. International Studies in the Philosophy of Science, 7(2):177–190, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573462.

Douven:2002:DAA

[DD02] Igor Douven and Henk W. De Regt. A Davidsonian argument against incommensurability. International Studies in the Philosophy of Science, 16(2):157–169, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145098.

DeMey:1992:SDC

[De 92] Marc De Mey. Scientific discovery: Cold fusion of ideas? International Studies in the Philosophy of Science, 6(1):23–27, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573406.

DeRegt:2002:BPS

[De 02] Henk W. De Regt. Beauty in physical science circa 2000. International Studies in the Philosophy of Science, 16(1):95–103, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120118855.

DeGroot:2015:LE

[De 15a] Jean De Groot. Letter to the Editor.

International Studies in the Philosophy of Science, 29(4):431–433, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1195146.

DePace:2015:CPA

[De 15b] Anna De Pace. Copernicus: Platonist astronomer–philosopher. Cosmic order, the movement of the Earth, and the scientific revolution. International Studies in the Philosophy of Science, 29(3):337–340, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1179044.

DelSanto:2022:BMD

[Del22] Flavio Del Santo. Beyond method: The diatribe between Feyerabend and Popper over the foundations of quantum mechanics. International Studies in the Philosophy of Science, 35(1):5– 22, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2022.
                                        2031430.

Deng:2015:DMT

[Den15] Natalja Deng. Debates in the metaphysics of time. International Studies in the Philosophy of Science, 29(3):340–344, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1179045.

Deroy:2010:WTP

[Der10] Ophelia Deroy. Worlds of truth: a philosophy of knowledge. International Studies in the Philosophy of Science, 24(4):446–448, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2010.543357.

DesRoches:2018:CCC

[Des18] C. Tyler DesRoches. On the concept and conservation of critical natural capital. International Studies in the Philosophy of Science, 32(3–4):207– 228, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2020.
1788347.

Detlefsen:1988:FHM

[Det89] Michael Detlefsen. Fregean hierarchies and mathematical explanation. International Studies in the Philosophy of Science, 3(1):97–116, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573327.

Dotson:1994:BPQ

[DF94] Allen Dotson and Henry Folse. Bearers of properties in the quantum mechanical description of nature. International Studies in the Philosophy of Science, 8(3):179–194, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573494.

Dieks:2001:STR

[Die01] Dennis Dieks. Space–time relationism in Newtonian and relativistic physics. International Studies in the Philosophy of Science, 15(1):5–17, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590020029279.

Dimitrakos:2016:KNK

[Dim16] Thodoris Dimitrakos. Kuhnianism and neo-Kantianism: On Friedman’s account of scientific change. International Studies in the Philosophy of Science, 30(4):361–382, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1331977.

DiSalle:1992:ENE

[DiS92] Robert DiSalle. Einstein, Newton and the empirical foundations of spacetime geometry. International Studies in the Philosophy of Science, 6(3):181–189, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573429.

Dowell:2008:BR

[DLVW08] Janice Dowell, Soazig Le Bihan, Janet Vertesi, and Sally Wyatt. Book reviews. International Studies in the Philosophy of Science, 22(3):327–339, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590802567431.

Dorato:2007:RTB

[Dor07] Mauro Dorato. Relativity theory between structural and dynamical explanations. International Studies in the Philosophy of Science, 21(1):95–102, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590701305891.

Douven:2000:TTP

[Dou00] Igor Douven. Theoretical terms and the Principle of the Benefit of Doubt. International Studies in the Philosophy of Science, 14(2):135–146, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590050045728.

Dowe:1992:EDC

[Dow92] Phil Dowe. An empiricist defence of the causal account of explanation. International Studies in the Philosophy of Science, 6(2):123–128, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573420.

dePaz:2015:PCH

[dP15] María de Paz. Poincaré’s classification of hypotheses and their role in natural science. International Studies in the Philosophy of Science, 29(4):369– 382, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1195142.

deRegt:2006:WSC

[dR06] Henk W. de Regt. Wesley Salmon’s complementarity thesis: Causalism and unificationism reconciled? International Studies in the Philosophy of Science, 20(2):129–147, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814308.

deRidder:2016:SNO

[dR16] Jeroen de Ridder. Scientism: The new orthodoxy. International Studies in the Philosophy of Science, 30(1):93– 95, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2016.
1240479.

deSousa:1988:KKI

[dS89] Ronald de Sousa. Kinds of kinds: Individuality and biological species. International Studies in the Philosophy of Science, 3(2):119–135, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573331.

deSousa:1990:SS

[dS90] Ronald de Sousa. The sociology of sociobiology. International Studies in the Philosophy of Science, 4(3):271–283, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573367.

deSousa:1991:DEK

[dS91] Ronald de Sousa. Does the eye know calculus? The threshold of representation in classical and connectionist models. International Studies in the Philosophy of Science, 5(2):171–185, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573388.

Dziurosz-Serafinowicz:2012:DOB

[DS12] Patryk Dziurosz-Serafinowicz. In defence of objective Bayesianism. International Studies in the Philosophy of Science, 26(3):348–351, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.731736.

Diez:2021:EFR

[DS21] José Díez and Albert Solé. On the elusive formalisation of the risky condition for hypothesis testing. International Studies in the Philosophy of Science, 34(4):199–219, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2084705.

Ducheyne:2005:MMN

[Duc05] Steffen Ducheyne. Mathematical models in Newton’s Principia: a new view of the ‘Newtonian style’. Comment on: The Newtonian revolution [Cambridge Univ. Press, Cambridge, 1980; MR0599360] by I. B. Cohen. International Studies in the Philosophy of Science, 19(1):1–19, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500051035.

Dudhwala:2017:BRB

[Dud17] Farzana Dudhwala. Book review: Bruno Latour [Gerard de Vries, Cambridge, Polity Press, 2016, viii + 221 pp., ISBN 978-0-7456-5062-3, £55.00, US$69.95, EUR 76.28 (hardback); ISBN 978-0-7456-5063-0, £16.99, US$24.95, EUR 23.56 (paperback)]. International Studies in the Philosophy of Science, 31(1):103–105, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2017.
1370928.

Duran:2017:VES

[Dur17] Juan M. Durán. Varying the explanatory span: Scientific explanation for computer simulations. International Studies in the Philosophy of Science, 31(1):27–45, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1370929.

Duwell:2016:BRB

[Duw16] Armond Duwell. Book review: Bananaworld: Quantum Mechanics for Primates [Jeffrey Bub, New York, Oxford University Press, 2016, xiv + 273 pp. ISBN 978-0-19-871853-6, US$44.95, £25.00 (hardback)]. International Studies in the Philosophy of Science, 30(3):309–311, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1316110.

Earman:2002:WTR

[Ear02] John Earman. What time reversal invariance is and why it matters. International Studies in the Philosophy of Science, 16(3):245–264, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859022000013328.

Earman:2004:CPS

[Ear04] John Earman. Curie’s Principle and spontaneous symmetry breaking. International Studies in the Philosophy of Science, 18(2–3):173–198, 2004. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/0269859042000311299.

Earman:2008:RPG

[Ear08] John Earman. Reassessing the prospects for a growing block model of the Universe. International Studies in the Philosophy of Science, 22(2):135–164, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802496680.

Eronen:2014:ISN

[EB14] Markus I. Eronen and Daniel S.

Brooks. Interventionism and supervenience: a new problem and provisional solution. International Studies in the Philosophy of Science, 28(2):185–202, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
932529.

Eder:2010:DTR

[Ede10] Anna-Maria A. Eder. Decision theory and rationality. International Studies in the Philosophy of Science, 24(3):326–329, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522415.

Ekweke:2015:IPS

[Ekw15] Evaristus Egwuatu Ekweke. An introduction to the philosophy of science. International Studies in the Philosophy of Science, 29(4):435–438, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1195147.

Elgin:2002:CRA

[Elg02] Catherine Z. Elgin. Creation as reconfiguration: Art in the advancement of science. International Studies in the Philosophy of Science, 16(1):13–25, 2002. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120118792.

Emedolu:2017:CIS

[Eme17] Christian C. Emedolu. A critical introduction to scientific realism. International Studies in the Philosophy of Science, 31(2):213–215, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2018.1424764.

Engler:2001:QFT

[Eng01] Gideon Engler. Quantum field theories and aesthetic disparity. International Studies in the Philosophy of Science, 15(1):51–63, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590020029305.

Engler:2002:EMB

[Eng02] Gideon Engler. Einstein and the most beautiful theories in physics. International Studies in the Philosophy of Science, 16(1):27–37, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120118800.

Engler:2005:EHT

[Eng05] Gideon Engler. Einstein, his theories, and his aesthetic considerations. International Studies in the Philosophy of Science, 19(1):21–30, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500051068.

Esfeld:2007:TRM

[ES07] Michael Esfeld and Christian Sachse.

Theory reduction by means of functional sub-types. International Studies in the Philosophy of Science, 21(1):1–17, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701305735.

Eser:2009:WB

[Ese09] Uta Eser. What is biodiversity? International Studies in the Philosophy of Science, 23(3):330–334, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903197823.

Esfeld:2009:MNS

[Esf09] Michael Esfeld. The modal nature of structures in ontic structural realism. International Studies in the Philosophy of Science, 23(2):179–194, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903006917.

Earman:2014:RCA

[EV14] John Earman and Giovanni Valente. Relativistic causality in algebraic quantum field theory. International Studies in the Philosophy of Science, 28(1):1–48, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915652.

Farrell:2000:WPF

[Far00] Robert P. Farrell. Will the Popperian Feyerabend please step forward: Pluralistic, Popperian themes in the philosophy of Paul Feyerabend. International Studies in the Philosophy of Science, 14(3):257–266, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437764.

Fard:2019:PNP

[Far19] Mohammad Reza Haghighi Fard. In praise of natural philosophy: A revolution for thought and life. International Studies in the Philosophy of Science, 32(1):67–69, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1615663.

Fasce:2018:PLS

[Fas18] Angelo Fasce. Are pseudosciences like seagulls? A discriminant metacriterion facilitates the solution of the demarcation problem. International Studies in the Philosophy of Science, 32(3–4):155–175, 2018. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1767891.

Fuller:2010:LE

[FD10] Steve Fuller and Willem B. Drees.

Letter to the Editor. International Studies in the Philosophy of Science, 24(2):217–221, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.484546.

Feher:1986:MAS

[Feh87] Marta Fehér. The method of analysissynthesis and the structure of causal explanation in Newton. International Studies in the Philosophy of Science, 1(1):60–84, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573282.

Feher:1987:RMA

[Feh88] Maria Feher. The role of metaphor and analogy in the birth of the principle of least action of Maupertuis (1698–1759). International Studies in the Philosophy of Science, 2(2):175–188, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573313.

Feher:1990:RAP

[Feh90] Mart Fehér. The role accorded to the public by philosophers of science. International Studies in the Philosophy of Science, 4(3):229–240, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573363.

Feher:1998:BAA

[Feh98a] Márta Fehér. Bad arguments against a good case (Laudan’s attack on the strong programme). International Studies in the Philosophy of Science, 12(3):233–238, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573598.

Feher:1998:PAG

[Feh98b] Marta Fehér. Patterns of argumentation in Galileo’s Discorsi. International Studies in the Philosophy of Science, 12(1):17–24, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573579.

Felline:2014:QIT

[Fel14] Laura Felline. Quantum information theory and the foundations of quantum mechanics. International Studies in the Philosophy of Science, 28(3):349– 352, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2014.
                                        953348.

Fernflores:2011:BSP

[Fer11] Francisco Fernflores. Bell’s spaceships problem and the foundations of special relativity. International Studies in the Philosophy of Science, 25(4):351– 370, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2011.
                                        623364.

Feyerabend:2022:SMN

[Fey22] Paul K. Feyerabend. Single magnetic northpoles and southpoles and their importance for science. International Studies in the Philosophy of Science, 35(1):95–117, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.

2158430. Ten lectures delivered at the University of Vienna during the summer semester of 1947.

Firt:2021:HDO

[FHS21] Erez Firt, Meir Hemmo, and Orly Shenker. Hempel’s Dilemma: Not only for physicalism*. International Studies in the Philosophy of Science, 34(2):101–129, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2041969.

Filomeno:2016:FEO

[Fil16] Aldo Filomeno. Fundamentality, effectiveness, and objectivity of gauge symmetries. International Studies in the Philosophy of Science, 30(1):19– 37, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2016.
1240459.

Fjelland:1991:TLO

[Fje91] Ragnar Fjelland. The theoryladenness of observations, the role of scientific instruments, and the Kantian a priori. International Studies in the Philosophy of Science, 5(3):269–280, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573399.

French:1999:RSA

[FL99] Steven French and James Ladyman. Reinflating the semantic approach. International Studies in the Philosophy of Science, 13(2):103–121, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573612.

Flannery:2003:AAF

[Fla03] Maura C. Flannery. Agnes Arber: form in the mind and the eye. International Studies in the Philosophy of Science, 17(3):281–300, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859032000169479.

Flores:1999:ETT

[Flo99] Francisco Flores. Einstein’s theory of theories and types of theoretical explanation. International Studies in the Philosophy of Science, 13(2):123–134, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573613.

Flores:2005:IEE

[Flo05] Francisco Flores. Interpretations of Einstein’s equation E = mc^{2}. International Studies in the Philosophy of Science, 19(3):245–260, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500462257. See re-

ply [Kra06] and response [Flo06].

Flores:2006:IER

[Flo06] Francisco Flores. On the interpretation of the equation E = mc^{2}: response to W. Krajewski: “On the interpretation of the equation E = mc^{2}: reply to F. Flores” [Int. Stud. Philos. Sci. 20 (2006), no. 2, 215–216; MR2253243]. International Studies in the Philosophy of Science, 20(2):217–218, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814480. See

[Kra06].

Foley:1986:ERS

[Fol87] Richard Foley. Epistemic rationality and scientific rationality. International Studies in the Philosophy of Science, 1(2):233–250, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573294.

Forrai:1993:MPR

[For93] Gábor Forrai. From the method of proofs and refutations to the methodology of scientific research programmes. International Studies in the Philosophy of Science, 7(2):161–175, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698599308573461.

Forge:1996:EQS

[For96] John Forge. Explanation and the quantum state. International Studies in the Philosophy of Science, 10(3):203–215, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573540.

Forge:2000:QQM

[For00] John Forge. Quantities in quantum mechanics. International Studies in the Philosophy of Science, 14(1):43– 56, 2000. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        /doi/abs/10.1080/026985900111891.

Franssen:2009:CMT

[Fra09] Maarten Franssen. Creations of the mind: Theories of artifacts and their representation. International Studies in the Philosophy of Science, 23(3):334–337, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903197856.

Freedman:1997:WNNa

[Fre97a] Karyn Freedman. What’s new on the net. International Studies in the Philosophy of Science, 11(2):205–206, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573564.

Freedman:1997:WNNb

[Fre97b] Karyn Freedman. What’s new on the net. International Studies in the Philosophy of Science, 11(3):303–304, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573573.

Freedman:1998:WNNa

[Fre98a] Karyn Freedman. What’s new on the net. International Studies in the Philosophy of Science, 12(1):87–89, 1998. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573583.

Freedman:1998:WNNb

[Fre98b] Karyn Freedman. What’s new on the net. International Studies in the Philosophy of Science, 12(2):193–195, 1998. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573592.

Freedman:1999:WNN

[Fre99] Karyn Freedman. What’s new on the net. International Studies in the Philosophy of Science, 13(2):193–194, 1999. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573618.

Freedman:2006:NNE

[Fre06] Karyn L. Freedman. Normative naturalism and epistemic Relativism. International Studies in the Philosophy of Science, 20(3):309–322, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600960952.

Frigg:2003:PSR

[Fri03] Roman Frigg. On the property structure of realist collapse interpretations of quantum mechanics and the socalled “counting anomaly”. International Studies in the Philosophy of Science, 17(1):43–57, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590305232.

Friederich:2017:FTO

[Fri17] Simon Friederich. Fine-tuning as old evidence, double counting, and the multiverse. International Studies in the Philosophy of Science, 31(4):363– 377, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.
1565214.

Fuchs:1993:AET

[Fuc93] Stephan Fuchs. Against essentialism in theories of rational action: a reply to Raymond Boudon. International Studies in the Philosophy of Science, 7(1):37–39, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573444. See

[Bou93b] and reply [Bou93a].

Fuller:2022:DFK

[Ful22] Gareth Fuller. A defence of functional kinds: Multiple realisability and explanatory counterfactuals. International Studies in the Philosophy of Science, 35(2):119–133, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2144436.

Futch:2002:LNT

[Fut02] Michael J. Futch. Leibniz’s nontensed theory of time. International Studies in the Philosophy of Science, 16(2):125–139, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145070.

Gao:2015:AEE

[Gao15] Jie Gao. Advances in experimental epistemology. International Studies in the Philosophy of Science, 29(1):101– 105, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1071554.

Gao:2017:USP

[Gao17] Shan Gao. Understanding scientific progress: Aim-oriented empiricism. International Studies in the Philosophy of Science, 31(4):435–438, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565211.

Garnham:1994:MM

[Gar94] Alan Garnham. March of the models.

International Studies in the Philosophy of Science, 8(1):37–39, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573477.

Gattei:2000:OJW

[Gat00] Stefano Gattei. Obituary: John W. N. Watkins, 1924–1999. International Studies in the Philosophy of Science, 14(1):77–81, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/026985900111918.

Gauthier:2005:HWM

[Gau05] Yvon Gauthier. Hermann Weyl on Minkowskian space–time and Riemannian geometry. International Studies in the Philosophy of Science, 19(3):261–269, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500462265.

Gelfert:2003:MSU

[Gel03] Axel Gelfert. Manipulative success and the unreal. International Studies in the Philosophy of Science, 17(3):245–263, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859032000169451.

Ghassib:2012:TKI

[Gha12] Hisham Ghassib. A theory of the knowledge industry. International Studies in the Philosophy of Science, 26(4):447–456, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.748499.

Ghassib:2017:RME

[Gha17] Hisham Ghassib. Reason and method in Einstein’s relativity. International Studies in the Philosophy of Science, 31(4):331–342, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565204.

Ghins:2003:TKE

[Ghi03] Michel Ghins. Thomas Kuhn on the existence of the world. International Studies in the Philosophy of Science, 17(3):265–279, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859032000169460.

Gijsbers:2009:DAS

[Gij09] Victor Gijsbers. Depth: An account of scientific explanation. International Studies in the Philosophy of Science, 23(2):225–228, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903007212.

Gillett:1991:LSE

[Gil91] Grant Gillett. Language, social ecology and experience. International Studies in the Philosophy of Science, 5(3):195–203, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573393.

Gillies:1992:CSD

[Gil92] Donald A. Gillies. Comments on ‘Scientific Discovery as Problem Solving’ by Herbert A. Simon. International Studies in the Philosophy of Science, 6(1):29–31, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573407. See

[Sim92a, Sim92b].

Ginev:2005:APP

[Gin05] Dimitri Ginev. Against the politics of postmodern philosophy of science. International Studies in the Philosophy of Science, 19(2):191–208, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500249522.

Ginev:2012:TAH

[Gin12] Dimitri Ginev. Two accounts of the hermeneutic fore-structure of scientific research. International Studies in the Philosophy of Science, 26(4):423–445, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
748498.

Girotto:1994:MTI

[Gir94] Vittorio Girotto. Is the model theory of induction also a theory of inductive reasoning? International Studies in the Philosophy of Science, 8(1):41–43, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573478.

Godman:2014:SEN

[God14] Marion Godman. Scientific enquiry and natural kinds: From planets to mallards. International Studies in the Philosophy of Science, 28(3):343–346, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
953346.

Goldstick:1993:LNP

[Gol93] D. Goldstick. Laws of nature and physical existents. International Studies in the Philosophy of Science, 7(3):255–265, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573469.

Golosz:1999:FAS

[Goł99] Jerzy Gołosz. On Field’s argument for substantivalism. International Studies in the Philosophy of Science, 13(1):5–16, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573604.

Goldstick:2020:FJP

[Gol20] Daniel Goldstick. De Facto and De Jure in the practice of induction. International Studies in the Philosophy of Science, 33(2):105–116, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1866478.

Gombay:1991:PDL

[Gom91] A. Gombay. Psychic determinism: Leibniz and Freud. International Studies in the Philosophy of Science, 5(3):205–213, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573394.

Gonzalez:1995:RCP

[Gon95] Wenceslao J. González. Reichenbach’s concept of prediction. International Studies in the Philosophy of Science, 9(1):37–58, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573505.

Gontier:2012:SAE

[Gon12] Nathalie Gontier. Selectionist approaches in evolutionary linguistics: An epistemological analysis. International Studies in the Philosophy of Science, 26(1):67–95, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.653114.

Gonzalez:2015:CEP

[Gon15] Wenceslao J. Gonzalez. From the characterization of ‘European philosophy of science’ to the case of philosophy of the social sciences. International Studies in the Philosophy of Science, 29(2):167–188, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1119418.

Gooding:1990:TOE

[Goo90] David Gooding. Theory and observation: The experimental nexus. International Studies in the Philosophy of Science, 4(2):131–148, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573353.

Gottlob:2000:NAP

[Got00] Rainer Gottlob. New aspects of the probabilistic evaluation of hypotheses and experience. International Studies in the Philosophy of Science, 14(2):147–163, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590050045737.

Gouk:1986:NMM

[Gou87] Penelope Gouk. Newton and music: From the microcosm to the macrocosm. International Studies in the Philosophy of Science, 1(1):36–59, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573281.

Grammenos:2018:MEP

[Gra18] Theophanes Grammenos. Metaphysical experiments — physics and the invention of the universe. International Studies in the Philosophy of Science, 32(3–4):232–234, 2018. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1767893.

Green:1994:IRS

[Gre94] David W. Green. Induction: Representation, strategy and argument. International Studies in the Philosophy of Science, 8(1):45–50, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573479.

Grobler:1991:VFM

[Gro91] Adam Grobler. Van Fraassen’s metaphysical move. International Studies in the Philosophy of Science, 5(1):21–34, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573375.

Gurova:2022:UTT

[Gur22] Lilia Gurova. The uses of truth: Is there room for reconciliation of factivist and non-factivist accounts of scientific understanding? International Studies in the Philosophy of Science, 35(3–4):211–221, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2022.2084329.

Grune-Yanoff:2011:ICM

[GY11] Till Grüne-Yanoff. Isolation is not characteristic of models. International Studies in the Philosophy of Science, 25(2):119–137, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.574854.

Hacker:1995:HTP

[Hac95] P. M. S. Hacker. Helmholtz’s theory of perception: An investigation into its conceptual framework. International Studies in the Philosophy of Science, 9(3):199–214, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573519.

Hansson:2009:CGK

[Han09] Sven Ove Hansson. Cutting the Gordian knot of demarcation. International Studies in the Philosophy of Science, 23(3):237–243, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903196007.

Hansson:2020:DDD

[Han20] Sven Ove Hansson. Disciplines, doctrines, and deviant science. International Studies in the Philosophy of Science, 33(1):43–52, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1831258.

Harre:1987:WMA

[Har88] R. Harre. Where models and analogies really count. International Studies in the Philosophy of Science, 2 (2):118–133, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573310.

Harre:1990:TAS

[Har90] R. Harré. Tracks and affordances:

The sources of a physical ontology. International Studies in the Philosophy of Science, 4(2):149–158, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573354.

Harre:1991:SRI

[Har91] Rom Harré. Some reflections on the individuation of events. International Studies in the Philosophy of Science, 5(1):49–63, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573377.

Harre:1995:ROP

[Har95] Rom Harré. Realism and an ontology of powerful particulars. International Studies in the Philosophy of Science, 9(3):285–300, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573524.

Harre:1997:FJR

[Har97a] Rom Harré, editor. Festschrift for J. R. Lucas. Carfax Publishing, Taylor & Francis Ltd., Basingstoke, UK, 1997. ISSN 0269-8595 (print), 1469-9281 (electronic). 1–108 pp. Internat. Stud. Philos. Sci. 11 (1997), no. 1.

Harre:1997:I

[Har97b] Rom Harré. Introduction. International Studies in the Philosophy of Science, 11(1):5, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573546.

Harre:1997:ODS

[Har97c] Rom Harré. The ontological duality of space–time variables. International Studies in the Philosophy of Science, 11(1):83–96, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573552. Festschrift

for J. R. Lucas.

Harbecke:2010:MCN

[Har10] Jens Harbecke. Mechanistic constitution in neurobiological explanations. International Studies in the Philosophy of Science, 24(3):267–285, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522409.

Hawley:2010:SO

[Haw10] Katherine Hawley. The structure of objects. International Studies in the Philosophy of Science, 24(3):336–339, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
522418.

Heidelberger:2001:OLT

[Hei01] Michael Heidelberger. Origins of the logical theory of probability: Von Kries, Wittgenstein, Waismann. International Studies in the Philosophy of Science, 15(2):177–188, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120059040.

Heidelberger:2006:AMF

[Hei06] Michael Heidelberger. Applying models in fluid dynamics. International Studies in the Philosophy of Science, 20(1):49–67, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600641016.

Hesse:1992:CHS

[Hes92] Mary Hesse. Comment on Herbert Simon, “Scientific discovery as problem solving”. International Studies in the Philosophy of Science, 6(1):33–34, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573408. See

[Sim92a].

Harre:1992:BR

[HF92] Rom Harré and Marta Fehér. Book reviews. International Studies in the Philosophy of Science, 6(3):249–253, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        /doi/abs/10.1080/02698599208573435.

Hibbert:2016:WIS

[Hib16] Ruth Hibbert. What is an immature science? International Studies in the Philosophy of Science, 30(1):1– 17, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2016.
                                        1240433.

Hoffmann-Kolss:2014:IHL

[HK14] Vera Hoffmann-Kolss. Interventionism and higher-level causation. International Studies in the Philosophy of Science, 28(1):49–64, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.915653.

Henderson:1999:BR

[HMO^{+}99] P. Henderson, Matthew MacKenzie, Dawn Ogden, Jutta Rockmann, Spas Spassov, and John Taylor. Book reviews. International Studies in the Philosophy of Science, 13(1):83–96, 1999. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573609.

Hochstein:2013:IME

[Hoc13] Eric Hochstein. Intentional models as essential scientific tools. International Studies in the Philosophy of Science, 27(2):199–217, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.813251.

Hodgson:1997:RQM

[Hod97] Peter Hodgson. Realism and quantum mechanics. International Studies in the Philosophy of Science, 11(1):53–65, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573550. Festschrift

for J. R. Lucas.

Hoefer:2000:KHE

[Hoe00] Carl Hoefer. Kant’s hands and Earman’s pions: Chirality arguments for substantival space. International Studies in the Philosophy of Science, 14(3):237–256, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437755.

Hohwy:2003:CEP

[Hoh03] Jakob Hohwy. Capacities, explanation and the possibility of disunity. International Studies in the Philosophy of Science, 17(2):179–190, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859031000160630.

Holst:2009:WPS

[Hol09] Cathrine Holst. What is philosophy of social science? International Studies in the Philosophy of Science, 23(3):313–321, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903197781.

Hon:1991:CME

[Hon91] Giora Hon. Can the monster errour be slain? International Studies in the Philosophy of Science, 5(3):257–268, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573398.

Hope:2013:STD

[Hop13] Kristin Lofthus Hope. Science transformed? Debating claims of an epochal break. International Studies in the Philosophy of Science, 27(2):228–231, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2013.
                                        813252.

Howson:2009:SNT

[How09] Colin Howson. Sorites is no threat to Modus Ponens: a reply to Kochan. International Studies in the Philosophy of Science, 23(2):209–212, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903007188. See

[Koc08, Koc09].

Howson:2011:NAH

[How11] Colin Howson. No answer to Hume [discussion of MR2680848]. International Studies in the Philosophy of Science, 25(3):279–284, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.605249. See

replies [Ste11, BTB16].

Howson:2016:HPH

[How16] Colin Howson. How pseudohypotheses defeat a non-Bayesian theory of evidence: Reply to Bandyopadhyay, Taper, and Brittan. International Studies in the Philosophy of Science, 30(3):299–306, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1316111.

Howson:2018:PCS

[How18] Colin Howson. The primacy of the classical? Saul Kripke meets Niels Bohr. International Studies in the Philosophy of Science, 32(3–4):141–153, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/

doi/full/10.1080/02698595.2020.
1767890.

Haufe:2009:WNM

[HS09] Chris Haufe and Matthew H. Slater. Where no mind has gone before: Exploring laws in distant and lonely worlds. International Studies in the Philosophy of Science, 23(3):265–276, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590903196049.

Hemmo:2015:LE

[HS15] Meir Hemmo and Orly Shenker. Letter to the Editor. International Studies in the Philosophy of Science, 29(1): 91–93, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1079075.

Huttemanna:2000:EP

[HT00] Andreas Hüttemanna and Orestis Terzidis. Emergence in physics. International Studies in the Philosophy of Science, 14(3):267–281, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437773.

Huggett:1999:WMS

[Hug99] Nick Huggett. Why manifold substantivalism is probably not a consequence of classical mechanics. International Studies in the Philosophy of Science, 13(1):17–34, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573605.

Hulton:2014:NCH

[Hul14] Ana Hulton. Natural categories and human kinds: Classification in the natural and social sciences. International Studies in the Philosophy of Science, 28(1):102–105, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915654.

Hunt:1990:LPU

[Hun90] G. M. K. Hunt. Laws: Projectability and uniformity. International Studies in the Philosophy of Science, 4(3):241–246, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573364.

Hunt:1994:CTI

[Hun94] Earl Hunt. Comments on a theory of induction. International Studies in the Philosophy of Science, 8(1):51–55, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573480.

Hung:2001:KPR

[Hun01] Edwin H.-C. Hung. Kuhnian paradigms as representational spaces: New perspectives on the problems of incommensurability, scientific explanation, and physical necessity. International Studies in the Philosophy of Science, 15(3):275–292, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120080217.

Hunt:2011:TSI

[Hun11] Shelby D. Hunt. Theory status, inductive realism, and approximate truth: No miracles, no charades. International Studies in the Philosophy of Science, 25(2):159–178, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.574858.

Iliescu:1991:BR

[Ili91] Adrian-Paul Iliescu. Book review. International Studies in the Philosophy of Science, 5(1):89–92, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573380.

Illari:2011:MED

[Ill11] Phyllis McKay Illari. Mechanistic evidence: Disambiguating the Russo–Williamson thesis. International Studies in the Philosophy of Science, 25(2):139–157, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.574856.

Inan:2014:RDT

[Ina14] Ilhan Inan. Rigid designation and theoretical identities. International Studies in the Philosophy of Science, 28(2):217–220, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.932531.

Irvine:1993:HBP

[Irv93] A. D. Irvine. How Braess’ paradox solves Newcomb’s problem. International Studies in the Philosophy of Science, 7(2):141–160, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573460. See re-

ply [Mar96].

Irvine:1997:PSW

[Irv97] A. D. Irvine. Philosophy of science in Western Canada. International Studies in the Philosophy of Science, 11(3):289–301, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573572.

IslasMondragon:2017:IRA

[Isl17] Damián Islas Mondragón. The inference rule of addition and the semantic view of scientific progress: Reply to Mizrahi. International Studies in the Philosophy of Science, 31(4):421– 425, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.

1565212. See [Miz17].

Ivanova:2011:FRP

[Iva11] Milena Ivanova. Friedman’s relativised A Priori and structural realism: In search of compatibility. International Studies in the Philosophy of Science, 25(1):23–37, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.552416.

Ivanova:2014:ESS

[Iva14] Milena Ivanova. Explaining science’s success: Understanding how scientific knowledge works. International Studies in the Philosophy of Science, 28(1):105–108, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.915655.

Jacquette:1993:PTP

[Jac93] Dale Jacquette. Pollock on token physicalism, agent materialism and strong artificial intelligence. International Studies in the Philosophy of Science, 7(2):127–140, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573459.

Jacobs:2001:LPS

[Jac01] Struan Jacobs. Limits to problem solving in science. International Studies in the Philosophy of Science, 15(3):231–242, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120080172.

Jacob:2011:CAR

[Jac11] Merle Jacob. The commodification of academic research: Science and the modern university. International Studies in the Philosophy of Science, 25(4):423–426, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.623373.

Jaeger:2002:HSB

[Jae02] Lydia Jaeger. Humean supervenience and best-system laws. International Studies in the Philosophy of Science, 16(2):141–155, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145089.

Jaeger:2017:ROA

[Jae17] Lydia Jaeger. Rethinking order: After the laws of nature. International Studies in the Philosophy of Science, 31(2):215–218, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1424765.

Janich:1986:NFP

[Jan87] Peter Janich. The normative foundation of physics. International Studies in the Philosophy of Science, 1 (2):251–261, 1986–1987. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573295.

Janich:1988:DBN

[Jan89] Peter Janich. Does biology need a relativistic revision? International Studies in the Philosophy of Science, 3(2):190–198, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573335.

Jansson:2015:NST

[Jan15] Lina Jansson. The nature of scientific thinking: On interpretation, explanation, and understanding. International Studies in the Philosophy of Science, 29(2):218–221, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1119425.

Jimenez-Buedo:2015:LDG

[JB15] María Jiménez-Buedo. The last dictator game? Dominance, reactivity, and the methodological artefact in experimental economics. International Studies in the Philosophy of Science, 29(3):295–310, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1179042.

Jones:1995:BR

[JBM^{+}95] W. Jones, James Robert Brown, W. J. Mander, Władysław Krajewski, John M. Preston, Stathis Psillos, Katherine Hawley, and John Taylor. Book reviews. International Studies in the Philosophy of Science, 9(2):157–188, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573515.

Jebeile:2017:CSE

[Jeb17] Julie Jebeile. Computer simulation, experiment, and novelty. International Studies in the Philosophy of Science, 31(4):379–395, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565205.

Jebeile:2015:EMR

[JK15] Julie Jebeile and Ashley Graham Kennedy. Explaining with models: The role of idealizations. International Studies in the Philosophy of Science, 29(4):383–392, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1195143.

Johnson-Laird:1994:MTI

[JL94a] Philip N. Johnson-Laird. A model theory of induction. International Studies in the Philosophy of Science, 8(1):5–29, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573474.

Johnson-Laird:1994:RCM

[JL94b] Philip N. Johnson-Laird. Reply to the commentators on a model theory of induction. International Studies in the Philosophy of Science, 8(1):73–96, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573485.

Johnson-Laird:1992:SHS

[JLL92] Philip Johnson-Laird and Paolo Legrenzi. Science as heuristic search: Some comments on Simon’s theory. International Studies in the Philosophy of Science, 6(1):35–39, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573409.

Johansen:2016:CSP

[JM16] Mikkel Willum Johansen and Morten Misfeldt. Computers as a source of a posteriori knowledge in mathematics. International Studies in the Philosophy of Science, 30(2):111–127, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2016.1265862.

Jones:2000:UET

[Jon00] Ward E. Jones. Underdetermination and the explanation of theory-acceptance: a response to Samir Okasha. International Studies in the Philosophy of Science, 14(3):299–304, 2000. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/026985900437791.  See

[Oka00b].

Joyce:2013:CDH

[Joy13] Kelly Joyce. Chagas disease: History of a continent’s scourge. International Studies in the Philosophy of Science, 27(4):459–461, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.868180.

Jones:1994:BR

[JSJ^{+}94] W. Jones, Helen Steward, Fiona Jenkins, Lee Mcintyre, Mary Hesse, C. W. Kilmister, Richard J. Hall, and Timothy Williamson. Book reviews. International Studies in the Philosophy of Science, 8(2):147–168, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573491.

Jraissati:2012:CCC

[JWDD12] Yasmina Jraissati, Elley Wakui, Lieven Decock, and Igor Douven. Constraints on colour category formation. International Studies in the Philosophy of Science, 26(2):171–196, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.703479.

Kaiser:2015:PM

[Kai15] Marie I. Kaiser. Philosophy of microbiology. International Studies in the Philosophy of Science, 29(2):224– 228, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1119427.

Kapusta:2010:PST

[Kap10] Paweł Kapusta. Philosophy of science for theologians: An introduction. International Studies in the Philosophy of Science, 24(4):443–446, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.543356.

Kardaun:2015:JQS

[Kar15] Maria Kardaun. Jung and the question of science. International Studies in the Philosophy of Science, 29(1):108– 111, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1071556.

Kastner:2004:SBN

[Kas04] Ruth E. Kastner. Shutters, boxes, but no paradoxes: Time symmetry puzzles in quantum theory. International Studies in the Philosophy of Science, 18(1):89–94, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590412331289279.

Kastner:2015:ECD

[Käs15] Lena Kästner. Epistemic cognition and development: The psychology of justification and truth. International Studies in the Philosophy of Science, 29(4):444–447, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1195150.

Kastner:2019:NMP

[Käs19] Lena Kästner. The new mechanical philosophy. International Studies in the Philosophy of Science, 32(1):69– 72, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2019.
                                        1615664.

Katzav:2013:DCP

[Kat13] Joel Katzav. Dispositions, causes, persistence as is, and general relativity. International Studies in the Philosophy of Science, 27(1):41–57, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.783974.

Kazumi:2014:DCC

[Kaz14] Inoue Kazumi. Dialectical contradictions and classical formal logic. International Studies in the Philosophy of Science, 28(2):113–132, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.932526.

Kleinhans:2005:TIE

[KBdR05] Maarten G. Kleinhans, Chris J. J.

Buskes, and Henk W. de Regt. Terra Incognita: explanation and reduction in Earth science. International Studies in the Philosophy of Science, 19(3):289–317, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500462356.

Keet:2012:AIB

[Kee12] C. Maria Keet. Artificial intelligence:

The basics. International Studies in the Philosophy of Science, 26(3):351– 354, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
731737.

Kendig:2011:DPB

[Ken11] Catherine Kendig. Debates in philosophy of biology: One long argument, or many? International Studies in the Philosophy of Science, 25(1): 73–81, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
552419.

Kennedy:2013:MSR

[Ken13] Ashley Graham Kennedy. Mathematics and scientific representation. International Studies in the Philosophy of Science, 27(1):95–98, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.783975.

Kent:2022:PFD

[Ken22] Rory Kent. Paul Feyerabend and the dialectical character of quantum mechanics: a lesson in philosophical dadaism. International Studies in the Philosophy of Science, 35(1):51– 67, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2022.
                                        2075687.

Kuby:2022:FQT

[KF22] Daniel Kuby and Patrick Fraser. Feyerabend on the quantum theory of measurement: a reassessment. International Studies in the Philosophy of Science, 35(1):23–49, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2022.2067811.

Khalidi:2005:AFR

[Kha05] Muhammad Ali Khalidi. Against functional reductionism in cognitive science. International Studies in the Philosophy of Science, 19(3):319–333, 2005. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590500462448.

Khalifa:2010:DPB

[Kha10] Kareem Khalifa. Default privilege and bad lots: Underconsideration and explanatory inference. International Studies in the Philosophy of Science, 24(1):91–105, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903467135.

Kidd:2013:HCI

[Kid13] Ian James Kidd. Historical contingency and the impact of scientific imperialism. International Studies in the Philosophy of Science, 27(3):315–324, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2013.
825494.

Kidd:2016:WFP

[Kid16] Ian James Kidd. Was Feyerabend a postmodernist? International Studies in the Philosophy of Science, 30(1): 55–68, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2016.
1240463.

Kindi:2011:CSR

[Kin11] Vasso Kindi. The challenge of scientific revolutions: Van Fraassen’s and Friedman’s responses. International Studies in the Philosophy of Science, 25(4):327–349, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.623363.

Kindi:2017:RRB

[Kin17] Vasso Kindi. A reconsideration of the relation between Kuhnian incommensurability and translation. International Studies in the Philosophy of Science, 31(4):397–414, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2019.1565206.

Kindi:2018:ER

[Kin18] Vasso Kindi. Editorial report 2019. International Studies in the Philosophy of Science, 32(3–4):235–236, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2019.1814040.

Kindi:2019:E

[Kin19] Vasso Kindi. Editorial. International Studies in the Philosophy of Science, 32(1):1, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2019.1650560.

Kindi:2020:ER

[Kin20] Vasso Kindi. Editorial report 2020. International Studies in the Philosophy of Science, 33(4):259–260, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2020.1986773.

Kistler:1999:MRR

[Kis99] Max Kistler. Multiple realization, reduction and mental properties. International Studies in the Philosophy of Science, 13(2):135–149, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573614.

Klein:2001:CRH

[Kle01] Carsten Klein. Conventionalism and realism in Hans Reichenbach’s philosophy of geometry. International Studies in the Philosophy of Science, 15(3):243–251, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120080181.

Kliemt:1993:PWR

[Kli93] Hartmut Kliemt. Perfect and workable rationality: a comment on Raymond Boudon’s paper. International Studies in the Philosophy of Science, 7(1):41–43, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573445. See

[Bou93b] and reply [Bou93a].

Knowles:2002:WRW

[Kno02] Jonathan Knowles. What’s really wrong with Laudan’s normative naturalism. International Studies in the Philosophy of Science, 16(2):171–186, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590220145106.

Knox:2011:BRA

[Kno11] Eleanor Knox. Book review: The Ashgate Companion to Contemporary Philosophy of Physics. International Studies in the Philosophy of Science, 25(2):199–202, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.574868.

Knuuttila:2010:SMP

[Knu10] Tarja Knuuttila. Scientific models in philosophy of science. International Studies in the Philosophy of Science, 24(4):437–440, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.543353.

Kochan:2008:RRS

[Koc08] Jeff Kochan. Realism, reliabilism, and the ‘strong programme’ in the sociology of scientific knowledge. International Studies in the Philosophy of Science, 22(1):21–38, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802280878. See re-

ply [How09] and response [Koc09].

Kochan:2009:EMR

[Koc09] Jeff Kochan. The exception makes the rule: Reply to Howson. International Studies in the Philosophy of Science, 23(2):213–216, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903007196. See

[Koc08, How09].

Kochan:2012:SSN

[Koc12] Jeff Kochan. Science studies as naturalized philosophy. International Studies in the Philosophy of Science, 26(1):121–124, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.653115.

Kosso:2000:FAS

[Kos00] Peter Kosso. Fundamental and accidental symmetries. International Studies in the Philosophy of Science, 14(2):109–121, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590050045700.

Kosso:2002:OBS

[Kos02] Peter Kosso. The omniscienter:

Beauty and scientific understanding. International Studies in the Philosophy of Science, 16(1):39–48, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120118819.

Kouw:2015:MBR

[Kou15] Matthijs Kouw. Model-based reasoning in science and technology: Theoretical and cognitive issues. International Studies in the Philosophy of Science, 29(1):105–108, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1071555.

Kowalenko:2014:CPL

[Kow14] Robert Kowalenko. Ceteris Paribus laws: a naturalistic account. International Studies in the Philosophy of Science, 28(2):133–155, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.932527.

Kowalenko:2017:MCP

[Kow17] Robert Kowalenko. Manipulationism, Ceteris Paribus laws, and the bugbear of background knowledge. International Studies in the Philosophy of Science, 31(3):261–283, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2018.1463693.

Krajewski:1992:QOK

[Kra92] Władysław Krajewski. Questions of the objects of knowledge and types of realism. International Studies in the Philosophy of Science, 6(3):205–213, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573431.

Krajewski:1997:IOM

[Kra97] Władysław Krajewski. Ideal objects as models in science. International Studies in the Philosophy of Science, 11(2):185–190, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573562.

Krajewski:2006:IER

[Kra06] Władysław Krajewski. On the interpretation of the equation E = mc^{2}: reply to F. Flores: “Interpretations of Einstein’s equation E = mc^{2}” [Int. Stud. Philos. Sci. 19 (2005), no. 3, 245–260; MR2201925]. International Studies in the Philosophy of Science, 20(2):215–216, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814449. See

[Flo05, Flo06].

Krauss:2021:AOV

[Kra21] Alexander Krauss. Assessing the overall validity of randomised controlled trials. International Studies in the Philosophy of Science, 34(3):159– 182, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2021.
2002676.

Kremakova:2017:AAC

[Kre17] Milena Kremakova. Accelerating academia: The changing structure of academic time. International Studies in the Philosophy of Science, 31(2):221–224, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1424766.

Krohs:2008:HDC

[Kro08] Ulrich Krohs. How digital computer simulations explain real-world processes. International Studies in the Philosophy of Science, 22(3):277–292, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590802567324.

Krause:2007:BR

[KSA07] Décio Krause, Eric Schliesser, and Hanne Andersen. Book reviews. International Studies in the Philosophy of Science, 21(3):345–357, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701589619.

Kujundzic:1998:RVT

[Kuj98] Nebojsa Kujundzic. The role of variation in thought experiments. International Studies in the Philosophy of Science, 12(3):239–243, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573599.

Kukla:1993:EB

[Kuk93] Andre Kukla. Epistemic boundedness. International Studies in the Philosophy of Science, 7(2):121–126, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573458.

Kukla:1995:TLI

[Kuk95] André Kukla. Is there a logic of incoherence? International Studies in the Philosophy of Science, 9(1):59–71, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573506.

Kuorikoski:2009:TCM

[Kuo09] Jaakko Kuorikoski. Two concepts of mechanism: Componential causal system and abstract form of interaction. International Studies in the Philosophy of Science, 23(2):143–160, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903006875.

Kutschmann:1986:SIS

[Kut87] Werner Kutschmann. Scientific instruments and the senses: Towards an anthropological historiography of the natural sciences. International Studies in the Philosophy of Science, 1(1):106–123, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573284.

Kutach:2010:CBM

[Kut10] Douglas Kutach. A connection between Minkowski and Galilean space– times in quantum mechanics. International Studies in the Philosophy of Science, 24(1):15–29, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903467093.

Kutach:2011:BI

[Kut11] Douglas Kutach. Backtracking influence. International Studies in the Philosophy of Science, 25(1):55–71, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.552418.

Kuukkanen:2009:RK

[Kuu09] Jouni-Matti Kuukkanen. Rereading Kuhn. International Studies in the Philosophy of Science, 23(2):217–224, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590903007204.

Kuukkanen:2012:AOS

[Kuu12] Jouni-Matti Kuukkanen. Autonomy and objectivity of science. International Studies in the Philosophy of Science, 26(3):309–334, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.731733.

Kuukkanen:2017:LRR

[Kuu17] Jouni-Matti Kuukkanen. Lakatosian rational reconstruction updated. International Studies in the Philosophy of Science, 31(1):83–102, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1370930.

Kvasz:2016:RMT

[Kva16] Ladislav Kvasz. Revisiting the mathematisation thesis: Galileo, Descartes, Newton, and the language of nature. International Studies in the Philosophy of Science, 30(4):399–406, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1331978.

Lam:2012:GP

[Lam12] Vincent Lam. Geometric possibility. International Studies in the Philosophy of Science, 26(2):226–229, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.703483.

Lam:2013:ESQ

[Lam13] Vincent Lam. The entanglement structure of quantum field systems. International Studies in the Philosophy of Science, 27(1):59–72, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.783976.

Langfeldt:2011:HPT

[Lan11] Liv Langfeldt. How professors think:

Inside the curious world of academic judgment. International Studies in the Philosophy of Science, 25(1):92– 95, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
552423.

Landes:2014:TIP

[Lan14] Jürgen Landes. Tychomancy: Inferring probability from causal structure. International Studies in the Philosophy of Science, 28(4):446–448, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.979675.

Larvor:1994:HME

[Lar94] Brendan Larvor. History, methodology and early algebra. International Studies in the Philosophy of Science, 8(2):113–124, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573488.

Laraudogoitia:2001:ICG

[Lar01] Jon Pérez Laraudogoitia. Indeterminism, classical gravitation and non-collision singularities. International Studies in the Philosophy of Science, 15(3):269–274, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120080208.

Laraudogoitia:2006:GIC

[Lar06] Jon Pérez Laraudogoitia. Global interaction in classical mechanics. International Studies in the Philosophy of Science, 20(2):173–183, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600814357.

Laudisa:2015:LD

[Lau15] Federico Laudisa. Laws are not descriptions. International Studies in the Philosophy of Science, 29(3):251– 270, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1179038.

Lawton:1993:ASC

[Law93] Philip Lawton. Art, science, and the clear blue sky. International Studies in the Philosophy of Science, 7(2):107–119, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573457.

Lazarovici:2017:MWF

[Laz17] Dustin Lazarovici. The meaning of the wave function: In search of the ontology of quantum mechanics. International Studies in the Philosophy of Science, 31(3):321–324, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1463694.

Leng:1998:BR

[LB98] Mary Leng and James Robert Brown. Book reviews. International Studies in the Philosophy of Science, 12(2):197–203, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573593.

Leech:2014:P

[Lee14] Jessica Leech. Properties. International Studies in the Philosophy of Science, 28(4):439–442, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.979673.

Legrenzi:1994:I

[Leg94] Paolo Legrenzi. Introduction. International Studies in the Philosophy of Science, 8(1):3, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573473.

Lelas:1987:RA

[Lel88] Srdjan Lelas. Review article. International Studies in the Philosophy of Science, 2(2):237–242, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573317.

Lelas:1988:ENR

[Lel89] Srdan Lelas. Evolutionary naturalist realism: Can this blend be coherent?

International Studies in the Philosophy of Science, 3(2):136–156, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573332.

Lennox:2001:AUD

[Len01] James G. Lennox. Aristotle on the unity and disunity of science. International Studies in the Philosophy of Science, 15(2):133–144, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120059004.

Leng:2005:PAP

[Len05] Mary Leng. Platonism and anti-Platonism: Why worry? International Studies in the Philosophy of Science, 19(1):65–84, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500051159.

Leonelli:2011:ECT

[Leo11] Sabina Leonelli. An epistemology of the concrete: Twentieth-century histories of life. International Studies in the Philosophy of Science, 25(4):420– 422, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
623372.

Leonelli:2012:CTD

[Leo12] Sabina Leonelli. Classificatory theory in data-intensive science: The case of open biomedical ontologies. International Studies in the Philosophy of Science, 26(1):47–65, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.653119.

Leplin:1987:RET

[Lep88] Jarrett Leplin. The role of experiment in theory construction. International Studies in the Philosophy of Science, 2(1):72–83, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573303.

Leroux:2001:PTF

[Ler01] Jean Leroux. “Picture theories”

as forerunners of the semantic approach to scientific theories. International Studies in the Philosophy of Science, 15(2):189–197, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120059059.

Leslie:1996:DEM

[Les96] John Leslie. A difficulty for Everett’s many-worlds theory. International Studies in the Philosophy of Science, 10(3):239–246, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573542.

Letrud:2019:GKD

[Let19] Kåre Letrud. The Gordian knot of demarcation: Tying up some loose ends. International Studies in the Philosophy of Science, 32(1):3–11, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1618031.

Letrud:2022:ISD

[Let22] Kåre Letrud. Incorrigible science and doctrinal pseudoscience. International Studies in the Philosophy of Science, 35(3–4):269–278, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2218264.

Lewis:2003:FSD

[Lew03] Peter J. Lewis. Four strategies for dealing with the counting anomaly in spontaneous collapse theories of quantum mechanics. International Studies in the Philosophy of Science, 17(2):137–142, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859031000160603.

Lewowicz:2016:SAP

[Lew16] Lucía Lewowicz. Science after the practice turn in the philosophy, history, and social studies of science. International Studies in the Philosophy of Science, 30(2):190–193, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1265863.

Lindenberg:1993:RBF

[Lin93] Siegwart Lindenberg. The rationality of beliefs at the frontier of social science: Comments on Boudon’s paper. International Studies in the Philosophy of Science, 7(1):51–53, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573447. See

[Bou93b] and reply [Bou93a].

Ling:1995:INP

[Lin95] Charles X. Ling. Introducing new predicates to model scientific revolution. International Studies in the Philosophy of Science, 9(1):19–36, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573504.

Lindholm:2022:SPS

[Lin22] Juho Lindholm. Scientific practices as social knowledge. International Studies in the Philosophy of Science, 35(3–4):223–242, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2196930.

Lipton:1990:PP

[Lip90] Peter Lipton. Prediction and prejudice. International Studies in the Philosophy of Science, 4(1):51–65, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573345.

Liu:1997:MTS

[Liu97] Chuang Liu. Models and theories.

I. The semantic view revisited. International Studies in the Philosophy of Science, 11(2):147–164, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573560.

Liu:1998:MTI

[Liu98] Chuang Liu. Models and theories. II. Issues and applications. International Studies in the Philosophy of Science, 12(2):111–128, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573587.

Liu:2003:GGU

[Liu03] Chuang Liu. Gauge gravity and the unification of natural forces. International Studies in the Philosophy of Science, 17(2):143–159, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859031000160612.

Liu:2015:RIC

[Liu15] Chuang Liu. Re-inflating the conception of scientific representation. International Studies in the Philosophy of Science, 29(1):41–59, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.979671.

Liu:2016:ANF

[Liu16] Chuang Liu. Against the new fictionalism: a hybrid view of scientific models. International Studies in the Philosophy of Science, 30(1):39–54, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2016.1240460.

Livanios:2021:TMM

[Liv21] Vassilis Livanios. The tools of metaphysics and the metaphysics of science.

International Studies in the Philosophy of Science, 34(4):249–252, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2098463.

Lipton:1993:BR

[LLOA93] Peter Lipton, Brendan Larvor, Hans Oberdiek, and Paul Abela. Book reviews. International Studies in the Philosophy of Science, 7(2):191–207, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573463.

Lanz:1995:RGC

[LM95] Peter Lanz and David Mcfarland. On representation, goals and cognition. International Studies in the Philosophy of Science, 9(2):121–133, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573512.

Lockwood:1997:TG

[Loc97] Michael Lockwood. As time goes by.

International Studies in the Philosophy of Science, 11(1):35–51, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573549. Festschrift

for J. R. Lucas.

Lodge:2001:DES

[Lod01] Paul Lodge. The debate over extended substance in Leibniz’s correspondence with De Volder. International Studies in the Philosophy of Science, 15(2):155–165, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120059022.

Logue:1997:RRR

[Log97] James Logue. Resiliency, robustness and rationality of probability judgements. International Studies in the Philosophy of Science, 11(1):21–34, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573548. Festschrift

for J. R. Lucas.

Longino:1997:AST

[Lon97] Helen E. Longino. Alan Sokal’s “transgressing boundaries. International Studies in the Philosophy of Science, 11(2):119–120, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573557.

Lorenzatti:2015:I

[Lor15] Joel J. Lorenzatti. Intuitions. International Studies in the Philosophy of Science, 29(2):221–224, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1119426.

Losee:1992:HSS

[Los92] John Losee. Herbert Simon on scientific discovery. International Studies in the Philosophy of Science, 6(1):41–43, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573410.

Leng:1999:BR

[LPR99] Mary Leng, Kent A. Peacock, and Andrew Reynolds. Book reviews. International Studies in the Philosophy of Science, 13(2):195–204, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573619.

Legrenzi:1993:RGR

[LR93] Paolo Legrenzi and Rino Rumiati. Rationality as “good reasons for”. International Studies in the Philosophy of Science, 7(1):45–49, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573446.

LaFollette:1994:AEL

[LS94] Hugh LaFollette and Niall Shanks.

Animal experimentation: The legacy of Claude Bernard. International Studies in the Philosophy of Science, 8(3):195–210, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573495.

Lucas:1995:PRQ

[Luc95] J. R. Lucas. Prospects for realism in quantum mechanics. International Studies in the Philosophy of Science, 9(3):225–234, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573521.

Lucas:1997:CRT

[Luc97] J. R. Lucas. Comments: Reality and time. International Studies in the Philosophy of Science, 11(1):97–108, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573553. Festschrift

for J. R. Lucas.

Lukes:1993:LI

[Luk93] Steven Lukes. The limits of intelligibility. International Studies in the Philosophy of Science, 7(1):55–59, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573448.

Lyons:2009:NCC

[Lyo09] Timothy D. Lyons. Non-competitor conditions in the scientific realism debate. International Studies in the Philosophy of Science, 23(1):65–84, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590902843393.

Lyons:2013:HIM

[Lyo13] Timothy D. Lyons. A historically informed Modus Ponens against scientific realism: Articulation, critique, and restoration. International Studies in the Philosophy of Science, 27(4):369–392, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.868181.

Lyre:2008:DHM

[Lyr08] Holger Lyre. Does the Higgs mechanism exist? International Studies in the Philosophy of Science, 22(2):119–133, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802496664.

Maleeh:2013:PBC

[MA13] Reza Maleeh and Parisa Amani.

Pragmatism, Bohr, and the Copenhagen interpretation of quantum mechanics. International Studies in the Philosophy of Science, 27(4):353–367, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2013.
868182.

MacIntosh:1992:RBE

[Mac92] J. J. MacIntosh. Robert Boyle’s epistemology: The interaction between scientific and religious knowledge. International Studies in the Philosophy of Science, 6(2):91–121, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573419.

MacCallum:2000:CRW

[Mac00] David MacCallum. Conclusive reasons that we perceive sets. International Studies in the Philosophy of Science, 14(1):25–42, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900111882.

MacIntosh:2001:BBC

[Mac01] J. J. MacIntosh. Boyle, Bentley and Clarke on God, necessity, frigorifick atoms and the void. International Studies in the Philosophy of Science, 15(1):33–50, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590020029297.

Machamer:2004:ACM

[Mac04] Peter Machamer. Activities and causation: The metaphysics and epistemology of mechanisms. International Studies in the Philosophy of Science, 18(1):27–39, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590412331289242.

Machamer:2007:DBJ

[Mac07] Peter Machamer. Daniela Bailer-Jones, 1969–2006. International Studies in the Philosophy of Science, 21(2):211–212, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701498159.

Magnus:2008:DIS

[Mag08] P. D. Magnus. Demonstrative induction and the skeleton of inference. International Studies in the Philosophy of Science, 22(3):303–315, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590802567373.

Maki:2013:SID

[Mäk13] Uskali Mäki. Scientific imperialism: Difficulties in definition, identification, and assessment. International Studies in the Philosophy of Science, 27(3):325–339, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.825496.

Malaterre:2007:ORC

[Mal07] Christophe Malaterre. Organicism and reductionism in cancer research: Towards a systemic approach. International Studies in the Philosophy of Science, 21(1):57–73, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701305792.

Malaterre:2012:WLI

[Mal12] Christophe Malaterre. What is life?

The intellectual pertinence of Erwin Schrödinger. International Studies in the Philosophy of Science, 26(2):229– 231, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
703484.

Mamchur:1986:HRA

[Mam87] Elena Mamchur. The heuristic role of aesthetics in science. International Studies in the Philosophy of Science, 1(2):209–222, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573292.

Mamchur:1990:TIT

[Mam90] E. Mamchur. Is there an ivory tower in reality? International Studies in the Philosophy of Science, 4(1):101–111, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573349.

Mamchur:2010:CSG

[Mam10] Elena Mamchur. Contradictions, synthesis, and the growth of knowledge [book review of MR2106565]. International Studies in the Philosophy of Science, 24(4):429–435, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2010.543352.

Mamchur:2016:ANP

[Mam16] Elena Mamchur. Aristotle’s notion of ‘place’ in the context of presentday physics. International Studies in the Philosophy of Science, 30(4):319– 326, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2017.
                                        1331979.

Mander:1990:BR

[Man90] W. J. Mander. Book review. International Studies in the Philosophy of Science, 4(3):321–323, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573371.

Mander:1991:FHB

[Man91] W. J. Mander. F. H. Bradley and the philosophy of science. International Studies in the Philosophy of Science, 5(1):65–78, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573378.

Marinov:1986:CVS

[Mar87] Marin Marinov. Cognitive values and scientific rationality. International Studies in the Philosophy of Science, 1(2):223–232, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573293.

Marinov:1988:IBE

[Mar89] Marin Marinov. Inference to the best explanation: Van Fraassen and the case of the ‘fifth force’. International Studies in the Philosophy of Science, 3(1):35–50, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573323.

Martinovic:1990:TIT

[Mar90] Ivica Martinović. Theories and intertheory relations in Bosković. International Studies in the Philosophy of Science, 4(3):247–262, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573365.

Marconi:1992:PSP

[Mar92] Diego Marconi. Planetary systems and pomegranites: Doubts about Simon’s essay. International Studies in the Philosophy of Science, 6(1):45–47, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573411.

Marinoff:1996:HBP

[Mar96] Louis Marinoff. How Braess’ paradox solves Newcomb’s problem: Not! International Studies in the Philosophy of Science, 10(3):217–237, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573541. See

[Irv93].

Marquis:1999:MEM

[Mar99] Jean-Pierre Marquis. Mathematical engineering and mathematical change. International Studies in the Philosophy of Science, 13(3):245–259, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573624.

Marcum:2001:CSP

[Mar01] James A. Marcum. Constructing a scientific paper: Howell’s prothrombin laboratory notebook and paper. International Studies in the Philosophy of Science, 15(3):293–310, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120080226.

Marcum:2005:MPS

[Mar05] James A. Marcum. Metaphysical presuppositions and scientific practices: Reductionism and organicism in cancer research. International Studies in the Philosophy of Science, 19(1):31–45, 2005. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590500051076.

Marcum:2008:ISD

[Mar08] James A. Marcum. Instituting science: Discovery or construction of scientific knowledge? International Studies in the Philosophy of Science, 22(2):185–210, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802496755.

Marcum:2010:HSP

[Mar10a] James A. Marcum. Horizon for scientific practice: Scientific discovery and progress. International Studies in the Philosophy of Science, 24(2):187– 215, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
484545.

Martijn:2010:N

[Mar10b] Marije Martijn. Neoplatonism. International Studies in the Philosophy of Science, 24(1):115–118, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903467150.

Martin:2012:TS

[Mar12] Brian Martin. The tyranny of science. International Studies in the Philosophy of Science, 26(1):118–121, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.653116.

Martin:2015:PB

[Mar15] August W. M. Martin. Philosophy of biology. International Studies in the Philosophy of Science, 29(4):441– 444, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1195149.

Martinez:2016:FIE

[Mar16] María Laura Martínez. Foucauldian imprints in the early works of Ian Hacking. International Studies in the Philosophy of Science, 30(1):69–84, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2016.
                                        1240470.

Massimi:2009:SRP

[Mas09] Michela Massimi. Scientific representation: Paradoxes of perspective. International Studies in the Philosophy of Science, 23(3):323–327, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903197799.

Matuszewski:1986:GTM

[Mat87] Roman Matuszewski. On Galileo’s theory of motion: an attempt at a coherent reconstruction. International Studies in the Philosophy of Science, 1(1):124–141, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573285.

Matheson:1998:WNM

[Mat98] Carl Matheson. Why the nomiracles argument fails. International Studies in the Philosophy of Science, 12(3):263–279, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573601.

Matarese:2021:MAW

[Mat21] Vera Matarese. On the methodological arguments for wave-function realism. International Studies in the Philosophy of Science, 34(2):63–80, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.2024786.

Maxwell:2009:MSA

[Max09] Nicholas Maxwell. The metaphysics of science: An account of modern science in terms of principles, laws and theories. International Studies in the Philosophy of Science, 23(2):228–232, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590903007220.

Mayoral:2022:TAK

[May22] Juan V. Mayoral. Travelling around Kuhn’s worlds. International Studies in the Philosophy of Science, 35(3–4):279–286, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2022.2163134.

McAllister:1988:ERR

[McA89] James W. McAllister. The explanative recourse to realism. International Studies in the Philosophy of Science, 3(1):2–18, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573321.

McAllister:1997:LNN

[McA97a] James W. McAllister. Laws of nature, natural history, and the description of the world. International Studies in the Philosophy of Science, 11(3):245–258, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573569.

McAllister:1997:PSN

[McA97b] James W. McAllister. Philosophy of science in The Netherlands. International Studies in the Philosophy of Science, 11(2):191–204, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573563.

McAllister:2001:NET

[McA01] James W. McAllister. New editorial team and policy statement. International Studies in the Philosophy of Science, 15(3):229–230, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120080163.

McAllister:2002:RWA

[McA02] James W. McAllister. Recent work on aesthetics of science. International Studies in the Philosophy of Science, 16(1):7–11, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120118783.

McAllister:2007:ER

[McA07] James W. McAllister. Editor’s report, 2006. International Studies in the Philosophy of Science, 21(2):119–122, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590701497995.

McAllister:2008:CEP

[McA08a] James W. McAllister. Contours of a European philosophy of science. International Studies in the Philosophy of Science, 22(1):1–3, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802280852.

McAllister:2008:ER

[McA08b] James W. McAllister. Editor’s report, 2007. International Studies in the Philosophy of Science, 22(2):115–117, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590802496649.

McArthur:2008:TCS

[McA08c] Dan McArthur. Theory change, structural realism, and the relativised A Priori. International Studies in the Philosophy of Science, 22(1):5–20, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590802280860.

McAllister:2009:ER

[McA09] James W. McAllister. Editor’s report, 2008. International Studies in the Philosophy of Science, 23(2):119–121, 2009. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590903006818.

McAllister:2010:ER

[McA10] James W. McAllister. Editor’s report, 2009. International Studies in the Philosophy of Science, 24(3):237– 239, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2010.
                                        522406.

McAllister:2011:ER

[McA11] James W. McAllister. Editor’s report, 2010. International Studies in the Philosophy of Science, 25(3):203– 204, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2011.
                                        605243.

McAllister:2012:ER

[McA12] James W. McAllister. Editor’s report, 2011. International Studies in the Philosophy of Science, 26(3):237– 239, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
731729.

McAllister:2013:ER

[McA13] James W. McAllister. Editor’s report, 2012. International Studies in the Philosophy of Science, 27(3):233– 234, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2013.
825495.

McAllister:2014:ER

[McA14] James W. McAllister. Editor’s report, 2013. International Studies in the Philosophy of Science, 28(3):231– 233, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
953340.

McAllister:2015:ER

[McA15] James W. McAllister. Editor’s report, 2014. International Studies in the Philosophy of Science, 29(3):229– 231, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1179033.

McAllister:2016:ER

[McA16] James W. McAllister. Editor’s report, 2015. International Studies in the Philosophy of Science, 30(3):197– 199, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2017.
1316112.

McAllister:2017:ER

[McA17] James W. McAllister. Editor’s report, 2016. International Studies in the Philosophy of Science, 31(3):225– 227, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2018.
1464099.

McFarland:1991:DMC

[McF91] David McFarland. Defining motivation and cognition in animals. International Studies in the Philosophy of Science, 5(2):153–170, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573387.

McFarland:1992:ACB

[McF92] David McFarland. Animals as cost-based robots. International Studies in the Philosophy of Science, 6(2):133–153, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573422.

Meier:2014:SWM

[Mei14] Thomas Meier. The structure of the world: Metaphysics and representation. International Studies in the Philosophy of Science, 28(4):443–445, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
979674.

Mellor:1990:LCP

[Mel90] D. H. Mellor. Laws, chances and properties. International Studies in the Philosophy of Science, 4(2):159–170, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573355.

Mellor:2009:DC

[Mel09] D. H. Mellor. Dispositions and causes.

International Studies in the Philosophy of Science, 23(3):327–330, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903197807.

Meynell:2008:WFD

[Mey08] Letitia Meynell. Why Feynman diagrams represent. International Studies in the Philosophy of Science, 22(1):39–59, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802280902.

Maltrana:2022:ETT

[MHB22] Diego Maltrana, Manuel Herrera, and Federico Benitez. Einstein’s theory of theories and mechanicism. International Studies in the Philosophy of Science, 35(2):153–170, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2130661.

Miller:1991:ICT

[Mil91a] Arthur I. Miller. Have incommensurability and causal theory of reference anything to do with actual science? — Incommensurability, no; causal theory, yes. International Studies in the Philosophy of Science, 5(2):97–108, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573383.

Miller:1991:IMC

[Mil91b] Arthur I. Miller. Imagery and meaning, the cognitive science connection. International Studies in the Philosophy of Science, 5(1):35–48, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573376.

Misak:1990:PB

[Mis90] Cheryl Misak. Pragmatism and bivalence. International Studies in the Philosophy of Science, 4(2):171–179, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573356.

Miscevic:1992:MMT

[Mis92] Nenad Miscević. Mental models and thought experiments. International Studies in the Philosophy of Science, 6(3):215–226, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573432.

Miscevic:1996:SRF

[Mis96] Nenad Miscević. Should reason be fragmented? International Studies in the Philosophy of Science, 10(1):23–36, 1996. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573527.

Miscevic:1997:RRN

[Mis97] Nenad Miscević. Resisting the rise of nationalism. International Studies in the Philosophy of Science, 11(3):221–222, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573567.

Miscevic:2001:SCP

[Mis01] Nenad Miscevic. Science, commonsense and philosophy: a defense of continuity (a critique of “network apriorism”). International Studies in the Philosophy of Science, 15(1):19–31, 2001. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590020029288.

Mitchell:1988:CBF

[Mit89] Sandra D. Mitchell. The causal background of functional explanation. International Studies in the Philosophy of Science, 3(2):213–229, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573337.

Mizrahi:2013:AUR

[Miz13] Moti Mizrahi. The argument from underconsideration and relative realism. International Studies in the Philosophy of Science, 27(4):393–407, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.868183.

Mizrahi:2015:HIN

[Miz15] Moti Mizrahi. Historical inductions: New cherries, same old cherry-picking. International Studies in the Philosophy of Science, 29(2):129–148, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1119413.

Mizrahi:2016:HSG

[Miz16] Moti Mizrahi. The history of science as a graveyard of theories: a philosophers’ myth? International Studies in the Philosophy of Science, 30(3):263– 278, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2017.
1316113.

Mizrahi:2017:SPW

[Miz17] Moti Mizrahi. Scientific progress: Why getting closer to truth is not enough. International Studies in the Philosophy of Science, 31(4):415–419, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.

1565213. See replies [Isl17, CT19].

Mizrahi:2020:HTS

[Miz20] Moti Mizrahi. Hypothesis testing in scientific practice: An empirical study. International Studies in the Philosophy of Science, 33(1):1–21, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1788348.

Mncube:2017:SR

[Mnc17] Zinhle Mncube. Is science racist? International Studies in the Philosophy of Science, 31(3):319–321, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1463695.

Mondadori:1993:CBP

[Mon93a] Marco Mondadori. Comments on Boudon’s paper. International Studies in the Philosophy of Science, 7(1):61–63, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573449. See

[Bou93b] and reply [Bou93a].

Montesano:1993:TMR

[Mon93b] Aldo Montesano. On the twofold meaning of rationality in economics. International Studies in the Philosophy of Science, 7(1):65–67, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573450.

Moore:2012:NRM

[Moo12] Dwayne Moore. A non-reductive model of component forces and resultant force. International Studies in the Philosophy of Science, 26(4):359– 380, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
748495.

Morrison:2006:ASA

[Mor06] Margaret Morrison. Applying science and applied science: What’s the difference? International Studies in the Philosophy of Science, 20(1):81–91, 2006. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590600641057.

Mormino:2011:LBS

[Mor11] Gianfranco Mormino. Leibniz: Body, substance, monad. International Studies in the Philosophy of Science, 25(2):191–194, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.574862.

Mosconi:1994:IDO

[Mos94] Giuseppe Mosconi. Induction, deduction and other forms of thought. International Studies in the Philosophy of Science, 8(1):61–64, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573482.

Mozersky:2000:TTS

[Moz00] Joshua M. Mozersky. Time, tense and special relativity. International Studies in the Philosophy of Science, 14(3):221–236, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/026985900437746.

Martinas:1987:AAM

[MR88] Katalin Martinás and László Ropolyi. Analogies: Aristotelian and modern physics. International Studies in the Philosophy of Science, 2(1):1–9, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573299.

Murthi:2009:MI

[MS09] S. K. Arun Murthi and Sundar Sarukkai. Multisemiosis and incommensurability. International Studies in the Philosophy of Science, 23(3):297–311, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903197773.

Muller:2015:PMI

[Mül15] Florian Müller. The pessimistic metainduction: Obsolete through scientific progress? International Studies in the Philosophy of Science, 29(4):393– 412, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1195144.

McAllister:1990:BR

[MV90] James W. McAllister and Boris M.

Velichkovsky. Book reviews. International Studies in the Philosophy of Science, 4(2):221–225, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573360.

Mozersky:1998:BR

[MWRB98] Joshua Mozersky, Rodney Watkins, Andrew Reynolds, and James Robert Brown. Book reviews. International Studies in the Philosophy of Science, 12(1):91–102, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573584.

Myrvold:2003:SEO

[Myr03] Wayne C. Myrvold. On some early objections to Bohm’s theory. International Studies in the Philosophy of Science, 17(1):7–24, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590305233.

Nadeau:1993:BAG

[Nad93] Robert Nadeau. A bad argument for good reasons. International Studies in the Philosophy of Science, 7(1):69–73, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573451.

Nartonis:2001:LSM

[Nar01] David K. Nartonis. Locke–Stewart– Mill: philosophy of science at Dartmouth College, 1771–1854. International Studies in the Philosophy of Science, 15(2):167–175, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120059031.

Navabi:2007:PSI

[Nav07] Ali Akbar Navabi. Philosophy of science in Iran. International Studies in the Philosophy of Science, 21(1):75–89, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590701305800.

Needham:1999:RAC

[Nee99] Paul Needham. Reduction and abduction in chemistry — a response to Scerri. International Studies in the Philosophy of Science, 13(2):169–184, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573616. See

[Sce98, Sce99, Sce00].

Needham:2000:RCS

[Nee00] Paul Needham. Reduction in chemistry — a second response to Scerri. International Studies in the Philosophy of Science, 14(3):317–323, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437827.

Needham:2002:DWH

[Nee02] Paul Needham. The discovery that water is H_{2}O. International Studies in the Philosophy of Science, 16(3):205–226, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859022000013300.

Needham:2004:WDA

[Nee04] Paul Needham. When did atoms begin to do any explanatory work in chemistry? International Studies in the Philosophy of Science, 18(2–3):199–219, 2004. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/0269859042000296521.

Needham:2012:NKT

[Nee12] Paul Needham. Natural kind thingamajigs. International Studies in the Philosophy of Science, 26(1):97–101, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2012.

653117. See reply [Bir12].

Needham:2013:QJP

[Nee13] Paul Needham. Questioning the justification of past science. International Studies in the Philosophy of Science, 27(1):85–93, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.783977.

Nell:2020:ETC

[Nel20] Miranda Nell. Exceptional technologies: A continental philosophy of technology. International Studies in the Philosophy of Science, 33(1):62–65, 2020. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2020.
                                        1829833.

Nesic:2022:TNS

[Nes22] Janko Nesić. Towards a neutralstructuralist theory of consciousness and selfhood. International Studies in the Philosophy of Science, 35(3–4):243–259, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2092825.

Newstead:1994:IRD

[New94] Stephen E. Newstead. Inductive reasoning, deductive reasoning and mental models. International Studies in the Philosophy of Science, 8(1):65–67, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573483.

Newman:2001:CAT

[New01] William R. Newman. Corpuscular alchemy and the tradition of Aristotle’s Meteorology, with special reference to Daniel Sennert. International Studies in the Philosophy of Science, 15(2):145–153, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120059013.

Newman:2012:IMS

[New12] Mark Newman. An inferential model of scientific understanding. International Studies in the Philosophy of Science, 26(1):1–26, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.653118.

Newman:2013:RIM

[New13] Mark Newman. Refining the inferential model of scientific understanding. International Studies in the Philosophy of Science, 27(2):173–197, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.813253.

Nielsen:2008:POB

[Nie08] Katherine Nielsen. The philosophy of Osman bin Bakar. International Studies in the Philosophy of Science, 22(1):81–95, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590802280936.

Niiniluoto:1990:STI

[Nii90] Ilkka Niiniluoto. Should technological imperatives be obeyed? International Studies in the Philosophy of Science, 4(2):181–189, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573357.

Nnaji:2016:CSD

[NL16] John Nnaji and José Luis Luján. The content of science debate in the historiography of the scientific revolution. International Studies in the Philosophy of Science, 30(2):99–109, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2016.1265864.

Norton:2000:WCW

[Nor00] John D. Norton. What can we learn about physical laws from the fact that we have memories only of the past? International Studies in the Philosophy of Science, 14(1):11–23, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900111873.

Norris:2001:PRR

[Nor01] Christopher Norris. Putnam on realism, reference and truth: The problem with quantum mechanics. International Studies in the Philosophy of Science, 15(1):65–91, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590020029314.

Norton:2007:DDB

[Nor07] John D. Norton. Disbelief as the dual of belief. International Studies in the Philosophy of Science, 21(3):231–252, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590701589536.

Novak:2010:TTM

[Nov10] Zsolt Novák. Truth and truthmaking. International Studies in the Philosophy of Science, 24(3):323–326, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
522414.

Newton-Smith:1992:SDP

[NS92] W. H. Newton-Smith. “Scientific discovery as problem solving” by H. A. Simon. International Studies in the Philosophy of Science, 6(1):49–52, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573412. See

[Sim92a].

Newton-Smith:1995:E

[NS95] W. H. Newton-Smith. Editorial. International Studies in the Philosophy of Science, 9(3):197, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599508573517.

Newton-Smith:1986:Ia

[NSW87a] W. H. Newton-Smith and K. V. Wilkes. Introduction. International Studies in the Philosophy of Science, 1(1):2, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598608573278.

Newton-Smith:1986:Ib

[NSW87b] W. H. Newton-Smith and K. V. Wilkes. Introduction. International Studies in the Philosophy of Science, 1(2):141–142, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573288.

Newton-Smith:1987:Ia

[NSW88a] W. H. Newton-Smith and K. V. Wilkes. Introduction. International Studies in the Philosophy of Science, 2(1):5, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573298.

Newton-Smith:1987:Ib

[NSW88b] W. H. Newton-Smith and K. V. Wilkes. Introduction. International Studies in the Philosophy of Science, 2(2):117, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573309.

Newton-Smith:1988:Ia

[NSW89a] W. H. Newton-Smith and K. V.

Wilkes. Introduction. International Studies in the Philosophy of Science, 3(1):1, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573320.

Newton-Smith:1988:Ib

[NSW89b] W. H. Newton-Smith and K. V.

Wilkes. Introduction. International Studies in the Philosophy of Science, 3(2):1, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573330.

Nugayev:1987:GSM

[Nug88] R. M. Nugayev. The genesis and structure of models in the modern theory of gravity. International Studies in the Philosophy of Science, 2(1):84–104, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573304.

Nugayev:1991:FLP

[Nug91] Renat Nugayev. The fundamental laws of physics can tell the truth. International Studies in the Philosophy of Science, 5(1):79–87, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573379.

Nugayev:1996:WDN

[Nug96] Rinat M. Nugayev. Why did the new physics force out the old? International Studies in the Philosophy of Science, 10(2):127–140, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573534.

Oberdiek:1990:TAN

[Obe90] Hans Oberdiek. Technology: Autonomous or neutral. International Studies in the Philosophy of Science, 4(1):67–77, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573346.

Oftedal:2010:DPN

[Oft10] Gry Oftedal. Darwinian populations and natural selection. International Studies in the Philosophy of Science, 24(3):333–336, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2010.522417.

OHear:1994:KEC

[O’H94] Anthony O’Hear. Knowledge in an evolutionary context. International Studies in the Philosophy of Science, 8(2):125–138, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573489.

Okasha:2000:ESB

[Oka00a] Samir Okasha. The explanation of scientific belief: Reply to W. E. Jones. International Studies in the Philosophy of Science, 14(3):305–306, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437809.

Okasha:2000:UTD

[Oka00b] Samir Okasha. The underdetermination of theory by data and the “strong programme” in the sociology of knowledge. International Studies in the Philosophy of Science, 14(3):283–297, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437782. See re-

sponse [Jon00].

Okasha:1996:BR

[OKHT96] Samir Okasha, Joel Katzav, Matthias Hild, and John Taylor. Book reviews. International Studies in the Philosophy of Science, 10(3):257–270, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573544.

Okruhlik:1998:BRO

[Okr98] Kathleen Okruhlik. Book review: Otto Neurath: Philosophy Between Science and Politics by Nancy Cartwright, Jordi Cat, Lola Fleck and Thomas E. Uebel. International Studies in the Philosophy of Science, 12(2):175–191, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573591.

Over:1994:IP

[OM94] D. E. Over and K. I. Manktelow. Induction and probability. International Studies in the Philosophy of Science, 8(1):57–60, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573481.

Okyere-Manu:2016:OLM

[OM16] Beatrice Okyere-Manu. Overpopulation and the lifeboat metaphor: a critique from an African worldview. International Studies in the Philosophy of Science, 30(3):279–289, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1316114.

Okyere-Manu:2017:EET

[OM17] Beatrice Okyere-Manu. Environmental ethics: From theory to practice. International Studies in the Philosophy of Science, 31(1):110–112, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1370931.

PerezOtero:2018:EPA

[P´ 18] Manuel Pérez Otero. An epistemic puzzle about knowledge and rational credence. International Studies in the Philosophy of Science, 32(3–4):195– 206, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/

doi/full/10.1080/02698595.2019.
1704615.

Paavola:2006:HHA

[Paa06] Sami Paavola. Hansonian and Harmanian abduction as models of discovery. International Studies in the Philosophy of Science, 20(1):93–108, 2006. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590600641065.

Pages:2002:DTA

[Pag02] Joan Pagès. The Dretske–Tooley– Armstrong theory of natural laws and the inference problem. International Studies in the Philosophy of Science, 16(3):227–243, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859022000013319.

Palacios:2016:CTA

[Pal16] Patricia Palacios. Chance and temporal asymmetry. International Studies in the Philosophy of Science, 30(2):185–187, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1265865.

Panagiotatou:2017:SRQ

[Pan17] Maria Panagiotatou. Scientific realism and quantum mechanics: Revisiting a controversial relation. International Studies in the Philosophy of Science, 31(3):243–259, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1463696.

Paoletti:2017:EPB

[Pao17] Michele Paolini Paoletti. Empirical physicalism and the boundaries of physics. International Studies in the Philosophy of Science, 31(4):343–362, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2019.
                                        1565207.

Papineau:1987:MF

[Pap88a] David Papineau. Mathematical fictionalism. International Studies in the Philosophy of Science, 2 (2):151–174, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573312.

Pappas:1987:SMB

[Pap88b] George S. Pappas. Science and metaphysics in Berkeley. International Studies in the Philosophy of Science, 2(1):105–114, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598708573305.

Papineau:1990:CMP

[Pap90] David Papineau. Causes and mixed probabilities. International Studies in the Philosophy of Science, 4(1):79–88, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573347.

Parusnikova:1990:PWH

[Par90] Dr Zuzana Parusnikova. Popper’s world 3 & human creativity. International Studies in the Philosophy of Science, 4(3):263–269, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573366.

Parker:2008:FHE

[Par08] Wendy S. Parker. Franklin, Holmes, and the epistemology of computer simulation. International Studies in the Philosophy of Science, 22(2):165–183, 2008. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590802496722.

Parigi:2015:EAD

[Par15] Silvia Parigi. Effluvia, action at a distance, and the challenge of the third causal model. International Studies in the Philosophy of Science, 29(4):351– 368, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1195141.

Parageau:2016:C

[Par16a] Sandrine Parageau. Cavendish. International Studies in the Philosophy of Science, 30(3):307–309, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1316115.

Parkkinen:2016:EDC

[Par16b] Veli-Pekka Parkkinen. Experiment, downward causation, and interventionist levels of explanation. International Studies in the Philosophy of Science, 30(3):245–261, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1316116.

Parigi:2020:MQH

[Par20] Silvia Parigi. Molyneux’s question and the history of philosophy. International Studies in the Philosophy of Science, 33(3):189–193, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.1940464.

Paul:2012:BRL

[Pau12] Herman Paul. Book review: The Life and Thought of Herbert Butterfield: History, Science and God [by Michael Bentley, Cambridge, Cambridge University Press, 2011. xv + 381 pp. ISBN 978-1-107-00397-2]. International Studies in the Philosophy of Science, 26(2):232–235, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.703485.

Paul:2016:TPI

[Pau16] Emily Paul. Time: a philosophical introduction. International Studies in the Philosophy of Science, 30(1):85– 88, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2016.
1240472.

Paya:1995:PAT

[Pay95] A. Paya. Philosophers against “truth”: The cases of Harré and Laudan. International Studies in the Philosophy of Science, 9(3):255–284, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573523.

Paya:2015:CAP

[Pay15] Ali Paya. A critical assessment of the programmes of producing ‘Islamic science’ and ‘Islamisation of science/ knowledge’. International Studies in the Philosophy of Science, 29(3):311– 335, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1179043.

Pietarinen:2014:NLP

[PB14] Ahti-Veikko Pietarinen and Francesco Bellucci. New light on Peirce’s conceptions of retroduction, deduction, and scientific reasoning. International Studies in the Philosophy of Science, 28(4):353–373, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.979667.

Peacock:1998:EPS

[Pea98] Kent A. Peacock. On the edge of a paradigm shift: Quantum nonlocality and the breakdown of peaceful coexistence. International Studies in the Philosophy of Science, 12(2):129–150, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573588.

Peacock:2002:BBQ

[Pea02] Kent A. Peacock. Bub and the barriers to quantum ontology. International Studies in the Philosophy of Science, 16(3):285–289, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859022000013346.

Pearce:2010:COT

[Pea10] Jacob V. Pearce. Constituting objectivity: Transcendental perspectives on modern physics. International Studies in the Philosophy of Science, 24(4):440–443, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2010.543354.

Pechenkin:2002:MIQ

[Pec02] A. A. Pechenkin. Mandelstam’s interpretation of quantum mechanics in comparative perspective. International Studies in the Philosophy of Science, 16(3):265–284, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859022000013337.

Peden:2018:SCA

[Ped18] William Peden. The selective confirmation answer to the paradox of the ravens. International Studies in the Philosophy of Science, 32(3–4):177– 193, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2020.
1768014.

Penrose:1997:UU

[Pen97] Roger Penrose. On understanding understanding. International Studies in the Philosophy of Science, 11(1):7–20, 1997. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573547. Festschrift

for J. R. Lucas.

Perini:2012:IIB

[Per12] Laura Perini. Image interpretation:

Bridging the gap from mechanically produced image to representation. International Studies in the Philosophy of Science, 26(2):153–170, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.703478.

Pernu:2013:DIN

[Per13a] Tuomas K. Pernu. Does the interventionist notion of causation deliver us from the fear of epiphenomenalism? International Studies in the Philosophy of Science, 27(2):157–172, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.813254.

Pero:2013:MMB

[Per13b] Francesca Pero. Models as makebelieve: Imagination, fiction and scientific representation. International Studies in the Philosophy of Science, 27(4):447–450, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.868184.

Perovic:2014:CBF

[Per14] Slobodan Perović. Causation and its basis in fundamental physics. International Studies in the Philosophy of Science, 28(3):347–349, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.953347.

Perovic:2021:OES

[Per21] Slobodan Perović. Observation, experiment, and scientific practice. International Studies in the Philosophy of Science, 34(1):1–20, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2021.1978038.

Petroni:1992:WHS

[Pet92] Angelo M. Petroni. Why have a heuristic of scientific discovery? International Studies in the Philosophy of Science, 6(1):53–55, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573413.

Petroni:1993:CSD

[Pet93a] Angelo M. Petroni. Conventionalism, scientific discovery and the sociology of knowledge. International Studies in the Philosophy of Science, 7(3):225–240, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573467.

Petroni:1993:I

[Pet93b] Angelo M. Petroni. Introduction.

International Studies in the Philosophy of Science, 7(1):3, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573438.

Paya:2006:PRS

[PG06] Ali Paya and Mohammad Amin Ghaneirad. The philosopher and the revolutionary state: How Karl Popper’s ideas shaped the views of Iranian intellectuals. International Studies in the Philosophy of Science, 20(2):185–213, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814381.

Phillips:2014:HCO

[Phi14] Denise Phillips. Hans Christian Ørsted: Reading Nature’s mind. International Studies in the Philosophy of Science, 28(3):341–343, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.953345.

Pietsch:2012:HUC

[Pie12] Wolfgang Pietsch. Hidden underdetermination: a case study in classical electrodynamics. International Studies in the Philosophy of Science, 26(2):125– 151, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
703477.

Pihlstrom:1996:GON

[Pih96] Sami Pihlström. Getting ontologically natural. International Studies in the Philosophy of Science, 10(3):247–256, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573543.

Pineda:2006:MCP

[Pin06] David Pineda. A mereological characterization of physicalism. International Studies in the Philosophy of Science, 20(3):243–266, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600960929.

Pineda:2010:NCC

[Pin10] David Pineda. Non-committal causal explanations. International Studies in the Philosophy of Science, 24(2):147– 170, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2010.
                                        484540.

Pincock:2011:FSP

[Pin11a] Christopher Pincock. Fictions in science: Philosophical essays on modeling and idealization. International Studies in the Philosophy of Science, 25(2):196–199, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.574867.

Pinto:2011:EEN

[Pin11b] Ana Teixeira Pinto. Emergence and embodiment: New essays on second-order systems theory. International Studies in the Philosophy of Science, 25(1):86–89, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.552421.

Piotrowska:2014:GPI

[Pio14] Monika Piotrowska. Genetics and philosophy: An introduction. International Studies in the Philosophy of Science, 28(2):223–226, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.932533.

Pizzorno:1993:AYC

[Piz93] Alessandro Pizzorno. All you can do with reasons. International Studies in the Philosophy of Science, 7(1):75–80, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573452.

Plaud:2010:BRW

[Pla10] Sabine Plaud. Book review: Wittgenstein and Analytic Philosophy: Essays for P. M. S. Hacker, [by Hans-Johan Glock and John Hyman (Eds), Oxford, Oxford University Press, 2009, xxii + 315 pp., ISBN 978-0-19-921323-8, £40.00 (hardback)]. International Studies in the Philosophy of Science, 24(2):223–226, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.484558.

Percival:1994:BR

[PLO94] Philip Percival, J. R. Lucas, and Anthony O’Hear. Book reviews. International Studies in the Philosophy of Science, 8(3):265–276, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573500.

Politi:2017:SII

[Pol17] Vincenzo Politi. Specialisation, interdisciplinarity, and incommensurability. International Studies in the Philosophy of Science, 31(3):301–317, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1463697.

Politi:2020:TNL

[Pol20] Vincenzo Politi. Taxonomies, networks, and lexicons: A study of Kuhn’s post-‘linguistic turn’ philosophy. International Studies in the Philosophy of Science, 33(2):87–103, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1865784.

Popa:2016:CPA

[Pop16] Elena Popa. Causal projectivism, agency, and objectivity. International Studies in the Philosophy of Science, 30(2):147–163, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1265866.

Price:2000:GPC

[Pri00] Carolyn Price. General-purpose content. International Studies in the Philosophy of Science, 14(2):123–133, 2000. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698590050045719.

Preston:1992:BR

[PS92] John M. Preston and Alan Soble. Book reviews. International Studies in the Philosophy of Science, 6(2):155–162, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698599208573423.

Pinnick:2001:CCP

[PS01] Cassandra L. Pinnick and Warren Schmaus. Changing conceptions of the philosophy of science. International Studies in the Philosophy of Science, 15(2):127–131, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120058997.

Psillos:1997:KR

[Psi97] Stathis Psillos. Kitcher on reference.

International Studies in the Philosophy of Science, 11(3):259–272, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599708573570.

Psillos:2000:AEV

[Psi00] Stathis Psillos. Agnostic empiricism versus scientific realism: Belief in truth matters. International Studies in the Philosophy of Science, 14(1):57–75, 2000. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900111909.

Palacio-Vera:2020:PRP

[PV20] Alfonso Palacio-Vera. Popper’s ‘rationality principle’ and ‘epistemic’ rationality: an attempt at reconciliation. International Studies in the Philosophy of Science, 33(4):195–209, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.1960108.

Petersen:2014:ENS

[PZ14] Graciana Petersen and Frank Zenker. From Euler to Navier–Stokes: a spatial analysis of conceptual changes in nineteenth-century fluid dynamics. International Studies in the Philosophy of Science, 28(3):235–253, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.953341.

Psillos:2020:ACI

[PZ20] Stathis Psillos and Lisa Zorzato. Against cognitive instrumentalism. International Studies in the Philosophy of Science, 33(4):247–257, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.1961420.

Qinggang:2016:HCS

[Qin16] Hao Qinggang. A history of Chinese science and technology, volume 1. International Studies in the Philosophy of Science, 30(2):188–190, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1265861.

Qiu:1988:MEE

[Qiu89] Ren-Zong Qiu. Models of explanation and explanation in medicine. International Studies in the Philosophy of Science, 3(2):199–212, 1988–1989. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573336.

Reutlinger:2016:AVC

[RA16] Alexander Reutlinger and Holly Andersen. Abstract versus causal explanations? International Studies in the Philosophy of Science, 30(2):129–146, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2016.
1265867.

Radder:2002:RAO

[Rad02] Hans Radder. Review article: The origin and nature of modern science [Science and modernity: Toward an Integral Theory of Science, Srdan Lelas, Dordrecht, Kluwer Academic Publishers, 2000. xiii + 290 pp., EUR 111, US$129, £80 (hardback), ISBN 0-7923-6305-5; EUR 45, US$42, £28 (paperback), ISBN 1-4020-0247-5]. International Studies in the Philosophy of Science, 16(3):291–295, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859022000013355.

Radder:2010:RSV

[Rad10] Hans Radder. Rethinking science and values. International Studies in the Philosophy of Science, 24(1):107–114, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590903467143.

Radder:2015:HIE

[Rad15] Hans Radder. How inclusive is European philosophy of science? International Studies in the Philosophy of Science, 29(2):149–165, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1119416.

Raley:2007:FEC

[Ral07] Yvonne Raley. The facticity of explanation and its consequences. International Studies in the Philosophy of Science, 21(2):123–135, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701498035.

Ramsey:1993:WRL

[Ram93] Jeffry L. Ramsey. When reduction leads to construction: Design considerations in scientific methodology. International Studies in the Philosophy of Science, 7(3):241–253, 1993. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573468.

Ratzsch:1988:ESS

[Rat89] Del Ratzsch. Explanation, subjunctives and statistical theories. International Studies in the Philosophy of Science, 3(1):80–96, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573326.

Ratcliffe:2005:EPE

[Rat05] Matthew Ratcliffe. An epistemological problem for evolutionary psychology. International Studies in the Philosophy of Science, 19(1):47–63, 2005. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500051126.

Rayan:2009:NDW

[Ray09] Sobhi Rayan. Nominal definition in the writings of Ibn Taymiyya. International Studies in the Philosophy of Science, 23(2):123–141, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903006859.

Root-Bernstein:2002:AC

[RB02] Robert S. Root-Bernstein. Aesthetic cognition. International Studies in the Philosophy of Science, 16(1):61–77, 2002. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120118837.

Read:2020:GCM

[Rea20] James Read. Geometrical constructivism and modal relationalism: Further aspects of the dynamical/geometrical debate. International Studies in the Philosophy of Science, 33(1):23–41, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2020.1813530.

Recio:2017:AVU

[Rec17] Gonzalo L. Recio. Adjusting Venus: The use of maximum elongations in the Almagest and Ptolemy’s Theory of Knowledge. International Studies in the Philosophy of Science, 31(2):113– 131, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2018.
1424762.

Redei:1992:WCN

[Réd92] Miklós Rédei. When can noncommutative statistical inference be Bayesian? International Studies in the Philosophy of Science, 6(2):129–132, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573421.

Reisch:2001:ATD

[Rei01] George A. Reisch. Against a third dogma of logical empiricism: Otto Neurath and “unpredictability in principle”. International Studies in the Philosophy of Science, 15(2):199–209, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120059068.

Renzi:2013:KES

[Ren13] Barbara G. Renzi. Kuhn’s evolutionary social epistemology. International Studies in the Philosophy of Science, 27(1):101–104, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.783978.

Resnik:1994:MCS

[Res94] D. Resnik. Methodological conservatism and social epistemology. International Studies in the Philosophy of Science, 8(3):247–264, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573499.

Reutlinger:2011:TNU

[Reu11] Alexander Reutlinger. A theory of non-universal laws. International Studies in the Philosophy of Science, 25(2):97–117, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.574853.

Reutlinger:2013:CIN

[Reu13] Alexander Reutlinger. Can interventionists be neo-Russellians? Interventionism, the open systems argument, and the arrow of entropy. International Studies in the Philosophy of Science, 27(3):273–293, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.825497.

Reynolds:1999:WH

[Rey99] Andrew Reynolds. What is historicism? International Studies in the Philosophy of Science, 13(3):275–287, 1999. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573626.

Ruttkamp:2005:RRP

[RH05] Emma Ruttkamp and Johannes Heidema. Reviewing reduction in a preferential model-theoretic context. International Studies in the Philosophy of Science, 19(2):123–146, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500249431.

Rini:2012:EPI

[Rin12] Regina A. Rini. Experimental philosophy: An introduction. International Studies in the Philosophy of Science, 26(4):457–460, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.748500.

RiojaNieto:2017:NBP

[Rio17] Ana Rioja Nieto. Niels Bohr and the philosophy of physics: Twentyfirst-century perspectives. International Studies in the Philosophy of Science, 31(4):429–432, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565209.

Robinson:1992:HHI

[Rob92] Don Robinson. On Healey’s holistic interpretation of quantum mechanics. International Studies in the Philosophy of Science, 6(3):227–240, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573433.

Robinson:1995:RO

[Rob95] Daniel N. Robinson. Radical ontologies. International Studies in the Philosophy of Science, 9(3):215–223, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573520.

Rockmann:1998:GLH

[Roc98] Jutta Rockmann. Gravitational lensing and Hacking’s extragalactic irreality. International Studies in the Philosophy of Science, 12(2):151–164, 1998. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573589.

Rockmore:2010:RTR

[Roc10] Tom Rockmore. Reason, truth, and reality. International Studies in the Philosophy of Science, 24(4):449–451, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
543359.

Roe:2017:JDS

[Roe17] Sarah M. Roe. The journey from discovery to scientific change: Scientific communities, shared models, and specialised vocabulary. International Studies in the Philosophy of Science, 31(1):47–67, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1370932.

Rohrlich:2004:RDC

[Roh04] Fritz Rohrlich. Realism despite cognitive antireductionism. International Studies in the Philosophy of Science, 18(1):73–88, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590412331289260.

Rolin:2015:EIE

[Rol15] Kristina Rolin. Economics imperialism and epistemic cosmopolitanism. International Studies in the Philosophy of Science, 29(4):413–429, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1195145.

Romeyn:2005:ET

[Rom05] Jan-Willem Romeyn. Enantiomorphy and time. International Studies in the Philosophy of Science, 19(2):167–190, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590500249498.

Roseberg:1988:HEM

[Rös89] Ulrich Röseberg. Historical explanations in modern physics? The lesson of quantum mechanics. International Studies in the Philosophy of Science, 3(1):68–79, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573325.

Rouse:2011:AWE

[Rou11] Joseph Rouse. Articulating the world: Experimental systems and conceptual understanding. International Studies in the Philosophy of Science, 25(3):243–254, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.605246.

Rowbottom:2008:BTC

[Row08] Darrell P. Rowbottom. The big test of corroboration. International Studies in the Philosophy of Science, 22(3):293–302, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802567357.

Rowbottom:2010:WSP

[Row10] Darrell P. Rowbottom. What scientific progress is not: Against Bird’s epistemic view. International Studies in the Philosophy of Science, 24(3):241–255, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.522407.

Radner:1995:CNS

[RR95] Daisie Radner and Michael Radner. Cognition, natural selection and the intentional stance. International Studies in the Philosophy of Science, 9(2):109–119, 1995. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573511.

Ryghaug:2010:GWC

[RS10] Marianne Ryghaug and Tomas Moe Skjølsvold. The global warming of climate science: Climategate and the construction of scientific facts. International Studies in the Philosophy of Science, 24(3):287–307, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2010.522411.

Rueger:2002:AAE

[Rue02] Alexander Rueger. Aesthetic appreciation of experiments: The case of 18th-century mimetic experiments. International Studies in the Philosophy of Science, 16(1):49–59, 2002. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120118828.

Ruetsche:2003:MSM

[Rue03] Laura Ruetsche. Modal semantics, modal dynamics and the problem of state preparation. International Studies in the Philosophy of Science, 17(1): 25–41, 2003. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        /doi/abs/10.1080/02698590305235.

Ruhmkorff:2013:GLP

[Ruh13] Samuel Ruhmkorff. Global and local pessimistic meta-inductions. International Studies in the Philosophy of Science, 27(4):409–428, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.868185.

Runehov:2011:RSC

[Run11] Anne L. C. Runehov. Religion and science in context: a guide to the debates. International Studies in the Philosophy of Science, 25(3):303–305, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2011.
605254.

Ruphy:2005:WMA

[Rup05] Stéphanie Ruphy. Why metaphysical abstinence should prevail in the debate on reductionism. International Studies in the Philosophy of Science, 19(2):105–121, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500249415.

Russo:2010:CAS

[Rus10] Federica Russo. Are causal analysis and system analysis compatible approaches? International Studies in the Philosophy of Science, 24(1):67–90, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

http://www.tandfonline.com/doi/
full/10.1080/02698590903467127.

Rush:2013:AMS

[Rus13] Penelope Rush. The applicability of mathematics in science: Indispensability and ontology. International Studies in the Philosophy of Science, 27(2):219–222, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.813255.

Russo:2014:WIH

[Rus14] Federica Russo. What invariance is and how to test for it. International Studies in the Philosophy of Science, 28(2):157–183, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.932528.

Ruyant:2021:SIP

[Ruy21] Quentin Ruyant. Symmetries, indexicality and the perspectivist stance. International Studies in the Philosophy of Science, 34(1):21–39, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2021.2001723.

Ruzzene:2011:CSR

[Ruz11] Attilia Ruzzene. The case study as research method: a practical handbook. International Studies in the Philosophy of Science, 25(3):293–296, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.605251.

Russo:2007:ICH

[RW07] Federica Russo and Jon Williamson. Interpreting causality in the health sciences. International Studies in the Philosophy of Science, 21(2):157–170, 2007. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL

                                        http://www.tandfonline.com/doi/
                                        full/10.1080/02698590701498084.

Saatsi:2007:LHN

[Saa07] Juha T. Saatsi. Living in harmony: Nominalism and the explanationist argument for realism. International Studies in the Philosophy of Science, 21(1):19–33, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701305743.

Salvaggio:1993:STR

[Sal93] Salvino A. Salvaggio. A synthetic theory of rationality: Propositions and aporia. International Studies in the Philosophy of Science, 7(1):81–84, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573453.

Sankey:1996:NNC

[San96] Howard Sankey. Normative naturalism and the challenge of relativism: Laudan versus Worrall on the justification of methodological principles. International Studies in the Philosophy of Science, 10(1):37–51, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573528.

Sankey:1998:TI

[San98] Howard Sankey. Taxonomic incommensurability. International Studies in the Philosophy of Science, 12(1):7–16, 1998. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573578.

Sarkar:2000:EEU

[Sar00] Husain Sarkar. Empirical equivalence and underdetermination. International Studies in the Philosophy of Science, 14(2):187–197, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590050045755.

Saridakis:2016:IRM

[Sar16] Emmanuel Saridakis. Information, reality, and modern physics. International Studies in the Philosophy of Science, 30(4):327–341, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1331980.

Schickore:2013:UMM

[SC13] Jutta Schickore and Klodian Coko. Using multiple means of determination. International Studies in the Philosophy of Science, 27(3):295–313, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.825498.

Scerri:1998:PNA

[Sce98] Eric R. Scerri. Popper’s naturalized approach to the reduction of chemistry. International Studies in the Philosophy of Science, 12(1):33–44, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573581. See

response [Nee99] and replies [Sce99, Sce00].

Scerri:1999:RN

[Sce99] Eric R. Scerri. Response to Needham. International Studies in the Philosophy of Science, 13(2):185–192, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573617. See

[Sce98, Nee99, Sce00].

Scerri:2000:SRP

[Sce00] Eric R. Scerri. Second response to Paul Needham. International Studies in the Philosophy of Science, 14(3):307–315, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900437818.  See

[Sce98, Nee99, Sce99].

Schneider:1998:GPH

[Sch98] Ulrich Johannes Schneider. GDR philosophy is history. International Studies in the Philosophy of Science, 12(3):213–215, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573596.

Schickore:2009:SJP

[Sch09] Jutta Schickore. Studying justificatory practice: An attempt to integrate the history and philosophy of science. International Studies in the Philosophy of Science, 23(1):85–107, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590902843401.

Schiaffonati:2011:PEE

[Sch11] Viola Schiaffonati. Philosophy and engineering: An emerging agenda. International Studies in the Philosophy of Science, 25(1):89–92, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.552422.

Schlosser:2012:FWM

[Sch12] Markus Schlosser. Free will and modern science. International Studies in the Philosophy of Science, 26(4):463– 466, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
748502.

Schyfter:2017:SSE

[Sch17] Pablo Schyfter. Science as social existence: Heidegger and the sociology of scientific knowledge. International Studies in the Philosophy of Science, 31(4):432–435, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2019.1565210.

Schliesser:2022:BBW

[Sch22] Eric Schliesser. Bad beliefs: Why they happen to good people. International Studies in the Philosophy of Science, 35(2):189–203, 2022. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2119492.

Sciortino:2016:SRH

[Sci16] Luca Sciortino. Styles of reasoning, human forms of life, and Relativism. International Studies in the Philosophy of Science, 30(2):165–184, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2016.1265868.

Seager:1986:SET

[Sea87] William E. Seager. Scientific explanation and the trial of Galileo. International Studies in the Philosophy of Science, 1(2):176–195, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573290.

Seager:1990:IP

[Sea90] William Seager. Instrumentalism in psychology. International Studies in the Philosophy of Science, 4(2):191–203, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573358.

Seager:1999:RN

[Sea99] William Seager. The reality of now.

International Studies in the Philosophy of Science, 13(1):69–82, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573608.

Sedivy:1996:CNP

[Sed96] Sonia Sedivy. Conventional naturalism: a perceptualist account of pictorial representation. International Studies in the Philosophy of Science, 10(2):103–125, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573533.

Seifert:2016:EPC

[Sei16] Vanessa A. Seifert. Essays in the philosophy of chemistry. International Studies in the Philosophy of Science, 30(4):415–417, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331981.

Sesardic:2011:MSB

[Ses11] Neven Sesardić. The mirage of a space between nature and nurture. International Studies in the Philosophy of Science, 25(4):417–420, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.623371.

Schank:1992:DSQ

[SH92] Roger C. Schank and Lucian P. Hughes. Discovering scientific questions. International Studies in the Philosophy of Science, 6(1):57–59, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573414.

Shah:2007:IJA

[Sha07] Mehul Shah. Is it justifiable to abandon all search for a logic of discovery? International Studies in the Philosophy of Science, 21(3):253–269, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701589544.

Shahvisi:2016:THP

[Sha16] Arianne Shahvisi. Tainted: How philosophy of science can expose bad science. International Studies in the Philosophy of Science, 30(2):193–196, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2016.
1265869.

Shaw:2020:DGS

[Sha20] Jamie Shaw. Duhem on good sense and theory pursuit: From virtue to social epistemology. International Studies in the Philosophy of Science, 33(2): 67–85, 2020. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2021.
1888191.

Sheehan:2007:MSS

[She07] Helena Sheehan. Marxism and science studies: a sweep through the decades. International Studies in the Philosophy of Science, 21(2):197–210, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701498126.

Shech:2021:DSE

[She21] Elay Shech. Darwinian-selectionist explanation, radical theory change, and the observable-unobservable dichotomy. International Studies in the Philosophy of Science, 34(4):221–241, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.

2092824. See reply [Wra21].

Shogenji:2018:HPS

[Sho18] Tomoji Shogenji. Hume’s problem solved: The optimality of metainduction. International Studies in the Philosophy of Science, 32(3–4):229– 231, 2018. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2020.
                                        1767892.

Sialaros:2019:GTB

[Sia19] Michalis Sialaros. Geometry: The third book of foundations. International Studies in the Philosophy of Science, 32(1):75–77, 2019. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2019.1615667.

Sikic:1992:DAS

[Sik92] Zvonimir Sikić. The diagonal argument — a study of cases. International Studies in the Philosophy of Science, 6(3):191–203, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573430.

Sikic:1996:WN

[Šik96] Zvonimir Šikić. What are numbers? International Studies in the Philosophy of Science, 10(2):159–171, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599608573536.

Simon:1992:SDPa

[Sim92a] Herbert A. Simon. Scientific discovery as problem solving. International Studies in the Philosophy of Science, 6(1):3–14, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573403. See

comments [Gil92, Hes92, NS92, Tra92] and response [Sim92b].

Simon:1992:SDPb

[Sim92b] Herbert A. Simon. Scientific discovery as problem solving: Reply to critics. International Studies in the Philosophy of Science, 6(1):69–88, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573417.

Simoes:2016:PCG

[Sim16] Ana Simões. Philosophical chemistry:

Genealogy of a scientific field. International Studies in the Philosophy of Science, 30(3):311–314, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1316117.

Simons:2019:KIS

[Sim19] Massimiliano Simons. The Kuhnian image of science: Time for a decisive transformation? International Studies in the Philosophy of Science, 32(1): 78–80, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2019.
1615668.

Sinaceur:2017:NN

[Sin17] Hourya Benis Sinaceur. Neurophilosophy of number. International Studies in the Philosophy of Science, 31(1): 1–25, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2017.
1370925.

Skow:2010:SAG

[Sko10] Bradford Skow. On a symmetry argument for the guidance equation in Bohmian mechanics. International Studies in the Philosophy of Science, 24(4):393–410, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2010.543350.

Shu-li:1991:ASD

[Sl91] Ji Shu-li. The antinomy of science and democracy in modern China. International Studies in the Philosophy of Science, 5(2):109–130, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573384.

Slater:2000:NNP

[Sla00] Ian Slater. Normative naturalism and Popperian views on reduction. International Studies in the Philosophy of Science, 14(3):325–326, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/026985900437836.

Slezak:1991:BBB

[Sle91] Peter Slezak. Bloor’s bluff: Behaviourism and the strong programme. International Studies in the Philosophy of Science, 5(3):241–256, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599108573397.

Sleutels:2009:BRP

[Sle09] Jan Sleutels. Beyond reduction: Philosophy of mind and post-reductionist philosophy of science. International Studies in the Philosophy of Science, 23(2):233–236, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903007279.

Slowik:2005:SOS

[Slo05] Edward Slowik. Spacetime, ontology, and structural realism. International Studies in the Philosophy of Science, 19(2):147–166, 2005. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590500249456.

Sypel:1991:BR

[SM91] Roland Sypel and Radim Marada.

Book reviews. International Studies in the Philosophy of Science, 5(2):187– 190, 1991. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
/doi/abs/10.1080/02698599108573389.

Sfendoni-Mentzou:1994:LNA

[SM94] Demetra Sfendoni-Mentzou. Laws of nature: Ante res or in rebus?. International Studies in the Philosophy of Science, 8(3):229–242, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573497. See re-

ply [Bro94].

Smith:1994:RMT

[Smi94] Edward E. Smith. Relating ‘a model theory’ to other research in induction. International Studies in the Philosophy of Science, 8(1):69–71, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573484.

Soe:2019:WI

[Søe19] Sille Obelitz Søe. What is information? International Studies in the Philosophy of Science, 32(1):73–75, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2019.1615665.

Soler:2014:ARS

[Sol14] Léna Soler. Against robustness? Strategies to support the reliability of scientific results. International Studies in the Philosophy of Science, 28(2):203–215, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2014.932530.

Spurrett:1999:FLC

[Spu99] David Spurrett. Fundamental laws and the completeness of physics. International Studies in the Philosophy of Science, 13(3):261–274, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573625.

Spurrett:2001:CLC

[Spu01] David Spurrett. Cartwright on laws and composition. International Studies in the Philosophy of Science, 15(3):253–268, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120080190.

Seselja:2014:CPV

[SS14] Dunja Seselja and Christian Straßer.

Concerning Peter Vickers’s recent treatment of ‘paraconsistencitis’. International Studies in the Philosophy of Science, 28(3):325–340, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.953344.

Sevilla:2015:KPN

[SS15] Ana Sevilla and Elisa Sevilla. Knowledge production in non-European spaces of modernity: The Society of Jesus and the circulation of Darwinian ideas in Postcolonial Ecuador, 1860– 1890. International Studies in the Philosophy of Science, 29(3):233–250, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1179036.

Shaw:2022:FPP

[SS22] Jamie Shaw and Michael T. Stuart.

Feyerabend and the philosophy of physics. International Studies in the Philosophy of Science, 35(1):1–4, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2022.2193369.

Sterrett:2006:MMM

[Ste06] Susan G. Sterrett. Models of machines and models of phenomena. International Studies in the Philosophy of Science, 20(1):69–80, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590600641024.

Steel:2010:WIP

[Ste10] Daniel Steel. What if the principle of induction is normative? Formal learning theory and Hume’s problem. International Studies in the Philosophy of Science, 24(2):171–185, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2010.484544.

Steel:2011:CPR

[Ste11] Daniel Steel. On not changing the problem: a reply to Howson [MR2845407]. International Studies in the Philosophy of Science, 25(3):285– 291, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2011.

605250. See [How11, BTB16].

Schmidt:1996:BR

[STL^{+}96] Thomas Schmidt, Jens Timmermann, Brendan Larvor, Samir Okasha, Peter Goldie, Louis Caruana, Hasok Chang, Paul Castell, and Rom Harré. Book reviews. International Studies in the Philosophy of Science, 10(1):67–90, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573530.

Strand:2017:MRB

[Str17] Anders Strand. The multiple realization book. International Studies in the Philosophy of Science, 31(2):218– 221, 2017. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2018.
1424767.

Suarez:2003:SRA

[Sua03] Mauricio Sua’rez. Scientific representation: against similarity and isomorphism. International Studies in the Philosophy of Science, 17(3):225–244, 2003. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/0269859032000169442.

Suarez:2013:FCS

[Suá13] Mauricio Suárez. Fictions, conditionals, and stellar astrophysics. International Studies in the Philosophy of Science, 27(3):235–252, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.825499.

Svetlova:2014:MBA

[Sve14] Ekaterina Svetlova. Modelling beyond application: Epistemic and nonepistemic values in modern science. International Studies in the Philosophy of Science, 28(1):79–98, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915656.

Svorcova:2015:TTD

[Svo15] Jana Svorcová. Towards a theory of development. International Studies in the Philosophy of Science, 29(3):344– 347, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2015.
                                        1179047.

Swan:2011:BML

[Swa11] Liz Stillwaggon Swan. The brain and the meaning of life. International Studies in the Philosophy of Science, 25(3):297–299, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.605252.

Szabo:2003:FSP

[Sza03] Laśzló E. Szabó. Formal systems as physical objects: a physicalist account of mathematical truth. International Studies in the Philosophy of Science, 17(2):117–125, 2003. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/0269859031000160568.

Teira:2013:PMC

[Tei13] David Teira. Philosophy of medicine: Causality, evidence and explanation. International Studies in the Philosophy of Science, 27(4):456–458, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.868186.

Tempczyk:1991:RDR

[Tem91] Michal Tempczyk. Random dynamics and the research programme of classical mechanics. International Studies in the Philosophy of Science, 5(3):227–239, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573396.

Tempczyk:1996:FGC

[Tem96] Michał Tempczyk. Fractal geometry — the case of a rapid career. International Studies in the Philosophy of Science, 10(1):53–65, 1996. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599608573529.

Tetens:1988:TCI

[Tet89] Holm Tetens. Teleology and the concept of information. International Studies in the Philosophy of Science, 3(2):230–237, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573338.

Thagard:2002:CCP

[Tha02] Paul Thagard. Curing cancer? Patrick Lee’s path to the reovirus treatment. International Studies in the Philosophy of Science, 16(1):79–93, 2002. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698590120118846.

Theodorou:2012:CEP

[The12] Panos Theodorou. Cognitive existentialism, phenomenology, and philosophy of science: Stimulating the dialogue. International Studies in the Philosophy of Science, 26(3):335–343, 2012. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2012.
731734.

Theurer:2014:CBT

[The14] Kari L. Theurer. Complexity-based theories of emergence: Criticisms and constraints. International Studies in the Philosophy of Science, 28(3):277– 301, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
953342.

Thebault:2015:PTM

[Thé15] Karim Thébault. Physical theory: Method and interpretation. International Studies in the Philosophy of Science, 29(4):438–441, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1195148.

Thompson:1992:MBS

[Tho92] Paul Thompson. Mathematics in the biological sciences. International Studies in the Philosophy of Science, 6(3):241–248, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573434.

Thoren:2014:RUC

[Tho14] Henrik Thorén. Resilience as a unifying concept. International Studies in the Philosophy of Science, 28(3):303– 324, 2014. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2014.
953343.

Thyssen:2022:WCE

[Thy22] Pieter Thyssen. What is a chemical element? A collection of essays by chemists, philosophers, historians, and educators. International Studies in the Philosophy of Science, 35(3–4):286– 289, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2023.
2195306.

Tiles:1986:NRI

[Til87a] James Tiles. The normative role of images of science. International Studies in the Philosophy of Science, 1(2):196–208, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598708573291.

Tiles:1986:MMB

[Til87b] Mary Tiles. Mathesis and the masculine birth of time. International Studies in the Philosophy of Science, 1(1):16–35, 1986–1987. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598608573280.

Tiles:1987:SDS

[Til88] Mary Tiles. Scientific dream space: Symbolic forms and scientific theories. International Studies in the Philosophy of Science, 2(2):189–204, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573314.

Tiles:1990:ITH

[Til90a] J. E. Tiles. Information technology from Homer to DENDRAL. International Studies in the Philosophy of Science, 4(2):205–220, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573359.

Tiles:1990:HBT

[Til90b] Mary Tiles. Of heroes and butterflies: Technological dreams and human realities. International Studies in the Philosophy of Science, 4(1):89–100, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599008573348.

Tiles:1998:ACD

[Til98] Mary Tiles. Adding a comparative dimension. International Studies in the Philosophy of Science, 12(2):109–110, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599808573586.

Tiles:1999:BAR

[Til99] Mary Tiles. Balancing acts: Rational agency and efficacious action. International Studies in the Philosophy of Science, 13(3):289–300, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599908573627.

Tishchenko:1988:SEB

[Tis89] V. J. Tishchenko. Social experiment and biological theory. International Studies in the Philosophy of Science, 3(2):238–247, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573339.

Talmont-Kaminski:2010:EE

[TK10] Konrad Talmont-Kaminski. Epistemology and emotions. International Studies in the Philosophy of Science, 24(2):229–233, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2010.484561.

Trpin:2022:I

[TO22] Borut Trpin and Barbara Osimani.

Introduction. International Studies in the Philosophy of Science, 35(3–4):209–210, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2023.2218190.

Todd:2008:UTB

[Tod08] Cain S. Todd. Unmasking the truth beneath the beauty: Why the supposed aesthetic judgements made in science May not be aesthetic at all. International Studies in the Philosophy of Science, 22(1):61–79, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590802280910.

Toon:2012:SSR

[Too12] Adam Toon. Similarity and scientific representation. International Studies in the Philosophy of Science, 26(3):241–257, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.731730.

Trautteur:1992:PSC

[Tra92] Giuseppe Trautteur. Problems with symbols. A commentary to Herbert Simon, “Scientific discovery as problem solving”. International Studies in the Philosophy of Science, 6(1):61–64, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599208573415. See

[Sim92a].

Trafimow:2020:TMP

[Tra20] David Trafimow. A taxonomy of major premises and implications for falsification and verification. International Studies in the Philosophy of Science, 33(4):211–229, 2020. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2021.1964845.

Tresker:2021:TER

[Tre21] Steven Tresker. Treatment effectiveness and the Russo–Williamson thesis, EBM+, and Bradford Hill’s viewpoints. International Studies in the Philosophy of Science, 34(3):131–158, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2054396.

Trobok:2011:MR

[Tro11] Majda Trobok. Mathematics and reality. International Studies in the Philosophy of Science, 25(4):413–417, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2011.623370.

Tsou:2007:HLE

[Tso07] Jonathan Y. Tsou. Hacking on the looping effects of psychiatric classifications: What is an interactive and indifferent kind? International Studies in the Philosophy of Science, 21(3):329–344, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701589601.

Tuchanska:1990:CRR

[Tuc90] Barbara Tuchańska. Can relativism be reconciled with realism and causalism? International Studies in the Philosophy of Science, 4(3):285–294, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573368.

Tugby:2013:GTM

[Tug13] Matthew Tugby. Graph-theoretic models of dispositional structures. International Studies in the Philosophy of Science, 27(1):23–39, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.783979.

Tulkki:2016:OSN

[Tul16] Leena Tulkki. Objectivity in science: New perspectives from science and technology studies. International Studies in the Philosophy of Science, 30(3):314–317, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1316118.

Tympas:2021:AIA

[Tym21] Aristotle Tympas. Artifictional intelligence: Against humanity’s surrender to computers. International Studies in the Philosophy of Science, 34(1): 59–61, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2021.
                                        1980650.

Uebel:2001:CNE

[Ueb01] Thomas E. Uebel. Carnap and Neurath in exile: Can their disputes be resolved? International Studies in the Philosophy of Science, 15(2):211–220, 2001. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698590120059077.

Ugaglia:2015:AEE

[Uga15] Monica Ugaglia. Aristotle’s empiricism: Experience and mechanics in the 4th Century BC. International Studies in the Philosophy of Science, 29(1):99– 101, 2015. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2015.
1071553.

Vahid:1994:EBH

[Vah94] H. Vahid. Experience and belief:

Haack on the problem of empirical basis. International Studies in the Philosophy of Science, 8(2):139–140, 1994. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599408573490.

Vaidman:1998:SEN

[Vai98] Lev Vaidman. On schizophrenic experiences of the neutron or why we should believe in the many-worlds interpretation of quantum theory. International Studies in the Philosophy of Science, 12(3):245–261, 1998. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599808573600.

Valentine:2012:HPP

[Val12] Elizabeth R. Valentine. History and philosophy of psychology. International Studies in the Philosophy of Science, 26(4):460–463, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2012.748501.

Vandevelde:1993:RR

[Van93] Toon Vandevelde. Rationality and reasonableness. International Studies in the Philosophy of Science, 7(1):85–86, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599308573454.

VanKerkhove:2010:PMP

[Van10] Bart Van Kerkhove. The philosophy of mathematical practice. International Studies in the Philosophy of Science, 24(1):118–122, 2010. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590903467168.

Vaz:2016:PPP

[Vaz16] Sofia Guedes Vaz. Philosophy and the precautionary principle: Science, evidence, and environmental policy. International Studies in the Philosophy of Science, 30(1):95–98, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1195152.

vanBrakel:1987:PIM

[vBG88] J. van Brakel and J. P. M. Geurts. Pragmatic identity of meaning and metaphor. International Studies in the Philosophy of Science, 2(2):205–226, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698598808573315.

vandenBerg:2014:KOE

[vdB14] Hein van den Berg. Kant’s organicism: Epigenesis and the development of critical philosophy. International Studies in the Philosophy of Science, 28(1):99–101, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915657.

vanderSchaar:2014:BRB

[vdS14] Maria van der Schaar. Book review: The Bloomsbury Companion to Analytic Philosophy. International Studies in the Philosophy of Science, 28(4):437–439, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.979672.

vanderTuin:2014:LWE

[vdT14] Iris van der Tuin. Life, war, Earth:

Deleuze and the sciences. International Studies in the Philosophy of Science, 28(2):226–229, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.932534.

Venturelli:2021:CCV

[Ven21] A. Nicolás Venturelli. Conceptual change in visual neuroscience: The receptive field concept. International Studies in the Philosophy of Science, 34(1):41–57, 2021. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2021.2002675.

Vetter:2011:LLS

[Vet11] Barbara Vetter. Laws and lawmakers: Science, metaphysics, and the laws of nature. International Studies in the Philosophy of Science, 25(1):83– 86, 2011. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2011.
                                        552420.

Vervoort:2015:MOD

[VG15] Louis Vervoort and Yves Gingras. Macroscopic oil droplets mimicking quantum behaviour: How far can we push an analogy? International Studies in the Philosophy of Science, 29(3):271–294, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1179040.

Votsis:2007:BR

[VHO^{+}07] Ioannis Votsis, Katherine Hawley, Robert J. O’Hara, Lesley B. Cormack, and Diane Greco. Book reviews. International Studies in the Philosophy of Science, 21(1):103–117, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590701306048.

Viale:1999:CCC

[Via99] Riccardo Viale. Causal cognition and causal realism. International Studies in the Philosophy of Science, 13(2):151–167, 1999. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599908573615.

Vicente:2006:CCP

[Vic06] Agustín Vicente. On the causal completeness of physics. International Studies in the Philosophy of Science, 20(2):149–171, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600814332.

Vicente:2013:WLE

[Vic13] Agustín Vicente. Where to look for emergent properties. International Studies in the Philosophy of Science, 27(2):137–156, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.813256.

Vistarini:2014:STS

[Vis14] Tiziana Vistarini. String theory and the scientific method. International Studies in the Philosophy of Science, 28(1):108–111, 2014. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2014.915658.

Vitasovi:2016:ET

[Vit16] Daria Vitasovi. Experiencing time. International Studies in the Philosophy of Science, 30(4):409–412, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331982.

Vorontsov:1988:PSS

[Vor89] Nikolay N. Vorontsov. The problem of species and speciation. International Studies in the Philosophy of Science, 3(2):173–189, 1988–1989. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598908573334.

Wagner:2009:MVI

[Wag09] Roy Wagner. Mathematical variables as indigenous concepts. International Studies in the Philosophy of Science, 23(1):1–18, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590902843351.

Wagner:2013:HPI

[Wag13] Roy Wagner. A historically and philosophically informed approach to mathematical metaphors. International Studies in the Philosophy of Science, 27(2):109–135, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.813257.

Walker:2013:ESI

[Wal13] Rebecca L. Walker. The ethics of species: An introduction. International Studies in the Philosophy of Science, 27(2):225–228, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.813258.

Waskan:2008:KCI

[Was08] Jonathan Waskan. Knowledge of counterfactual interventions through cognitive models of mechanisms. International Studies in the Philosophy of Science, 22(3):259–275, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590802567308.

Watkins:1992:BVK

[Wat92] John Watkins. Has BACON vindicated Kant? International Studies in the Philosophy of Science, 6(1):65–67, 1992. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573416.

Wayne:2015:CRE

[Way15] Andrew Wayne. Causal relations and explanatory strategies in physics. International Studies in the Philosophy of Science, 29(1):75–89, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1071552.

Walker:2006:AFT

[WĆ06] Mark A. Walker and Milan M.

Ćirković. Astrophysical fine tuning, naturalism, and the contemporary design argument. International Studies in the Philosophy of Science, 20(3):285–307, 2006. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590600960945.

Weber:2009:HPC

[Web09] Erik Weber. How probabilistic causation can account for the use of mechanistic evidence. International Studies in the Philosophy of Science, 23(3):277–295, 2009. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590903197757.

Weinert:1995:DQT

[Wei95] F. Weinert. The Duhem–Quine thesis revisited. International Studies in the Philosophy of Science, 9(2):147– 156, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        /doi/abs/10.1080/02698599508573514.

Weinstein:2013:STS

[Wei13] Steven Weinstein. Space, time, and stuff. International Studies in the Philosophy of Science, 27(1):98–101, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2013.783980.

Weinert:2016:TSG

[Wei16] Friedel Weinert. The time-symmetric Gold universe reconsidered. International Studies in the Philosophy of Science, 30(3):231–243, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2017.1316119.

Whiteley:2015:BVS

[Whi15] Ella Whiteley. Beyond versus: The struggle to understand the interaction of nature and nurture. International Studies in the Philosophy of Science, 29(3):347–350, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1179048.

Whyte:1990:PPL

[Why90] J. T. Whyte. The psycho-physical laws of intentionality. International Studies in the Philosophy of Science, 4(3):295–304, 1990. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573369.

Wilkes:1991:MMC

[Wil91] K. V. Wilkes. Of mice and men: The comparative assumption in psychology. International Studies in the Philosophy of Science, 5(1):3–19, 1991. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599108573374.

Wilkes:1992:ESW

[Wil92] Kathy Wilkes. Editorial special war issue. International Studies in the Philosophy of Science, 6(3):171, 1992. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599208573426.

Wilson:1995:OMD

[Wil95] Fred Wilson. Once more to dissolve the ravens. International Studies in the Philosophy of Science, 9(2):135–146, 1995. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599508573513.

Williams:2011:FVD

[Wil11] James Williams. The force of the virtual: Deleuze, science, and philosophy. International Studies in the Philosophy of Science, 25(3):300–302, 2011. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2011.605253.

Williamson:2019:ECC

[Wil19] Jon Williamson. Establishing causal claims in medicine. International Studies in the Philosophy of Science, 32(1): 33–61, 2019. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/02698595.2019.
                                        1630927.

Wong:1997:BR

[WMB97] James Wong, David Morris, and James Robert Brown. Book reviews. International Studies in the Philosophy of Science, 11(2):207–214, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573565.

Wolters:2015:GPC

[Wol15] Gereon Wolters. Globalized parochialism: Consequences of English as Lingua Franca in philosophy of science. International Studies in the Philosophy of Science, 29(2):189–200, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2015.1119420.

Wong:1997:MTP

[Won97] James Wong. The “making” of teenage pregnancy. International Studies in the Philosophy of Science, 11(3):273–288, 1997. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599708573571.

Woodward:2004:CCE

[Woo04] James Woodward. Counterfactuals and causal explanation. International Studies in the Philosophy of Science, 18(1):41–72, 2004. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590412331289251.

Woodcock:2007:BPN

[Woo07] Brian A. Woodcock. Bloch’s Paradox and the nonlocality of chance. International Studies in the Philosophy of Science, 21(2):137–156, 2007. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590701498050.

Worrall:1990:RSS

[Wor90] John Worrall. Rationality, sociology and the symmetry thesis. International Studies in the Philosophy of Science, 4(3):305–319, 1990. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599008573370.

Webster:1993:BR

[WR93] Charles Webster and Don Robinson. Book reviews. International Studies in the Philosophy of Science, 7(3):267–272, 1993. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698599308573470.

Wray:2008:AUG

[Wra08] K. Brad Wray. The argument from underconsideration as grounds for anti-realism: a defence. International Studies in the Philosophy of Science, 22(3):317–326, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802567399.

Wray:2015:PIF

[Wra15] K. Brad Wray. Pessimistic inductions: Four varieties. International Studies in the Philosophy of Science, 29(1):61–73, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1071551.

Wray:2021:IRR

[Wra21] K. Brad Wray. Induction, rationality, and the realism/anti-realism debate: a reply to Shech. International Studies in the Philosophy of Science, 34(4):243–247, 2021. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.

2113627. See [She21].

Wylie:1987:SAR

[Wyl88] Alison Wylie. ‘Simple’ analogy and the role of relevance assumptions: Implications of archaeological practice. International Studies in the Philosophy of Science, 2(2):134–150, 1987–1988. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573311.

Wylie:2000:RON

[Wyl00] Alison Wylie. Rethinking objectivity: Nozick’s neglected third option.

International Studies in the Philosophy of Science, 14(1):5–9, 2000. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/026985900111864.

Yaghmaie:2017:HCP

[Yag17] Aboutorab Yaghmaie. How to characterise pure and applied science. International Studies in the Philosophy of Science, 31(2):133–149, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1424763.

Yang:2008:MDP

[Yan08] Andrew S. Yang. Matters of demarcation: Philosophy, biology, and the evolving fraternity between disciplines. International Studies in the Philosophy of Science, 22(2):211–225, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802496789.

Yan:2013:REV

[Yan13] Wong Muk Yan. Reliability and external validity of neurobiological experiments. International Studies in the Philosophy of Science, 27(4):429– 446, 2013. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2013.
868188.

Ye:2010:NAE

[Ye10] Feng Ye. Naturalism and abstract entities. International Studies in the Philosophy of Science, 24(2):129–146, 2010. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2010.
484539.

Yaghmaie:2022:RSH

[YKMM22] Aboutorab Yaghmaie, Bijan Ahmadi Kakavandi, Saeed Masoumi, and Morteza Moniri. Representation and spacetime: The hole argument revisited. International Studies in the Philosophy of Science, 35(2):171–188, 2022. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/02698595.2022.
2144457.

Ye:2008:BR

[YTP08] Feng Ye, Emma Tobin, and Isabelle Peschard. Book reviews. International Studies in the Philosophy of Science, 22(2):227–236, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698590802496805.

Zemplen:2013:HOG

[Zem13] Gábor Á. Zemplén. A history of optics from Greek antiquity to the nineteenth century. International Studies in the Philosophy of Science, 27(4):450–453, 2013. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2013.868189.

Zenker:2016:CTS

[ZG16] Frank Zenker and Peter Gärdenfors. Continuity of theory structure: a conceptual spaces approach. International Studies in the Philosophy of Science, 30(4):343–360, 2016. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2017.1331983.

Zhang:2015:BRL

[Zha15] Lu Zhang. Book review: Leibniz [Richard T. W. Arthur, Cambridge, Polity Press, 2014. xvi + 231 pp. ISBN 978-0-7456-5374-7, £55.00, EUR 68.80, US$69.95 (hardback); ISBN 978-0-7456-5375-4, £16.99, EUR 21.30, US$24.95 (paperback)]. International Studies in the Philosophy of Science, 29(2):215–218, 2015. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2015.1119423.

Zheng:1987:ISR

[Zhe88] Lan Zheng. Incommensurability and scientific rationality. International Studies in the Philosophy of Science, 2(2):227–236, 1987–1988. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com//doi/abs/
10.1080/02698598808573316.

Zhu:2017:EWE

[Zhu17] Qin Zhu. Ethics within engineering: An introduction. International Studies in the Philosophy of Science, 31(3):327–330, 2017. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/
10.1080/02698595.2018.1463698.

Zibakalam:1994:RDU

[Zib94] Saeid Zibakalam. Relativism due to underdetermination of theory by data. International Studies in the Philosophy of Science, 8(3):211–228, 1994. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com//doi/abs/
                                        10.1080/02698599408573496.

Zinkernagel:2008:DTB

[Zin08] Henrik Zinkernagel. Did time have a beginning? International Studies in the Philosophy of Science, 22(3):237–258, 2008. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698590802567282.

Zwicky:2012:WI

[Zwi12] Jan Zwicky. What is ineffable? International Studies in the Philosophy of Science, 26(2):197–217, 2012. CO-DEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

                                        //www.tandfonline.com/doi/full/
                                        10.1080/02698595.2012.703480.